### v1

# Granulometric Smoothing on Manifolds

Abstract
Given a random sample from a density function supported on a manifold $M$, a new method for the estimating highest density regions of the underlying population is introduced. The new proposal is based on the empirical version of the opening operator from mathematical morphology combined with a preliminary estimator of density function. This results in an estimator that is easy-to-compute since it simply consists of a list of carefully selected centers and a radius. The new estimator is shown to be consistent, and its convergence rate in terms of the Hausdorff distance are provided. All consistency results are established uniformly on the level of the set and for any Riemannian manifold $M$ satisfying mild assumptions. The applicability of the procedure is shown by means of some illustrative examples.
1 Introduction
In data analysis, one of the first questions that the statistician has to answer is: “Where are the data concentrated?” Some of the statistical techniques that provide an answer to the previous question include density estimation (Parzen,, 1962; Cao et al.,, 1994; Wand and Jones,, 1994; Loader,, 1996), mode hunting (Silverman,, 1981; Hartigan and Hartigan,, 1985; Minnotte and Scott,, 1993; Chaudhuri and Marron,, 1999; Burman and Polonik,, 2009; Ameijeiras-Alonso et al.,, 2019), support estimation (Devroye and Wise,, 1980; Mammen and Tsybakov,, 1995; Cuevas and Fraiman,, 1997; Klemelä,, 2004; Biau et al.,, 2008) or cluster analysis (Hartigan,, 1975; Cuevas et al.,, 2001; Chacón and Duong,, 2013; Wierzchoń and Kłopotek,, 2018). For most of these exploratory techniques, highest density regions (HDRs in short) play a relevant role. HDRs are the sets where the density function of the data exceeds a given level. More precisely, if $f$ is the population density function, HDRs are the sets such that
$L(\lambda)=f^{-1}\big{(}[\lambda,+\infty)\big{)},\quad\lambda>0.$ (1.1)
Modes can be seen as limiting cases of HDRs where the value of $\lambda$ converges to a local maxima of the density function. In practice, the specific value of the level $\lambda$ is often irrelevant and the interest lies in recovering the HDR that satisfies a given probability content. That is, given any $\gamma\in(0,1)$, the objective is to estimate the set $L(\lambda_{\gamma})$, where $\lambda_{\gamma}>0$ is such that
$\mathbb{P}\big{[}L(\lambda_{\gamma})\big{]}=1-\gamma.$ (1.2)
Values of $\gamma$ close to 00 lead to large HDRs that approach to the population support (since their probability content is close to one), and, therefore, their corresponding level $\lambda_{\gamma}$ is close to zero. Estimation of these sets provides a natural way of detecting outliers (Devroye and Wise,, 1980; Baíllo et al.,, 2001). Reciprocally, if $\gamma$ is close to $1$, the corresponding level $\lambda_{\gamma}$ is large and $L(\lambda_{\gamma})$ converges to the largest modes of the population. This connection allows to investigate multimodality via HDR estimation, see Müller and Sawitzki, (1991), Cuevas et al., (2001), and Rinaldo and Wasserman, (2010).
HDR estimation provides an easy-to-understand way to visualize a density in a general setting. Indeed, there are several practical problems regarding non-Euclidean data where HDR estimators may provide useful information. One example, which will be analyzed in detail, is the study of long-period comet orbits, a task previously addressed by Jupp et al., (2003). Long-period comets (comets with periods larger than 200 years) are assumed to originate in a celestial formation called Oort cloud. The Oort cloud (Weissman,, 1990) is a theoretical cloud of planetesimals supposed to surround the Solar System at very large distances. It is though to be sphere shaped, with the asteroids roughly uniformly distributed within it. The Oort cloud is influenced by galactic tidal forces and gravity perturbations due to passing stars, bringing some of these planetesimals close enough to the sun to become long-period comets (see Heisler and Tremaine,, 1986). This formation process explains why the orbits of these comets are nearly uniformly distributed, while short-period comets are highly concentrated around the ecliptic plane.
The uniformity of orbits can be assessed by testing the uniformity of their unit normal vectors. This way, each orbit orientation corresponds to a point on the 2-dimensional sphere $\mathbb{S}^{2}=\{x\in\mathbb{R}^{3}\colon\left\|x\right\|=1\}$, and one can address this problem by testing the uniformity of a given sample in $\mathbb{S}^{2}$. Figure 1 shows the normal unit vectors of all the long-period comets detected up to 30th of May 2022. A more detailed explanation of these data can be found in Section 6.1.
Figure 1: Comet orbits data represented on the sphere using orthogonal projections centered on the north pole (left) and south pole (right).
The problem of testing the uniformity of the long-period comets has been a widely considered example in statistical literature, and some authors like García-Portugués et al., (2023) noticed that their unit vectors do not follow a uniform spherical distribution. This is suggested to be caused by an observational bias (see Jupp et al.,, 2003). When astronomers look for objects in the space, they do not search uniformly in all the celestial sphere. Actually, they tend to focus on where it is more likely to find something, that is, the ecliptic plane. Since most of the observational effort is concentrated around it, long-period comets near the ecliptic plane are more likely to be detected, and therefore, they are overrepresented in the observed sample.
One possible way to check this observational bias hypothesis is via HDR estimation. If the HDRs of the observed normal unit vectors are located around the north and south poles of the sphere, then this will mean that comets near the ecliptic are overrepresented in the registered data, which agrees with the observational bias theory.
Furthermore, this analysis cannot be performed ignoring the geometry of the data support. See, for example, Figure 2, where the comet orbits orientations are represented in spherical coordinates, that is, each unit vector $v\in\mathbb{S}^{2}$ is represented as the pair $(\varphi,\psi)\in[-\pi,\pi]\times[-\pi/2,\pi/2]$ such that $v=\big{(}\cos(\varphi)\cos(\psi),\sin(\varphi)\cos(\psi),\sin(\psi)\big{)}^{\prime}$. This representation is known to distort distances in the sphere, specially in the regions near the north and south poles. Since these areas are precisely the areas of interest in this situation, any conclusion based on such a “flat” representation of the data will be misleading, and a HDR technique that takes into account the natural curvature and distance of the sphere should be employed.
Of course, this problematic is not specific of the sphere and data supported on any non-flat Riemannian manifold will suffer from it. This kind of datasets are a current trend in modern science with practical examples on the sphere, as previously illustrated, but also on the torus (Di Marzio et al.,, 2011, Ameijeiras-Alonso and Ley,, 2020), cylinder (Jammalamadaka and Lund,, 2006, Alonso-Pena et al.,, 2023) and polysphere (García-Portugués and Meilán-Vila,, 2023). Therefore, a global theory of HDR estimation is required, establishing consistency results for any Riemannian manifold $M$.
Figure 2: Comet orbits data represented in spherical coordinates. $\varphi$ is the longitude, whereas $\psi$ is the latitude. This representation distort areas and distances: points with $\psi$ close to ±$\pm\pi/2$ appear further apart than they really are.
Due to its practical utility, HDR estimation in the euclidean framework has been widely treated in the literature. Some references in this topic include Polonik, (1995), Tsybakov, (1997), Cadre, (2006) and Rodríguez-Casal and Saavedra-Nieves, (2022). However, HDR estimation in other contexts has not been considered until very recently. Cuevas et al., (2006) prove some theoretical consistency results for level set estimation in metric spaces (for a general function, not necessarily a density), and Saavedra-Nieves and Crujeiras, (2021) and Cholaquidis et al., (2022) give proposals of estimators for HDRs on the $d$-dimensional sphere and on a compact manifold, respectively.
To the best of the authors’ knowledge, all contributions on HDR estimation in a general framework found in the literature follow a plug-in approach. That is, they propose an estimator for the underlying density $f$, say $f_{n}$, and then the HDR is reconstructed with the set $f_{n}^{-1}\big{(}[\lambda,+\infty)\big{)}$. This estimation technique has some advantages, since its asymptotic properties can be related to those of $f_{n}$ as estimator of $f$. However, this approach has some issues in practice. For example, visualizing this estimator requires to evaluate $f_{n}$ in a large grid of values, which might be computationally expensive. In addition, sometimes one is interested not in the set itself, but in some of its characteristics, like its border or the number of connected components, and computing those for $f_{n}^{-1}\big{(}[\lambda,+\infty)\big{)}$ is not straightforward. Finally, this approach ignores the inner geometry of the problem: if one knows that the HDRs of $f$ fulfill some geometric property (e.g. some shape condition), there is no guarantee that the plug-in estimator will satisfy it too.
In the euclidean context geometrically based/conditioned HDR estimators have been proposed in the literature. For instance, an HDR estimator of a convex HDR should intuitively be convex too. To do that, one possibility is to maximize the empirical estimator of the excess mass within a given family of sets $\mathcal{D}$, see Müller and Sawitzki, (1991) and Polonik, (1995). Of course, this approach guarantees that the HDR estimator satisfies the required geometric restriction. Nevertheless, for a general family of sets the optimization problem is not feasible. This fact hampers the utilization of this estimation technique for most practical purposes.
To overcome this issue, an easy to compute new estimator for HDRs on Riemannian manifolds is introduced in this paper. The new approach combines an estimator of the underlying density with some a priori geometric information, which is appropriately included in the estimation method to simplify the final form of the estimator. More specifically, the new proposal can be viewed as an extension of the euclidean HDRs estimation technique, called granulometric smoothing, introduced by Walther, (1997) to Riemannian manifolds. It is assumed that the population HDR satisfies some smoothness assumptions which can be established in terms of some simple geometrical operations from mathematical morphology, see Serra, (1982). The proposed estimator is easy to compute in the very general framework of data on manifolds such as the sphere or the torus. Results are established for a general Riemannian manifold $M$.
The distribution of this paper is as follows. First, Section 2 includes the basic notation and definitions that are required throughout the manuscript. Then, Section 3 introduces the new proposal of HDR estimator. In Section 4 the consistency of the new estimator is proved, and its convergence rates (uniform on $\lambda$) are derived. Section 5 deals with the problem of estimating a HDR when, instead of a level $\lambda$, a probability content is given. A general procedure for estimating $\lambda_{\gamma}$ is proposed and analyzed. Section 6 illustrates the proposed estimator with two real-data examples. Finally, Section 7 provides a final discussion and future work.
Three appendices collect all the proofs. Appendix A includes some technical results regarding Minkowski operations in manifolds that are required along the proofs. Appendix B contains all the proofs related with the consistency and convergence rates of the proposed HDR estimator. Appendix C investigates the assumptions that are required on the manifold $M$ and the HDRs of the population for guaranteeing the consistency of the estimator. Particularly, in Appendix C, these assumptions are shown to hold for any manifold with non-negative and bounded sectional curvature and any differentiable density function, supporting the generality of our method.
2 Preliminaries and notation
The HDR estimator proposed by Walther, (1997) is based on Minkowski operations. So, extending this estimation method to Riemannian manifolds requires the generalization of these operations to the manifold setting. This section provides an appropriate generalization of these concepts jointly with some basic notation that will be used throughout this manuscript.
2.1 Riemannian manifolds and density functions
Let $M$ be a connected Riemannian manifold with dimension $d\in\mathbb{N}$. Given a point $x\in M$, we denote by $T_{x}M$ the tangent space to $M$ at $x$. Given any $v\in T_{x}M$, its norm is denoted by $\left\|v\right\|_{M}$ (or $\left\|v\right\|$ when there is no need to specify the manifold).
We denote by $\mathrm{Vol}_{M}$ the Riemann-Lebesgue volume measure of $M$ (see Amann and Escher,, 2009, Ch. XII). $\mathrm{Vol}_{M}$ is a Radon measure over the $\sigma$-algebra of Borel sets of $M$, and it allows to extend the concept of density function to the manifold setting.
Let $(\Omega,\mathcal{A},\mathbb{P})$ be a probability space and $X:\Omega\rightarrow M$ a Borel-measurable function. $X$ is called a random point of $M$. $X$ has density function $f$, where $f:M\rightarrow[0,+\infty)$ is a Borel-measurable function, if
[…]
for all Borel sets $A\subset M$. Actually, the Radon-Nikodym theorem (see Billingsley,, 1995, Th. 32.2) ensures that $X$ has a density function if and only if the probability measure induced by $X$ is absolutely continuous with respect to $\mathrm{Vol}_{M}$, and the density function $f$ coincides with the Radon–Nikodym derivative.
2.2 Distance and topology
Let $\mathcal{d}(\cdot,\cdot)$ be the geodesic distance of $M$ (note that the geodesic distance function is well defined since $M$ is connected, see Lee,, 2012, Ch. 13). The open and closed ball of center $x\in M$ and radius $r>0$ will be denoted as $B_{r}(x)$ and $B_{r}[x]$, respectively. That is, given a point $x\in M$ and a radius $r>0$, define
$B_{r}(x)=\left\{y\in M\colon\mathcal{d}(x,y)<r\right\},\qquad B_{r}[x]=\left\{y\in M\colon\mathcal{d}(x,y)\leq r\right\}.$
[…]
2.3 Minkowski operations
Given a subset $A\subset M$ and a radius $r>0$, the Minkowski sum and difference are defined as
[…]
where $A^{c}$ is the complement of the set $A$. The notation of the Minkowski operations is directly taken from Walther, (1997). Note that there is a straightforward relation between these two operations via the De Morgan’s laws:
$(A\ominus rB)^{c}=A^{c}\oplus rB.$ (2.4)
2.4 Distance between two sets and between a set and a point
Let $A_{1},A_{2}\subset M$ be two non empty subsets of $M$. The point distance can be used to define a distance between two sets, as
$\mathcal{d}(A_{1},A_{2})=\inf\left\{\mathcal{d}(a_{1},a_{2})\colon a_{1}\in A_{1},a_{2}\in A_{2}\right\}.$
One can define the distance between a point $x\in M$ and a non empty set $A_{1}\subset M$ using the previous definition:
$\mathcal{d}(x,A_{1})=\mathcal{d}\big{(}\{x\},A_{1}\big{)}=\inf\{\mathcal{d}(x,a_{1})\colon a_{1}\in A_{1}\}.$
2.5 Hausdorff distance
The distance between sets introduced in Section 2.4 is indeed useful, but it does not provide a metric space structure to the set
$F(M)=\big{\{}A\subset M\colon A\textit{ is compact and non empty}\big{\}}.$
And, since our aim is to make inference with the subsets of $M$, a metric space structure on $F(M)$ is required. We must then introduce a new distance between two sets. Given two compact and non empty subsets of $M$, $A_{1},A_{2}\in F(M)$, the Hausdorff distance between $A_{1}$ and $A_{2}$ is given by
$\mathcal{d}_{H}(A_{1},A_{2})=\inf\big{\{}r>0\colon A_{1}\subset A_{2}\oplus rB\textit{ and }A_{2}\subset A_{1}\oplus rB\big{\}}.$
3 The new estimator
Let $X$ be a random point on a manifold $M$ with density $f$. The main objective of this section is to introduce an estimator of the HDRs of $f$, that is, given a constant $\lambda$ such that $0<\lambda<\sup f$, the aim is to recover the set
$L(\lambda)=f^{-1}\big{(}[\lambda,+\infty)\big{)},$ (3.1)
from a finite i.i.d. sample of $X$.
As it has been pointed out in the Introduction, the new proposal does not follow a plug-in approach. Instead, the true HDRs are supposed to satisfy some shape conditions, and this information is included in the estimation method. These shape conditions are based on the assumptions made by Walther, (1997) for HDR estimation on Euclidean spaces, that are strongly related to the notion of “granulometry” introduced by Matheron, (1975). Given a subset $A\subset M$ and a radius $r>0$, define the functional
$\Psi_{r}(A)=\bigcup_{B_{r}[x]\subset A}B_{r}[x]=(A\ominus rB)\oplus rB.$ (3.2)
The shape conditions required by the new estimation method can be expressed in terms of this functional $\Psi_{r}$. Precisely, the HDRs are assumed to be fixed-points of $\Psi_{r}$, i.e., it is assumed that there exists a positive constant $r$ such that
$L(\lambda)=\Psi_{r}\big{[}L(\lambda)\big{]}=\bigcup_{B_{r}[x]\subset L(\lambda)}B_{r}[x]$ (3.3)
for some levels $\lambda>0$. Hence, the HDRs can be written as the union of closed balls of the same positive radius $r$. The proposed estimation method takes advantage of this shape condition, recovering $L(\lambda)$ as a union of balls of the same radius too. The idea is to “inflate” the high density points of the sample, where high and low density points are distinguished employing a density function estimator.
Let $\mathcal{X}_{n}=\{X_{1},\ldots,X_{n}\}$ be an i.i.d. sample from $X$. The estimation technique consists of the following steps.
  1. Step 1:
    First, consider an estimator of the underlying density, say $f_{n}$, and an estimator of the radius $r$, namely $r_{n}$.
  2. Step 2:
    Divide the sample $\mathcal{X}_{n}$ into low and high density points using $f_{n}$:
    $\mathcal{X}^{+}_{n}(\lambda)=\{X_{i}\colon f_{n}(X_{i})\geq\lambda\},\qquad\mathcal{X}^{-}_{n}(\lambda)=\{X_{i}\colon f_{n}(X_{i})<\lambda\}.$
  3. Step 3:
    The proposed estimator of $L(\lambda)$ is
    $L_{n}(\lambda)$ $=\Big{[}\mathcal{X}^{+}_{n}(\lambda)\cap\big{(}\mathcal{X}^{-}_{n}(\lambda)\oplus r_{n}B\big{)}^{c}\Big{]}\oplus r_{n}B$
    $=\bigcup_{\scriptsize\begin{array}[]{c}X_{i}\in\mathcal{X}^{+}_{n}(\lambda)\\ B_{r_{n}}[X_{i}]\cap\mathcal{X}^{-}_{n}(\lambda)=\emptyset\end{array}}B_{r_{n}}[X_{i}].$ (3.6)
It follows from the definition of the Minkowski difference in (2.3) that $L(\lambda)\ominus rB\subset L(\lambda)$. Hence, if $L(\lambda)=\Psi_{r}\big{[}L(\lambda)\big{]}$, then the definition of $\Psi_{r}$ in (3.2) jointly with (2.4) yield that
$L(\lambda)=\Big{[}L(\lambda)\cap\big{(}L(\lambda)^{c}\oplus rB\big{)}^{c}\Big{]}\oplus rB.$ (3.7)
The estimator $L_{n}(\lambda)$ defined in (3.6) exploits the shape condition (3.7) to recover $L(\lambda)$ by using $\mathcal{X}^{+}_{n}(\lambda)$ and $\mathcal{X}^{-}_{n}(\lambda)$ as discretizations of the sets $L(\lambda)$ and $L(\lambda)^{c}$, respectively. Therefore, provided that $f_{n}$ and $r_{n}$ are consistent estimators of $f$ and $r$, $L_{n}(\lambda)$ will eventually recover $L(\lambda)$.
Figure 3: HDR estimation of a mixture of two von Mises-Fisher distributions. All three graphs show the border of the estimator $L_{n}(\lambda)$ with $\lambda=0.45$ for a mixture of the distributions $\mathrm{M}_{2}(\mu_{1},10)$ and $\mathrm{M}_{2}(\mu_{2},10)$ with equal weights (see equation (3.8) for a definition of $\mu_{1}$ and $\mu_{2}$). Three sample sizes are considered, one for each graph: $n=400$ (a), $n=800$ (b) and $n=1600$ (c). The border of $L_{n}(\lambda)$ is plotted in red in all graphs, whereas the border of the true HDR $L(\lambda)$ is depicted as a dashed blue line. For reference, the points of the subsamples $\mathcal{X}^{+}_{n}(\lambda)$ and $\mathcal{X}^{-}_{n}(\lambda)$ are shown in the graphs as ‘$+$’ and $\circ$’, respectively. In all cases, $r_{n}=0.05$ and $f_{n}$ is the kernel density estimator with von Mises-Fisher kernel (García–Portugués,, 2013) and concentration parameter chosen with cross-validation.
Figure 3 illustrates the behavior in practice of this HDR estimator. The three plots show $L_{n}(\lambda)$ with $\lambda=0.45$ for a mixture of two different von Mises-Fisher distributions. The mixture consists of the distributions $\mathrm{M}_{2}(\mu_{1},10)$ and $\mathrm{M}_{2}(\mu_{2},10)$ with equal weights, where
$\begin{array}[]{l}\mu_{1}=\big{(}\cos(-\pi/6),\sin(-\pi/6),0\big{)}^{\prime},\\ \mu_{2}=\big{(}\cos(\pi/6)\cos(\pi/6),\cos(\pi/6)\sin(\pi/6),\sin(\pi/6)\big{)}^{\prime}.\end{array}$ (3.8)
The sample size is different in each plot, with $n=400$ in Figure 3(a), $n=800$ in Figure 3(b), and $n=1600$ in Figure 3(c). In all cases, $r_{n}=0.05$ and $f_{n}$ is the kernel density estimator with von Mises-Fisher kernel (García–Portugués,, 2013) and concentration parameter chosen by cross-validation. For the smallest size, the estimator $L_{n}(\lambda)$ (red line) is very different from the true HDR $L(\lambda)$ (blue line), with an irregular border and empty spaces within it. As $n$ grows, $\mathcal{X}^{+}_{n}(\lambda)$ and $\mathcal{X}^{-}_{n}(\lambda)$ become denser discretizations of $L(\lambda)$ and $L(\lambda)^{c}$ and, consequently, $L_{n}(\lambda)$ approaches $L(\lambda)$.
It should be remarked that the estimator $L_{n}(\lambda)$ does not rely solely on the shape assumption, but also on two auxiliary estimators: an estimator of the density function, $f_{n}$, and an estimator of the radius, $r_{n}$. There are several proposals for density estimators in the manifold setting. For example, if $M$ is a $d$-dimensional submanifold of the Euclidean space $\mathbb{R}^{D}$ with $d<D$, then Wu and Wu, (2021) propose the following estimator of $f$
$f_{n}(x)=\frac{1}{nh_{n}^{d}}\mathcal{K}\bigg{(}\frac{\left\|x-X_{i}\right\|_{\mathbb{R}^{D}}}{h_{n}}\bigg{)},$ (3.9)
where the bandwidth $h_{n}$ converges to zero at a suitable rate and $\mathcal{K}:[0,+\infty)\rightarrow\mathbb{R}$ is a kernel function. If $f$ is Hölder continuous and $\mathcal{K}$ satisfies some regularity conditions, Wu and Wu, (2021) proved the global consistency of such an estimator and computed its convergence rate.
On the contrary, estimating the radius $r$ has received almost no attention in statistical literature, even in the Euclidean setting. To the best of our knowledge, the only proposal for Euclidean data is provided by Rodríguez-Casal and Saavedra-Nieves, (2022). Their proposed estimator for $r$ relies on the relation between $r$ and the gradient of $f$, namely $\nabla f$, shown by Walther, (1997). However, if $f$ is a density defined on a Riemannian manifold, the relation between $\nabla f$ and $r$ remains unclear, and therefore this estimator cannot be directly extended to the manifold setting. Consequently, no data-driven method for estimating $r$ in Riemannian manifolds has been developed yet, and the statistician may choose $r_{n}$ with their own criterion.
Trying to overcome this issue, the theoretical results of this chapter do not depend on a particular choice of the estimators $r_{n}$ and $f_{n}$. Instead, in the derived results both items are assumed to fulfill some mild conditions, namely that $\Psi_{r_{n}}\big{[}L(\lambda)\big{]}=L(\lambda)$ and $f_{n}$ is globally consistent estimator of $f$. Then, the consistency of the estimator $L_{n}(\lambda)$ based on those is shown, and its consistency rate is derived as a function of the global error of $f_{n}$. This approach broadens the applicability of our results, since the consistency of $L_{n}(\lambda)$ is guaranteed for a large variety of estimators $f_{n}$ and $r_{n}$. Besides, if a new estimator of $f$ converging faster than the one defined on (3.9) is found, then our results will also guarantee that the estimator $L_{n}(\lambda)$ relying on such a new estimator will converge faster to $L(\lambda)$.
4 Consistency of the new proposal
In this section, the consistency of the estimator $L_{n}(\lambda)$ defined at (3.6) is derived. First, Sections 4.1 and 4.2 detail the required conditions for the consistency of $L_{n}(\lambda)$. Specifically, Section 4.1 describes the manifold setting that is addressed throughout this chapter, while Section 4.2 details the shape conditions on the HDRs of $f$ that are required for the consistence of the HDR estimator. Finally, in Section 4.3 the consistency of $L_{n}(\lambda)$ is shown and its asymptotic convergence rate is derived.
4.1 Assumptions on the manifold $M$
As we have seen in Section 3, $L_{n}(\lambda)$ uses the Minkowski operations defined in (2.2) and (2.3) to recover the HDRs of the underlying density $f$. To guarantee the good behavior of the Minkowski operations, and therefore the consistency of $L_{n}(\lambda)$, one needs to impose the following assumptions on the manifold $M$.
Assumptions M.
$M$ is a $d$-dimensional Riemannian manifold verifying:
  1. (M1)
    $M$ is complete and connected.
  2. (M2)
    Let $r_{1},r_{2}>0$. Let $x_{1},x_{2}\in M$ such that $\mathcal{d}(x_{1},x_{2})<r_{1}+r_{2}$. There exist two constants, $a,\rho>0$, (both not depending on $x_{1}$ and $x_{2}$) such that
    $\mathrm{Vol}_{M}\big{(}B_{r_{1}}[x_{1}]\cap B_{r_{2}}[x_{2}]\big{)}\geq a\min\{r_{1},r_{2},\rho\}^{(d-1)/2}\min\{\varepsilon,\rho\}^{(d+1)/2}.$
    for all $\varepsilon\leq r_{1}+r_{2}-\mathcal{d}(x_{1},x_{2})$.
  3. (M3)
    For every bounded subset, $A\subset M$, define:
    $D(\varepsilon,A)=\max\left\{\mathrm{card}(T)\colon T\subset A,\mathcal{d}(x,y)>\varepsilon\textit{ for all }x,y\in T,x\neq y\right\}.$
    There exist two constants, $\rho>0$ (which does not depend on $A$) and $K>0$, such that $D(\varepsilon,A)\leq K\varepsilon^{-d}$ for all $\varepsilon<\rho$.
The role of Assumption (M1) is twofold: the connectedness of $M$ ensures that the geodesic distance is well defined, while the completeness guarantees a good behavior of the Minkowski operations and the mapping $\Psi_{r}$ (see Appendix A). Assumption (M2) is required for showing the consistency of $L_{n}(\lambda)$. By (3.6), $L_{n}(\lambda)$ is a union of balls of the same radius. So, in order to compute the convergence rate of $L_{n}(\lambda)$, one needs to find a lower bound of the probability of the intersection of two balls. Since probability and volume are related through the density function $f$, Assumption (M2) guarantees the existence of such a bound. Finally, Assumption (M3) is a technical condition that simplifies the proofs. Actually, it is not independent of the others: if $M$ fulfills Assumption (M2), then it also satisfies (M3).
Proposition 4.1.
Suppose that $M$ is a $d$-dimensional Riemannian manifold satisfying Assumption (M2). Let $A$ be a bounded subset of $M$. Then, there exists a constant $K$ such that
$D(\varepsilon,A)\leq K\varepsilon^{-d}$
for all $\varepsilon\leq\rho$, where $\rho>0$ is the same constant than Assumption (M2).
Therefore, $M$ satisfies Assumption (M3).
At first sight, Assumption (M2) may seem an ad-hoc condition for this problem. Nevertheless, any manifold satisfying some mild conditions verify Assumption (M2), as the next result shows.
Theorem 4.1.
Let $M$ a complete and connected $d$-dimensional Riemannian manifold satisfying:
  1. 1.
    All sectional curvatures of $M$, $\kappa_{M}$, are non-negative and bounded:
    $0\leq\kappa_{M}<\kappa.$
  2. 2.
    $M$ has positive injectivity radius: $i_{M}>0$.
Then $M$ satisfies Assumption (M2).
Theorem 4.1 jointly with Proposition 4.1 ensure that any manifold with bounded and non negative sectional curvatures fulfill Assumptions M. Of course, this includes the Euclidean space $\mathbb{R}^{d}$, but also manifolds like the hypersphere $\mathbb{S}^{d}$, or the $d$-dimensional torus $\mathbb{T}^{d}$. The proofs of these two results can be found in Appendix C.1.
4.2 Assumptions on the HDRs
As discussed in Section 3, at least the condition $L(\lambda)=\Psi_{r}\big{[}L(\lambda)\big{]}$ is required for the consistency of the new proposal. In practice, stronger assumptions on the HDRs of $f$ are made.
Assumptions L.
Let $M$ be a connected Riemannian manifold, $f:M\rightarrow\mathbb{R}$ a continuous density function, and $l$ and $u$ two constants such that $0<l\leq u<\sup f$. There exist a radius, $r>0$, and two positive constants $k$ and $\delta$, with $k>0$ and $20<\delta<r/$ 2, verifying
  1. (L1)
    $L(\lambda)=\Psi_{r}[L(\lambda)]$ for all $\lambda\in[l-\delta,u+\delta]$.
  2. (L2)
    $L(\lambda)^{c}=\Psi_{\delta}\big{[}L(\lambda)^{c}\big{]}$ for all $\lambda\in[l-\delta,u+\delta]$.
  3. (L3)
    $L(\lambda)$ is a compact set for all $\lambda\in[l-\delta,u+\delta]$.
  4. (L4)
    $\mathcal{d}_{H}\big{(}L(\lambda),L(\lambda\pm\eta)\big{)}\leq k\eta$ for all $\lambda\in[l,u]$ when $\eta\in[0,\delta]$.
Remark 4.1.
The quantity $\delta$ in Assumptions (L1)(L4) should be understood as a “small but positive constant”. Its specific value is not important, since it will not affect the convergence rate of the estimator $L_{n}(\lambda)$.
Furthermore, assuming that $\delta$ is equal for all of Assumptions A is not restrictive. For example, if $f$ verifies (L1) with $\delta_{1}$, (L2) with $\delta_{2}$, (L3) with $\delta_{3}$, and (L4) with $\delta_{4}$; then $f$ also verifies all four assumptions with $\delta=\min\{\delta_{1},\delta_{2},\delta_{3},\delta_{4}\}$. Then, they will be assumed as equal to simplify notation.
The necessity of Assumption (L1) for the consistency of $L_{n}(\lambda)$ is clear, since the estimation technique relies entirely on it. Assumption (L3) serves merely as a technical condition to guarantee that the Hausdorff distance is well defined over the HDRs of $f$. The consistency of $L_{n}(\lambda)$ is going to be derived in terms of the Hausdorff distance. Hence, it is required that $\mathcal{d}_{H}$ is well defined on the highest density regions of $f$, and $\mathcal{d}_{H}$ is only defined for compact sets.
As mentioned in Section 3, $L_{n}(\lambda)$ uses $\mathcal{X}^{+}_{n}(\lambda)$ and $\mathcal{X}^{-}_{n}(\lambda)$ as discretizations of the sets $L(\lambda)$ and $L(\lambda)^{c}$, respectively. The intuition is that, as $n$ gets larger, $\mathcal{X}^{+}_{n}(\lambda)$ and $\mathcal{X}^{-}_{n}(\lambda)$ should become finer and finer discretizations of $L(\lambda)$ and $L(\lambda)^{c}$, and, eventually, $L_{n}(\lambda)$ will recover $L(\lambda)$. However, $\mathcal{X}^{+}_{n}(\lambda)$ and $\mathcal{X}^{-}_{n}(\lambda)$ depend on the estimator $f_{n}$ and not on the underlying density $f$. So, without any further assumptions, it is not guaranteed that $\mathcal{X}^{+}_{n}(\lambda)\subset L(\lambda)$ and $\mathcal{X}^{-}_{n}(\lambda)\subset L(\lambda)^{c}$. Assumption (L4) solves this problem, ensuring that, if $f_{n}$ is a consistent estimator of $f$, $\mathcal{X}^{+}_{n}(\lambda)\subset L(\lambda)$ and $\mathcal{X}^{-}_{n}(\lambda)\subset L(\lambda)^{c}$ will eventually hold. Thus, $\mathcal{X}^{+}_{n}(\lambda)$ and $\mathcal{X}^{-}_{n}(\lambda)$ will be valid discretizations of $L(\lambda)$ and $L(\lambda)^{c}$ for $n$ large.
In addition, for the consistency of $L_{n}(\lambda)$ both discretizations are required to become “fine enough” as the sample size grows. That is, $\mathcal{X}^{+}_{n}(\lambda)$ and $\mathcal{X}^{-}_{n}(\lambda)$ are required to be eventually dense in $L(\lambda)$ and $L(\lambda)^{c}$, respectively. For $\mathcal{X}^{-}_{n}(\lambda)$, this means that, given any point $x\in L(\lambda)^{c}$, there should be a point in $\mathcal{X}^{-}_{n}(\lambda)$ arbitrarily close to $x$. If one assumes (L2), then there exists a ball $B_{\delta}[p]\in L(\lambda)^{c}$ such that $x\in B_{\delta}[p]$ and $\mathbb{P}\big{(}B_{\delta}[p]\big{)}>0$, and this will eventually occur. Therefore, Assumption (L2) is needed to guarantee that $\mathcal{X}^{-}_{n}(\lambda)$ will become dense in $L(\lambda)^{c}$. For $\mathcal{X}^{+}_{n}(\lambda)$ to be dense in $L(\lambda)$ no further assumptions are required, since Assumption (L1) will ensure this by similar arguments.
$L(\lambda)$
$L(\lambda)$
Figure 4: Example of a HDR $L(\lambda)$ satisfying (left) and not satisfying (right) Assumptions (L1) and (L2).
Actually, Assumptions L can be viewed as smoothness conditions on the HDRs $L(\lambda)$. One readily see this for Assumption (L4), since it is a Lipschitz condition on the mapping $L$: (L4) states that $L$ is a (locally) Lipschitz continuous function with respect the Hausdorff distance. Assumptions (L1) and (L2) are smoothness conditions in a more geometric sense. If both $L(\lambda)$ and $L(\lambda)^{c}$ can be recovered as a union of balls of the same fixed radius, then the border of $L(\lambda)$ is smooth, without sharp edges (see Figure 4). This fact implies that the border of the sets $L(\lambda)$ are regular submanifolds of $M$.
Since they are smoothness conditions on the HDRs of $f$, it is no surprise that they translate into smoothness conditions on the density function $f$. For example, if $f$ is continuous, then the HDRs are guaranteed to be compact and Assumption (L3) holds.
Proposition 4.2.
Let $M$ be a Riemannian manifold under Assumptions M, let $\lambda>0$ and $f:M\rightarrow\mathbb{R}$ a continuous density function. Suppose that $L(\lambda)=\Psi_{\delta}\big{[}L(\lambda)\big{]}$ for some $\delta>0$. Then, $L(\lambda)$ is a compact set.
If $f$ is a differentiable function, the gradient $\nabla f$ is the direction in which $f$ varies faster, and $\left\|\nabla f\right\|_{M}$ measures the rate of variation in that direction. Hence, controlling $\left\|\nabla f\right\|_{M}$ allows to control the distance between the HDRs, as the next result shows.
Proposition 4.3.
Let $M$ be a $d$-dimensional Riemannian manifold under Assumptions M, and $f:M\rightarrow\mathbb{R}$ a density function. Let $0<\delta<l\leq u<\sup f$, and let $U$ be an open set such that $f^{-1}\big{(}[l-\delta,u+\delta]\big{)}\subset U$.
Suppose that $f$ is continuously differentiable on $U$ and that there exists a positive constant $c>0$ such that $\left\|\nabla f\right\|_{M}\geq c$ in $U$. Then:
$\mathcal{d}_{H}\big{(}L(\lambda),L(\lambda\pm\eta)\big{)}\leq\frac{\eta}{c}$
for all $\eta\leq\delta$ and $\lambda\in[l,u]$.
The proofs of Propositions 4.2 and 4.3 are collected in Appendix C.2. These results ensure that any sufficiently smooth density function $f$ verifies Assumptions (L3) and (L4). If $M=\mathbb{R}^{d}$, Walther, (1997) showed that this is also true for Assumptions (L1) and (L2) (see Walther,, 1997, Th. 2). In the manifold setting, the relation between the differentiability of $f$ and Assumptions (L1) and (L2) is unclear, yet it is conjectured that all densities $f$ under the conditions of Proposition 4.3 will satisfy them too.
4.3 Convergence rate of the new estimator
Under Assumptions M and Assumptions L, the uniform behavior of the $L_{n}(\lambda)$ estimator can be studied. Theorem 4.2, which is the our main result, ensures that $L_{n}(\lambda)$ is a consistent estimator of $L(\lambda)$, and its convergence rate (in Hausdorff distance) is derived in terms of the global error of the density estimator $f_{n}$. The proof of this result is contained in Appendix B.1.
Theorem 4.2.
Let $M$ be a $d$-dimensional Riemannian manifold under Assumptions M. Let $X$ be a random point from $M$ with density function $f$, and suppose that $f$ satisfies Assumptions L. Let $\mathcal{X}_{n}=\{X_{1},\ldots,X_{n}\}$ be an i.i.d. sample from $X$, and let $f_{n}$ be an estimator of $f$ verifying:
$D_{n}=\sup_{x\in M}\left|f(x)-f_{n}(x)\right|\rightarrow 0$
almost surely. Furthermore, let $r_{n}$ be a sequence of positive random variables verifying
$\delta<r_{n}\leq r-\delta$ (4.1)
e.a.s., where $\delta>0$ is the same constant as Assumptions L.
Then, if $L_{n}(\lambda)$ is the (random) set defined in (3.6),
$\sup_{\lambda\in[l,u]}\mathcal{d}_{H}\big{(}L(\lambda),L_{n}(\lambda)\big{)}=O\left(\max\left\{D_{n},\left(\frac{\log n}{n}\right)^{\frac{2}{d+1}}\right\}\right)$
almost surely. That is, with probability $1$, there exists $n_{0}\in\mathbb{N}$ such that, if $n\geq n_{0}$, then
$\sup_{\lambda\in[l,u]}\mathcal{d}_{H}\big{(}L(\lambda),L_{n}(\lambda)\big{)}\leq C\max\left\{D_{n},\left(\frac{\log n}{n}\right)^{\frac{2}{d+1}}\right\}$
for some constant $C$.
If the density function $f$ is a Hölder continuous function with exponent $\alpha\in(0,1]$, and $f_{n}$ is the kernel density estimator defined in (3.9) with bandwidth $h_{n}$ of order $(\log(n)/n)^{1/(d+2\epsilon)}$ with $\epsilon>0$, Theorem 2.4 of Wu and Wu, (2021) guarantees that, for some kernels $\mathcal{K}$,
$D_{n}=O\bigg{(}\left(\frac{\log n}{n}\right)^{\frac{\alpha\epsilon}{d+2\epsilon}}\bigg{)}$
almost surely. Hence, in this case, Theorem 4.2 yields
$\sup_{\lambda\in[l,u]}\mathcal{d}_{H}\big{(}L(\lambda),L_{n}(\lambda)\big{)}=O\left(\max\left\{\left(\frac{\log n}{n}\right)^{\frac{\alpha\epsilon}{d+2\epsilon}},\left(\frac{\log n}{n}\right)^{\frac{2}{d+1}}\right\}\right)$
almost surely.
Theorem 4.2, jointly with Assumptions M and L, imposes some restrictions on the primary estimators $f_{n}$ and $r_{n}$. The condition required for $f_{n}$ was already discussed in Section 3: $f_{n}$ must be globally consistent, otherwise $L_{n}(\lambda)$ will not recover $L(\lambda)$. However, the restriction imposed on $r_{n}$ is stronger than the one introduced in Section 3. When the construction of $L_{n}(\lambda)$ was motivated, the only required condition for $r_{n}$ was $\Psi_{r_{n}}\big{[}L(\lambda)\big{]}=L(\lambda)$, and this holds whenever $r_{n}\leq r$ (see Proposition A.2d in Appendix A). Nevertheless, in equation (4.1) $r_{n}$ is forced to be strictly smaller than $r$. This condition is required to ensure that $\mathbb{P}\big{[}L(\lambda)\ominus r_{n}B\big{]}\neq 0$, otherwise $L_{n}(\lambda)$ will not recover $L(\lambda)$.
Moreover, equation (4.1) forces the sequence of random variables $r_{n}$ to not converge to zero, but this is not thought to be a limitation in practice. For instance, the proposal of $r_{n}$ given by Rodríguez-Casal and Saavedra-Nieves, (2022) for Euclidean data does not converge to zero if $f$ is smooth enough. Analogously, it is expected that any reasonable proposal of $r_{n}$ in the manifold setting will be also bounded away from zero and, consequently, Theorem 4.2 still ensures the consistency of the estimator $L_{n}(\lambda)$ based on such a choice of $r_{n}$.
5 HDRs by probability content
In Sections 3 and 4, we have studied HDRs as sets where the value of a given density function $f$ exceeds some threshold $\lambda$, as expressed in equation (3.1). This is convenient from a theoretical point of view, but the specific value of $\lambda$ is often irrelevant in practice. Instead, one is usually more interested in recovering the HDR verifying a (previously fixed) probability content. That is, given a probability $\gamma\in(0,1)$, the goal is to recover the set
$L(\lambda_{\gamma})=f^{-1}\big{(}[\lambda_{\gamma},+\infty)\big{)}=\{x\in M\colon f(x)\geq\lambda_{\gamma}\}$ (5.1)
where
$\lambda_{\gamma}=\sup\Big{\{}\lambda\in\mathbb{R}\colon\mathbb{P}\big{[}L(\lambda)\big{]}\geq 1-\gamma\Big{\}}.$ (5.2)
Thus, $L(\lambda_{\gamma})$ is the smallest HDR such that $\mathbb{P}\big{[}L(\lambda_{\gamma})\big{]}\geq 1-\gamma$. This perspective links HDR estimation with the problem of estimating the minimum volume set satisfying a probability content (see Polonik,, 1997).
Theorem 4.2 states that the convergence of the estimator $L_{n}(\lambda)$ is uniform with respect to the level $\lambda$. This fact allows to address the problem of recovering the set $L(\lambda_{\gamma})$ using this estimator via a plug-in process. First, one computes an estimator of $\lambda_{\gamma}$, namely $\tilde{\lambda}_{\gamma,n}$, and then, $L(\lambda_{\gamma})$ is recovered by $L_{n}(\tilde{\lambda}_{\gamma,n})$. In Section 5.1, the consistency of such a plug-in technique is studied, while Section 5.2 discusses the possible approaches to estimate the level $\lambda_{\gamma}$.
5.1 Consistency of the plug-in estimation technique
If $\tilde{\lambda}_{\gamma,n}$ is a consistent estimator of $\lambda_{\gamma}$, then $L_{n}(\tilde{\lambda}_{\gamma,n})$ is also a consistent estimator of $L(\lambda_{\gamma})$, as the following theorem proves.
Theorem 5.1.
Let $M$ be a $d$-dimensional Riemannian manifold under Assumptions M. Let $X$ be a random point from $M$ with density function $f$, and suppose that $f$ satisfies Assumptions L. Let $\mathcal{X}_{n}=\{X_{1},\ldots,X_{n}\}$ be an i.i.d. sample from $X$, and let $f_{n}$ be an estimator of $f$ verifying:
$D_{n}=\sup_{x\in M}\left|f(x)-f_{n}(x)\right|\rightarrow 0$
almost surely. Let $r_{n}$ be a sequence of positive random variables verifying
$\delta<r_{n}\leq r-\delta$
e.a.s, where $\delta>0$ is the same constant as Assumptions L.
Let $10<\underline{\gamma}\leq\overline{\gamma}<$ 1 such that $[\lambda_{\underline{\gamma}},\lambda_{\overline{\gamma}}]\subset(l,u)$. Let $\tilde{\lambda}_{\gamma,n}$ be an estimator of $\lambda_{\gamma}$ verifying:
$T_{n}=\sup_{\gamma\in[\underline{\gamma},\overline{\gamma}]}\left|\tilde{\lambda}_{\gamma,n}-\lambda_{\gamma}\right|\rightarrow 0$
almost surely.
Hence:
$\sup_{\gamma\in[\underline{\gamma},\overline{\gamma}]}\mathcal{d}_{H}\big{(}L(\lambda_{\gamma}),L_{n}(\tilde{\lambda}_{\gamma,n})\big{)}=O\left(\max\left\{D_{n},T_{n},\left(\frac{\log n}{n}\right)^{\frac{2}{d+1}}\right\}\right)$
almost surely.
Theorem 5.1 (see the proof in Appendix B.2) shows that $L_{n}(\tilde{\lambda}_{\gamma,n})$ is a consistent estimator of $L(\lambda_{\gamma})$ and its convergence rate depends on the global error of the primary estimator of the level $\tilde{\lambda}_{\gamma,n}$, in such a way that, if $\tilde{\lambda}_{\gamma,n}$ converges slowly to $\lambda_{\gamma}$, the estimator $L_{n}(\tilde{\lambda}_{\gamma,n})$ will then inherit that convergence rate and it will also converge slowly to $L(\lambda_{\gamma})$. Therefore, the choice of $\tilde{\lambda}_{\gamma,n}$ is crucial. Not only it must be a consistent estimator of the true level $\lambda_{\gamma}$, but it should converge as fast as possible to avoid any penalization on the convergence rate of estimator $L_{n}(\tilde{\lambda}_{\gamma,n})$. The next section explores this problem, investigating the possible options for the estimator $\tilde{\lambda}_{\gamma,n}$ found in the literature, and providing an estimator of the level that do not penalize the convergence rate of the final HDR estimator.
5.2 Estimation of the level $\lambda_{\gamma}$
As Theorem 5.1 states, estimating $L(\lambda_{\gamma})$ reduces to choosing a consistent estimator of the level $\lambda_{\gamma}$ defined in (5.2). Obviously, there is no unique choice, and one can select the most suitable estimator for the problem at hand. For example, Walther, (1997) proposes to estimate $\lambda_{\gamma}$ with
$\tilde{\lambda}_{\gamma,n}=\sup\Big{\{}\lambda\in\mathbb{R}\colon\mathbb{P}_{n}\big{[}L_{n}(\lambda)\big{]}\geq 1-\gamma\Big{\}}$ (5.3)
where $\mathbb{P}_{n}$ is the empirical probability distribution. This estimator requires to compute $L_{n}(\lambda)$ for a grid of values of $\lambda$, which is computationally intensive. Additionally, since equation (5.3) relies on the HDR estimator $L_{n}(\lambda)$, this particular choice of $\tilde{\lambda}_{\gamma,n}$ depends on the primary estimator of the radius $r_{n}$, and this fact makes it difficult to analyze theoretically.
Another possible choice is given by Cadre, (2006), who takes $\tilde{\lambda}_{\gamma,n}$ as the solution in $\lambda$ for the equation
[…]
Another proposal is to estimate $\lambda_{\gamma}$ with
$\hat{\lambda}_{\gamma,n}=\sup\left\{\lambda\in\mathbb{R}\colon\mathbb{P}_{n}\Big{[}f_{n}^{-1}\big{(}[\lambda,+\infty)\big{)}\Big{]}\geq 1-\gamma\right\},$ (5.5)
where $\mathbb{P}_{n}$ is the empirical probability distribution. This estimator, previously introduced by Hyndman, (1996) in the Euclidean setting, has several advantages with respect to the proposals given by Walther, (1997) and Cadre, (2006). For example, $\hat{\lambda}_{\gamma,n}$ does not depend on the estimator of the radius $r_{n}$ since equation (5.5) does not involve this parameter. Moreover:
$\mathbb{P}_{n}\Big{[}f_{n}^{-1}\big{(}[\lambda,+\infty)\big{)}\Big{]}=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\big{(}f_{n}(X_{i})\geq\lambda\big{)},$
and therefore
$\hat{\lambda}_{\gamma,n}=\sup\left\{\lambda\in\mathbb{R}\colon\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\big{(}f_{n}(X_{i})\geq\lambda\big{)}\geq 1-\gamma\right\}.$
So, the estimator $\hat{\lambda}_{\gamma,n}$ is the $\gamma$-quantile of the empirical distribution of the sample $f_{n}(X_{1}),\ldots,f_{n}(X_{n})$ and, consequently, $\hat{\lambda}_{\gamma,n}$ can be computed quite efficiently in practice.
In order to guarantee the uniform convergence of the estimator $\hat{\lambda}_{\gamma,n}$ defined in (5.5), a smoothing assumption on $\lambda_{\gamma}$ is required.
Assumption P.
Define $\lambda_{\gamma}=\sup\Big{\{}\lambda\in\mathbb{R}\colon\mathbb{P}\big{[}L(\lambda)\big{]}\geq 1-\gamma\Big{\}}.$ There exist $0<\underline{\gamma}\leq\overline{\gamma}$ such that
$\sup_{\gamma\in[\underline{\gamma},\overline{\gamma}]}\left|\lambda_{\gamma}-\lambda_{\gamma\pm\eta}\right|\leq k\eta$
for all $\eta\in[0,\delta]$, where $\delta$ and $k$ are the same constants as Assumptions L.
Assumption P states that $\lambda_{\gamma}$ is a (locally) Lipschitz continuous function of $\gamma$. Hence, it can be seen as another smoothness condition on the level sets of $f$. Like Assumptions (L3) and (L4), $f$ will verify Assumption P if it is smooth enough. More specifically, the next result holds (see Appendix C.3).
Proposition 5.1.
Let $M$ be a Riemannian manifold under Assumptions M, and let $f:M\rightarrow\mathbb{R}$ be a density function. Let $0<l\leq u<\sup f$, and let $U$ be an open set such that $f^{-1}\big{(}[l,u]\big{)}\subset U$. Suppose that $f$ is a $\mathcal{C}^{1}$ function in $U$ and there exists a positive constant $c>0$ satisfying such that $\left\|\nabla f\right\|_{M}\geq c$ in $U$.
Let $10<\underline{\gamma}\leq\overline{\gamma}<$ 1 and $\delta>0$ such that $l<\lambda_{\underline{\gamma}-\delta}\leq\lambda_{\overline{\gamma}+\delta}<u$. Then, there exists a constant $k>0$ such that
$\sup_{\gamma\in[\underline{\gamma},\overline{\gamma}]}\left|\lambda_{\gamma}-\lambda_{\gamma+\eta}\right|\leq k\left|\eta\right|,\quad\text{ for all }\eta\in[-\delta,\delta].$
Remark 5.1.
Notice that Proposition 5.1 does not establish any relation between $c$ and $k$, whereas Proposition 4.3 in Section 4.2 ensures that Assumption (L4) holds for $k=1/c$.
Under Assumption P, $\hat{\lambda}_{\gamma,n}$ in (5.5) is a globally consistent estimator of $\lambda_{\gamma}$ and its convergence rate is known, as the next theorem shows.
Theorem 5.2.
Let $M$ be a Riemannian manifold and $\mathcal{X}_{n}=\{X_{1},\ldots,X_{n}\}$ an i.i.d. sample from the random point $X$. Suppose that $X$ has a density function $f$ satisfying Assumption P. Let $f_{n}$ be an estimator of $f$ verifying:
$D_{n}=\sup_{x\in M}\left|f(x)-f_{n}(x)\right|\rightarrow 0$
almost surely.
Then
$\sup_{\gamma\in[\underline{\gamma},\overline{\gamma}]}\left|\lambda_{\gamma}-\hat{\lambda}_{\gamma,n}\right|=O\left(\max\left\{D_{n},\sqrt{\frac{\log n}{n}}\right\}\right)$
almost surely.
The proof of Theorem 5.2 can be found in Appendix B.2. From Theorems 5.1 and 5.2 it immediately follows that $L_{n}(\hat{\lambda}_{\gamma,n})$ is a consistent estimator of $L(\lambda_{\gamma})$ and its convergence rate is derived.
Theorem 5.3.
Let $M$ be a $d$-dimensional Riemannian manifold under Assumptions M. Let $X$ be a random point from $M$ with density function $f$, and suppose that $f$ satisfies Assumptions L and Assumption P with $[\lambda_{\underline{\gamma}},\lambda_{\overline{\gamma}}]\subset(l,u)$. Let $\mathcal{X}_{n}=\{X_{1},\ldots,X_{n}\}$ be an i.i.d. sample from $X$, and let $f_{n}$ be an estimator of $f$ verifying:
$D_{n}=\sup_{x\in M}\left|f(x)-f_{n}(x)\right|\rightarrow 0$
almost surely. Let $r_{n}$ be a sequence of positive random variables verifying
$\delta<r_{n}\leq r-\delta$
e.a.s.
Hence:
$\sup_{\gamma\in[\underline{\gamma},\overline{\gamma}]}\mathcal{d}_{H}\big{(}L(\lambda_{\gamma}),L_{n}(\hat{\lambda}_{\gamma,n})\big{)}=O\left(\max\left\{D_{n},\sqrt{\frac{\log n}{n}},\left(\frac{\log n}{n}\right)^{\frac{2}{d+1}}\right\}\right)$
almost surely.
The term $(\log(n)/n)^{1/2}$ in the convergence rate of $L_{n}(\hat{\lambda}_{\gamma,n})$ comes from the estimation of the level and does not suppose a penalization in practice, since it gets neglected by $D_{n}$ under some mild conditions. As already mentioned in Section 4.3, if $f$ is a Hölder continuous function with exponent $\alpha\in(0,1]$, and $f_{n}$ is kernel density estimator defined in (3.9) with bandwidth $h_{n}$ of order $(\log(n)/n)^{1/(d+2\epsilon)}$ with $\epsilon>0$, Theorem 2.4 of Wu and Wu, (2021) guarantees that, for some kernels $\mathcal{K}$,
$D_{n}=O\bigg{(}\left(\frac{\log n}{n}\right)^{\frac{\alpha\epsilon}{d+2\epsilon}}\bigg{)}$
almost surely, and, consequently, Theorem 5.3 ensures that
$\sup_{\gamma\in[\underline{\gamma},\overline{\gamma}]}\mathcal{d}_{H}\big{(}L(\lambda_{\gamma}),L_{n}(\hat{\lambda}_{\gamma,n})\big{)}=O\left(\max\left\{\left(\frac{\log n}{n}\right)^{\frac{\alpha\epsilon}{d+2\epsilon}},\left(\frac{\log n}{n}\right)^{\frac{2}{d+1}}\right\}\right)$
almost surely.
6 Real data illustration
The performance in practice of the new estimator $L_{n}(\hat{\lambda}_{\gamma,n})$ is illustrated with two real data examples. The manifold holding the data is different in each example. In the first case data are a point cloud in the 2–dimensional sphere ($M=\mathbb{S}^{2}$), whereas in the second data are registered in a torus ($M=\mathbb{S}^{1}\times\mathbb{S}^{1}$).
6.1 Long-period comets
The first example is the comet orbits data introduced in Section 1. These data are available from the JPL Small-Body Database (https://ssd.jpl.nasa.gov/tools/sbdb_query.html) by NASA. In this repository, orbits of objects in the solar system are registered measuring two angles called inclination, $i\in[0,\pi]$, and longitude of the ascending node, $\Omega\in[0,2\pi)$. Given an orbit with parameters $(i,\Omega)$, its normal unit vector is $v=(\sin(i)\sin(\Omega),-\sin(i)\cos(\Omega),\cos(i))^{\prime}\in\mathbb{S}^{2}$, where the direction of the vector depends on the rotation direction of the comet (see Jupp et al.,, 2003). We considered the normal unit vectors of all the long-period comets detected up to 30th of May 2022. This data set is known to have duplicate entries, so, similarly to García-Portugués et al., (2023), only comets with distinct $(i,\Omega)$ up to the second decimal position are considered, obtaining $n=612$ comets in total. The normal unit vector of these $n=612$ comet orbits are represented in Figure 1.
As explained in Section 1, the orbits of this kind of comets are expected to be uniformly distributed due to their formation process. Nevertheless, the observed sample have been shown to be not uniform (see García-Portugués et al.,, 2023), a situation that is suggested to be caused by an observational bias: astronomers search for new objects near the ecliptic more frequently than elsewhere on the celestial sphere, and therefore observations corresponding to comets located near the ecliptic plane are overrepresented within the sample. HDR estimation allows to check whether this hypothesis has empirical evidence or not, since, if true, the HDRs of the populations should be located near the north and south poles, as such normal vectors correspond to orbits near the ecliptic.
[…]
Figure 5 shows the border of the estimator $L_{n}(\hat{\lambda}_{\gamma,n})$ for $\gamma=0.8$ and $\gamma=0.9$. For $\gamma=0.8$ (respectively, for $\gamma=0.9$), the HDR estimator was computed with $r_{n}=0.3$ (respectively, for $r_{n}=0.275$). In both cases, $f_{n}$ was the kernel density estimator with von Mises-Fisher kernel (García–Portugués,, 2013) and concentration parameter chosen with Likelihood Cross-Validation. For $\gamma=0.8$, $L_{n}(\hat{\lambda}_{\gamma,n})$ consists of two connected components, each of them approximately centered near the north and south pole, which agrees with the observational bias hypothesis. For $\gamma=0.9$, $L_{n}(\hat{\lambda}_{\gamma,n})$ only has one connected component located near the south pole. Surprisingly, this seems to point out that one direction of rotation is slightly more common than the other. However, since normal vectors near the south pole correspond to orbits near the ecliptic plane, this behavior still provides empirical evidence supporting the observational bias theory.
We also compute the plug-in HDR estimator for directional data introduced by Saavedra-Nieves and Crujeiras, (2021) for comparison. Figure 6 shows the border of $f_{n}^{-1}\big{(}[\hat{\lambda}_{\gamma,n},+\infty]\big{)}$ for $\gamma=0.8$ and $\gamma=0.9$, where $f_{n}$ is the kernel density estimator with von Mises-Fisher kernel (García–Portugués,, 2013) and concentration parameter chosen with Likelihood Cross-Validation. Surprisingly, the shape of the plug-in HDR estimators is quite different from the ones shown in Figure 5. For $\gamma=0.8$, $f_{n}^{-1}\big{(}[\hat{\lambda}_{\gamma,n},+\infty]\big{)}$ has a further third component close to the equator, whereas for $\gamma=0.9$ $f_{n}^{-1}\big{(}[\hat{\lambda}_{\gamma,n},+\infty]\big{)}$ preserves a small second connected component near the north pole. These differences may be caused by the well-known fact that cross-validation techniques lead to undersmoothed estimators of the density function (see Hall et al.,, 1987). Of course, this behavior of the cross-validation bandwidth especially affects the plug-in estimator, while in the new proposal the incorporation of the shape conditions into the estimation method reduces the effect that the undersmoothed density estimator $f_{n}$ has on the final form of the HDR estimator $L_{n}(\hat{\lambda}_{\gamma,n})$.
[…]
6.2 Circadian gene expressions
Several organisms present biological processes that are synchronized with the solar day. For example, mammals have multiple of physiological parameters that fluctuate in an approximately 24 hours period, like blood pressure, heartbeat rate, and liver metabolism (see, for instance, Balsalobre et al.,, 1998). These biochemical oscillators are called circadian clocks, and they allow the organism to prepare to the usual activities that occur during a day.
There are a large variety of genes involved in circadian clocks. The expression of these genes is characterized by exhibiting a periodical behavior, with a time of maximum expression that repeats every 24 hours and receives the name of phase. This periodicity should be taken into account to correctly analyze circadian gene data. Thus, phases are often regarded as circular data, i.e., points of the unit circle $\mathbb{S}^{1}$. Indeed, circadian clocks are classical example of circular data, see Brown and Chow, (1976) and Batschelet, (1981).
The genetic circadian behavior has been shown to be tissue-specific. That is, the same gene may have a circadian expression in one tissue and not in another. Or, even when a gene shows a circadian behavior in both tissues, its phase may be different in every tissue. In order to explore this tissue-specific behavior, Liu et al., (2006) analyzed 48 common circadian genes present in the heart and liver of mice detected by Storch et al., (2002). Liu et al., (2006) studied the difference between the phases in both tissues and they discovered two groups of genes: one consisting of 30 genes with similar phases in heart and liver, and a smaller group of 10 genes with an average lag of 8 hours between the two.
The new HDR estimator allows to analyze both phases jointly, and therefore to perform a more detailed exploration of the data. Since both phases are points in the unit circle $\mathbb{S}^{1}$, the joint sample is a point cloud in the two-dimensional torus $\mathbb{T}^{2}=\mathbb{S}^{1}\times\mathbb{S}^{1}$. Figure 7 shows the heart and liver phases of all 48 genes and the border of the estimator $L_{n}(\hat{\lambda}_{\gamma,n})$ for three different values of $\gamma$: $\gamma=0.25$, $\gamma=0.5$, and $\gamma=0.75$. In all cases, $f_{n}$ was the kernel density estimator with von Mises kernel and concentration parameter chosen with maximum Likelihood Cross-Validation (Di Marzio et al.,, 2011). The value of $r_{n}$ was different for the three values $\gamma$: $r_{n}=0.6$ for $\gamma=0.25$, $r_{n}=0.45$ for $\gamma=0.5$, and $r_{n}=0.3$ for $\gamma=0.75$.
[…]
Figure 7 shows that $L_{n}(\hat{\lambda}_{\gamma,n})$ is relatively close to the diagonal for the three values of $\gamma$. This fact seems to point out that circadian genes commonly exhibit a similar behavior in both tissues. In addition, the HDR estimators reveal the presence of two different clusters within the data: one centered at $(0,0)$ and other around $(12,12)$. One group consists of morning genes, i.e., genes with both phases near 00, indicating that the expression of these genes peaks around the beginning of circadian day in both tissues. The second group is composed of night genes, with phases close to the start of circadian night in heart and liver.
[…]
For comparison, we also estimate the HDRs of the circadian gene data using the plug-in estimator for manifold data studied by Cholaquidis et al., (2022). Figure 8 shows the border of $f_{n}^{-1}\big{(}[\hat{\lambda}_{\gamma,n},+\infty]\big{)}$ for $\gamma=0.25$, $\gamma=0.5$, and $\gamma=0.75$, where $f_{n}$ is the kernel density estimator with von Mises kernel and concentration parameter chosen with Likelihood Cross-Validation (Di Marzio et al.,, 2011). The general conclusions drawn from Figure 7 still hold with the plug-in HDR estimator. The three estimators in Figure 8 are located near the diagonal, which indicates that most circadian genes have a similar behavior in heart and liver. The distinction between morning and night genes described above is also appreciable in Figure 8. Furthermore, Figure 8 shows the “undersmoothing” problematic of the plug-in estimator explained in Section 6.1. For example, for $\gamma=0.5$, $f_{n}^{-1}\big{(}[\hat{\lambda}_{\gamma,n},+\infty]\big{)}$ has a tiny connected component near $(12,6)$ that was not present in $L_{n}(\hat{\lambda}_{\gamma,n})$ (see Figure 7). This component has no sample points in it, which may indicate that it is a consequence of a undersmooth estimation of $f$ and not a proper component of the HDR of the population.
7 Discussion and future work
We have introduced a new estimation technique for HDRs in Riemannian manifolds. This new proposal is based on the HDR estimator for Euclidean data introduced by Walther, (1997) and adapts its ideas to any manifold satisfying very nonrestrictive conditions. The central result is Theorem 4.2, which proves the consistency of the proposed estimator and establishes its convergence rate. Theorem 4.2 ensures that the convergence of this estimator is uniform with respect to the level $\lambda$. This allows the application of the new estimator for recovering the smallest set that satisfies a specified probability content through a plug-in approach. This technique is explored in Section 5.1, where its consistency is proved. Additionally, Section 5.2 introduces an estimator for the corresponding level, and its convergence rate is derived. To highlight its flexibility, the estimator is then applied to analyze two real datasets, each from a different manifold.
Perspectives for new research are rich and diverse. The clearest line of future research is to propose an estimator for the radius $r$ in the manifold setting. A possible approach is to generalize the estimator for Euclidean data introduced by Rodríguez-Casal and Saavedra-Nieves, (2022), which is in turn based on the relation between $r$ and the gradient of $f$ given by Walther, (1997). Nevertheless, this relation between $r$ and $f$ remains unclear in the manifold setting. Another possible line to estimate $r$ is to rely on maximal spacing. Given a set $A$ and a sample $\mathcal{X}_{n}$, the maximal spacing of $A$ (with respect to $\mathcal{X}_{n}$) is the radius of the largest ball contained in $A$ that does not intersect $\mathcal{X}_{n}$, that is
$\mathrm{MS}(A,\mathcal{X}_{n})=\sup\big{\{}r\geq 0\colon\exists x\in A\text{ such that }B_{r}(x)\subset A\setminus\mathcal{X}_{n}\big{\}}.$
Aaron et al., (2017) use this concept to test the convexity of the support of the data. The maximal spacing of the convex hull of the data converges to zero if the support is convex, and to a positive constant if not. This suggests rejecting the null hypothesis of convexity when the maximal spacing of the convex hull takes large values. A similar idea might be applied to estimate $r$, choosing $r_{n}$ in a way that $\mathrm{MS}\big{(}L_{n}(\lambda),\mathcal{X}_{n}^{+}(\lambda)\big{)}$ is not too large.
Another interesting approach for estimating $r_{n}$ relies on the area of $L_{n}(\lambda)$. Notice that, if $r_{n}$ is too large of too small, the estimator $L_{n}(\lambda)$ will be the empty set for some $\lambda$. This fact seems to point out that choosing the value $r_{n}$ that maximizes the area of $L_{n}(\lambda)$ will provide consistent estimators for $r$. However, this approach will require to compute the area of $L_{n}(\lambda)$ for a grid of values of $r_{n}$ and, therefore, will be unfeasible in practice (or, at least, unpractical).
The analysis of the two data examples in Section 6 raises the question of how to visualize variability in this context, i.e., how to make inference on the HDRs of a population supported on a manifold. In the Euclidean setting, Mason and Polonik, (2009) show the asymptotic normality of the HDR plug-in estimators in terms of Lebesgue measure distance:
$\mathcal{d}_{\mathrm{Vol}_{\mathbb{R}^{d}}}\big{(}L(\lambda),f_{n}^{-1}[\lambda,+\infty)\big{)}=\mathrm{Vol}_{\mathbb{R}^{d}}\big{(}L(\lambda)\triangle f_{n}^{-1}[\lambda,+\infty)\big{)}$
where $A\triangle B=(A\setminus B)\cup(B\setminus A)$ for all $A,B\subset\mathbb{R}^{d}$. This distance, yet useful, does not allow to develop confidence regions for the HDRs of the population. In this sense, Chen et al., (2017) derive the asymptotic distribution of the Hausdorff distance between the HDR of level $\lambda$ and its plug-in estimator, $\mathcal{d}_{H}\big{(}L(\lambda),f_{n}^{-1}[\lambda,+\infty)\big{)}$, and develop a bootstrap method to establish a confidence region for the border of $L(\lambda)$. Mammen and Polonik, (2013) also propose a method based on the (Euclidean) plug-in HDR estimator to construct confidence regions for $L(\lambda)$ but relying on a entirely different approach that does not involve the Hausdorff distance. To the best of our knowledge, all the contributions in this line are restricted to the Euclidean context and the plug-in estimator, and this problem remains unexplored for other kinds of HDR estimators in a general manifold. Deriving the asymptotic distribution of $\mathcal{d}_{H}\big{(}L(\lambda),L_{n}(\lambda)\big{)}$ will not only provide a reliable technique to construct confidence regions for $L(\lambda)$, but will also allow to find a lower bound of the number of connected components of $L(\lambda)$, which is related with the number of clusters within the population.
The HDR estimator introduced in (3.6) can be an interesting preliminary tool for several tasks in data analysis. For example, as mentioned above, the different connected components of $L(\lambda)$ can be regarded as population clusters, and consequently, HDR estimation can be used for cluster analysis. This approach has been explored in the literature for euclidean data, see Cuevas et al., (2001), but, to the best of the authors’ knowledge, it has not been treated in the manifold context yet. In addition, the number of connected components of $L_{n}(\lambda)$ is an estimator of the number of clusters, which is a crucial initial parameter to set in many clustering techniques. HDR estimation is also useful for outlier detection. For small values of $\tau$, $L(\lambda_{\tau})$ approaches the substantial support of the population, and a new observation is classified as an outlier if lies outside of it. Similarly to cluster analysis, this application of HDR estimation has been investigated in the Euclidean context (see Devroye and Wise,, 1980 and Baíllo et al.,, 2001), but it has not been considered yet for data on manifolds. Comparison of two (or more) populations is another possible application, as one can distinguish density functions by measuring the Hausdorff distance between their HDRs. This approach allows for different weights to be assigned to different values of $\lambda$, thereby prioritizing differences in the central parts of the populations over the tails, or vice versa.
Appendix A Properties of Minkowski operations on manifolds
Matheron, (1975) studied the properties of the $\Psi_{r}$ mapping defined in equation (3.2) for subsets of the Euclidean space. For example, in the Euclidean setting $\Psi_{r}$ satisfies:
  • If $s\leq r$, then $\Psi_{r}(A)\subset\Psi_{s}(A)$.
  • If $A\subset A^{\prime}$, then $\Psi_{r}(A)\subset\Psi_{s}(A^{\prime})$.
  • If $s\leq r$, then $\Psi_{s}\circ\Psi_{r}=\Psi_{r}$.
In this appendix, we prove that the properties of monotonicity and idempotence are preserved in the manifold setting. First, some simple statements regarding the monotonicity of $\Psi_{r}$ are shown.
Proposition A.1.
Let $M$ be a Riemannian manifold, $A,A^{\prime}\subset M$ two arbitrary subsets, and $r>0$ a positive constant. Then:
  1. (a)
    $\Psi_{r}(A)\subset A$.
  2. (b)
    If $A\subset A^{\prime}$, then $\Psi_{r}(A)\subset\Psi_{r}(A^{\prime})$.
  3. (c)
    $\cup_{A\in\mathcal{F}}\Psi_{r}(A)\subset\Psi_{r}(\cup_{A\in\mathcal{F}}A)$, where $\mathcal{F}\subset\mathcal{P}(M)$.
  4. (d)
    $\Psi_{r}(A\oplus rB)=A\oplus rB$.
Proof.
The proof of this result is simple and is left to the reader. ∎
If $M$ is assumed to be a connected and complete manifold, the Hopf-Rinow Theorem (Lee,, 2018, Cor. 6.21) guarantees that each pair of points in $M$ can be joined with a length-minimizing geodesic. Specifically, given two points in $M$, $x,y\in M$, there exists a geodesic curve $\alpha:[a,b]\rightarrow M$ such that
[…]
Every admissible curve can be parameterized by arc length, so it will be always assumed that $\alpha:[0,\mathcal{d}(x,y)]\rightarrow M$ without loss of generality. Furthermore, since $\alpha$ minimizes the distance between $x$ and $y$, it is also minimizes the distance between any two points within its path. That is to say
$\mathcal{d}\big{(}\alpha(t),\alpha(s)\big{)}=\left|t-s\right|$
for all $0\leq t,s\leq\mathcal{d}(x,y)$.
This simple property guarantees that the $\Psi_{\bullet}$ mapping and Minkowski operations behave satisfactorily when two radii are used, as the following two propositions show.
Proposition A.2.
Let $M$ be a connected and complete Riemann manifold, $A\subset M$ and $0<s\leq r$. The following statements hold:
  1. (a)
    $B_{r}[x]=\Psi_{s}\big{(}B_{r}[x]\big{)}$ for all $x\in M$.
  2. (b)
    $\Psi_{r}(A)=\Psi_{s}\big{(}\Psi_{r}(A)\big{)}$.
  3. (c)
    $\Psi_{r}(A)\subset\Psi_{s}(A).$
  4. (d)
    If $A=\Psi_{r}(A)$, then $A=\Psi_{s}(A)$.
Proof.
Step by step.
  1. (a)
    Let $y\in B_{r}[x]$. Denote by $\alpha$ the length-minimizing geodesic joining $x$ with $y$. We distinguish two cases:
    • If $\mathcal{d}\big{(}x,y\big{)}\leq s$, then $y\in B_{s}[x]\subset B_{r}[x]$ and therefore $y\in\Psi_{s}\big{(}B_{r}[x]\big{)}$.
    • If $\mathcal{d}\big{(}x,y\big{)}>s$, the value $t=\mathcal{d}\big{(}x,y\big{)}-s$ verifies that $\mathcal{d}\big{(}x,\alpha(t)\big{)}=\mathcal{d}\big{(}x,y\big{)}-s$ and $\mathcal{d}\big{(}y,\alpha(t)\big{)}=s$. Furthermore, given any $z\in B_{s}\big{[}\alpha(t)\big{]}$:
      $\mathcal{d}\big{(}y,z\big{)}\leq\mathcal{d}\big{(}y,\alpha(t)\big{)}+\mathcal{d}\big{(}\alpha(t),z\big{)}\leq\mathcal{d}\big{(}x,y\big{)}-s+s=\mathcal{d}\big{(}x,y\big{)}\leq r.$
      So, $y\in B_{s}\big{[}\alpha(t)\big{]}\subset B_{r}[x]$ and that implies $y\in\Psi_{s}\big{(}B_{r}[x]\big{)}$.
  2. (b)
    From (a) and Propositions A.1a and A.1c, it follows the next chain of inclusions holds:
    $\Psi_{r}(A)=\left(\bigcup_{B_{r}[x]\subset A}B_{r}[x]\right)=\left(\bigcup_{B_{r}[x]\subset A}\Psi_{s}\big{(}B_{r}[x]\big{)}\right)\subset\\ \subset\Psi_{s}\left(\bigcup_{B_{r}[x]\subset A}B_{r}[x]\right)=\Psi_{s}\big{(}\Psi_{r}(A)\big{)}\subset\Psi_{r}(A).$
    Hence, the equality also holds.
  3. (c)
    It is straightforward from (b), $\Psi_{r}(A)\subset A$ and Proposition A.1b.
  4. (d)
    From (c) and Proposition A.1a, we deduce the next chain of inclusions
    $A=\Psi_{r}(A)\subset\Psi_{s}(A)\subset A,$
    so the equality holds.∎
Proposition A.3.
Let $M$ be a connected and complete Riemann manifold, $A\subset M$ and $r,s>0$. Then:
  1. (a)
    $(A\oplus rB)\oplus sB=A\oplus(r+s)B$.
  2. (b)
    $(A\ominus rB)\ominus sB=A\ominus(r+s)B$.
Proof.
Let $x\in(A\oplus rB)\oplus sB$. Then, there exists a point $y\in A\oplus rB$ verifying that $x\in B_{s}[y]$. In addition, there exists a point $z\in A$ satisfying that $y\in B_{r}[z]$. From the triangular inequality, one derives
$\mathcal{d}(x,z)\leq\mathcal{d}(x,y)+\mathcal{d}(y,z)\leq r+s.$
Therefore $x\in B_{r+s}[z]$ and then $x\in A\oplus(r+s)B$.
Now let $x\in A\oplus(r+s)B$. There is a point $y\in A$ such that $x\in B_{r+s}[y]$. We distinguish two cases:
  • If $\mathcal{d}(x,y)\leq r$, then $x\in A\oplus rB\subset(A\oplus rB)\oplus sB$.
  • Assume that $\mathcal{d}(x,y)>r$. Let $\alpha:\big{[}0,\mathcal{d}(x,y)\big{]}\rightarrow M$ be the length-minimizing geodesic joining $y$ with $x$, that is, $\alpha$ verifies
[…]
    Thus, $\alpha(r)\in B_{r}[y]$ and then $\alpha(r)\in A\oplus rB$. Moreover,
    $d\big{(}\alpha(r),x\big{)}=d\big{(}x,y\big{)}-r\leq s$
    Therefore $x\in B_{s}\big{[}\alpha(r)\big{]}$ and so $x\in(A\oplus rB)\oplus sB$.
Hence, (a) holds. Part (b) is a direct consequence of (a):
$(A\ominus rB)\ominus sB=\big{[}(A^{c}\oplus rB)\oplus sB\big{]}^{c}=\big{[}A^{c}\oplus(r+s)B\big{]}^{c}=A\ominus(r+s)B.\qed$
For closed subsets, Minkowski operations are closely related with the distance function.
Proposition A.4.
Let $M$ be a connected and complete manifold, $A\subset M$ a closed subset, and $r>0$. The following statements hold.
  1. (a)
    $A\oplus rB=\{x\in M\colon\mathcal{d}(x,A)\leq r\}.$
  2. (b)
    $A\ominus rB=\{x\in M\colon\mathcal{d}(x,A^{c})\geq r\}.$
  3. (c)
    $A\oplus rB$, $A\ominus rB$ and $\Psi_{r}(A)$ are closed subsets of $M$.
Proof.
Let $x\in A\oplus rB$. Then, there exists $a\in A$ such that $x\in B_{r}[a]$ and so $\mathcal{d}(x,A)\leq r$. Reciprocally, let $x\in M$ such that $\mathcal{d}(x,A)\leq r$. Then, $B_{r+1/n}[x]\cap A\neq\emptyset$ holds for all $n\in\mathbb{N}$. From Cantor’s intersection theorem, it follows that
$B_{r}[x]\cap A=\bigcap_{n\in N}B_{r+1/n}[x]\cap A\neq\emptyset.$
So, there exists a point $a\in B_{r}[x]\cap A$. Therefore, $x\in B_{r}[a]$ and $x\in A\oplus rB$. This concludes the proof of (a).
The proof of (b) is similar. Let $x\in A\ominus rB$. Then $B_{r}[x]\subset A$. Therefore, $B_{r}[x]\cap A^{c}=\emptyset$ and $\mathcal{d}(x,A^{c})\geq r$. Assume then that $x\in M$ verifies $\mathcal{d}(x,A^{c})\geq r$. Consequently, $B_{r}(x)\cap A^{c}=\emptyset$, that is to say, $B_{r}(x)\subset A$. Since $A$ is closed, it follows that $\overline{B_{r}(x)}=B_{r}[x]\subset A$ and then $x\in A\ominus rB$.
Part (c) is a direct consequence of (a) and (b). ∎
Remark A.1.
Note that, since $(A\oplus rB)^{c}=A^{c}\ominus rB$ for every $A\subset M$, the previous proposition guarantees that Minkowski operations and $\Psi_{r}$ also preserve open subsets.
In fact, distance and Minkowski operations have a stronger connection than Proposition A.4 shows. For example, given a closed subset $A$, the distances $\mathcal{d}(x,A^{c})$ and $\mathcal{d}(x,A\ominus rB)$ are linked by a simple and useful equation.
Proposition A.5.
Let $M$ be a connected and complete manifold, $A\subset M$ be a closed set and $r,\delta>0$. Assume that $\Psi_{r}(A)=A$ and $\Psi_{\delta}(A^{c})=A^{c}$. Then
$\mathcal{d}(x,A^{c})+\mathcal{d}(x,A\ominus rB)=r,$
for all $x\in A\setminus\big{[}A\ominus rB\big{]}$.
Proof.
Let $x$ be an arbitrary point in $A\setminus\big{[}A\ominus rB\big{]}$, and let $p\in\partial A$ such that $\mathcal{d}(x,p)=d(x,A^{c})$. First, we need to show that there exists a point $y\in A^{c}$ verifying $p\in B_{\delta}[y]$ and $B_{\delta}(y)\subset A^{c}$.
Since $p\in\partial A$, there exists a sequence of points of $A^{c}$ satisfying $p_{n}\rightarrow p$. Keeping in mind that $A^{c}=\Psi_{\delta}(A^{c})$, for every $p_{n}$ there exists a center $y_{n}$ such that $p_{n}\in B_{\delta}[y_{n}]\subset A^{c}$. The sequence $p_{n}$ converges to $p$, so there is a constant $R>0$ verifying $\mathcal{d}(p,p_{n})<R$ for all $n\in\mathbb{N}$. Thus, $\mathcal{d}(p,y_{n})\leq\mathcal{d}(p,p_{n})+\mathcal{d}(p_{n},y_{n})<R+\delta$ holds for all $n\in\mathbb{N}$ and the sequence $y_{n}$ is bounded. $M$ is a complete metric space, and so $y_{n}$ has a convergent subsequence. For simplicity, the convergent subsequence of centers and the corresponding $p_{n}$’s subsequence are named as the original ones, $\{y_{n}\}_{n\in\mathbb{N}}$ and $\{p_{n}\}_{n\in\mathbb{N}}$, $y_{n}\rightarrow y$, $p_{n}\rightarrow p$. Then $\mathcal{d}(y,p)=\lim_{n\rightarrow+\infty}\mathcal{d}(y_{n},p_{n})\leq\delta$ and $\mathcal{d}(y,A)=\lim_{n\rightarrow+\infty}\mathcal{d}(y_{n},A)\geq\delta$. From the previous statements it follows that $p\in B_{\delta}[y]$ and $B_{\delta}(y)\subset A^{c}$, and since $p\in\partial A\subset A$, $\mathcal{d}(p,y)=\delta$.
On the other hand, since $p\in A$ and $\Psi_{r}(A)=A$, there exists a point $q\in A$ such that $p\in B_{r}[q]\subset A$. Furthermore, $p\in\partial A$, so $d(p,q)=r$.
$y$
$p$
$x$
$q$
$\beta_{x}$
$\beta_{q}$
$\alpha$
Figure 9: The four points $p,x,q$ and $y$, together with the length-minimizing segments $\alpha$, $\beta_{x}$ and $\beta_{q}$.
The next step is to prove that $p,x,q$ and $y$ lie on the same length-minimizing geodesic segment. This can be derived from uniqueness of geodesics. Let $\alpha$ be a geodesic segment joining $y$ and $p$; and $\beta_{x}$ a geodesic segment joining $p$ and $x$ (see Figure 9 for an illustration). Since $p\in B_{\delta}[y]\cap B_{\mathcal{d}(x,p)}[x]$ and $B_{\delta}(y)\cap B_{\mathcal{d}(x,p)}(x)\subset A^{c}\cap A=\emptyset$, one derives that
$\mathcal{d}(x,y)=\mathcal{d}(x,p)+\delta=\mathcal{d}(x,p)+\mathcal{d}(p,y)$
and therefore, the curve
$\gamma_{1}(t)=\left\{\begin{array}[]{ll}\alpha(t),&t\in[0,\delta]\\ \beta_{x}(t-\delta),&t\in[\delta,\delta+\mathcal{d}(x,p)]\end{array}\right.$
verifies that
[…]
So, $\gamma_{1}$ is a length-minimizing curve joining $y$ and $x$, and then $\gamma_{1}$ is a geodesic segment.
Let now $\beta_{q}$ be a geodesic segment from $p$ to $q$ (see Figure 9). Following a similar argument, one shows that
$\gamma_{2}(t)=\left\{\begin{array}[]{ll}\alpha(t),&t\in[0,\delta]\\ \beta_{q}(t-\delta),&t\in[\delta,\delta+\mathcal{d}(q,p)]\end{array}\right.$
is a length-minimizing curve and a geodesic segment from $y$ to $q$. But $\gamma_{1}(t)=\gamma_{2}(t)$ for all $t\in[0,\delta]$, and since geodesics are unique, it follows that $\gamma_{1}=\gamma_{2}$ in their common support. Particularly,
$\beta_{q}\big{(}\mathcal{d}(x,p)\big{)}=\gamma_{2}\big{(}\delta+\mathcal{d}(x,p)\big{)}=\gamma_{1}\big{(}\delta+\mathcal{d}(x,p)\big{)}=\beta_{x}\big{(}\mathcal{d}(x,p)\big{)}=x;$
so the geodesic segment $\beta_{q}$ joins the points $p,x$ and $q$.
Finally, one derives that
$r\geq\mathcal{d}(p,q)=\mathcal{d}(p,x)+\mathcal{d}(x,q)=\mathcal{d}(x,A^{c})+\mathcal{d}(x,q)\geq\\ \geq\mathcal{d}(x,A^{c})+\mathcal{d}(x,A\ominus rB)\geq\mathcal{d}(A^{c},A\ominus rB)\geq r,$
where the last inequality follows from $\Psi_{r}(A)=A$ and Proposition A.4a. Consequently, $\mathcal{d}(x,A\ominus rB)+\mathcal{d}(x,A^{c})=r$ holds. ∎
This section ends with three lemmas exploring how $\Psi_{r}$, Minkowski operations, and set difference are related. The first one yields that, if a set $A\subset M$ verifies $\Psi_{r+s}(A)=A$, then $\Psi_{s}(A\ominus rB)=A\ominus rB$.
Lemma A.1.
Let $M$ be a connected and complete Riemannian manifold. Let $A\subset M$ be a closed set and $s,r,\delta>0$ some positive constants. Assume that $\Psi_{r+s}(A)=A$ and $\Psi_{\delta}(A^{c})=A^{c}$. Then $\Psi_{s}(A\ominus rB)=A\ominus rB$.
Proof.
The content $\Psi_{s}(A\ominus rB)=A\ominus rB$ is straightforward by Proposition A.1a. It follows from the definition of $\Psi_{s}$ that $A\ominus(r+s)B\subset\Psi_{s}(A\ominus rB)$. Let $x\in A\setminus\big{[}A\ominus(r+s)B\big{]}$. Proposition A.5 guarantees that
$\mathcal{d}(x,A^{c})+\mathcal{d}\big{(}x,A\ominus(r+s)B\big{)}=r+s.$ (A.1)
Therefore, for all $x\in\big{[}A\ominus rB\big{]}\setminus\big{[}A\ominus(r+s)B\big{]}$, it holds that
$\mathcal{d}\big{(}x,A\ominus(r+s)B\big{)}=r+s-\mathcal{d}(x,A^{c})\leq s;$
where the last inequality is derived from Proposition A.4b. Then, taking into account that Proposition A.3b ensures that $A\ominus(r+s)B=(A\ominus rB)\ominus sB$, Proposition A.4a yields that
$x\in\big{[}(A\ominus rB)\ominus sB\big{]}\oplus sB=\Psi_{s}(A\ominus rB).\qed$
Remark A.2.
It may not be obvious at first, but the hypothesis that $\Psi_{\delta}(A^{c})=A^{c}$ for some $\delta>0$ is necessary for Lemma A.1 to hold. For example, consider the closed set $A\subset\mathbb{R}^{2}$ represented in Figure 10(a) and defined by
$A=B_{\sqrt{2}}\big{[}(-1,0)\big{]}\cup B_{\sqrt{2}}\big{[}(1,0)\big{]}.$
This set verifies that $\Psi_{\sqrt{2}}(A)=\Psi_{1+(\sqrt{2}-1)}(A)=A$, but $\Psi_{\delta}(A^{c})=A^{c}$ does not hold for any $\delta>0$.
The set $A\ominus 1B$ is represented in Figure 10(b). For this set, one can easily show that $\Psi_{\sqrt{2}-1}(A\ominus 1B)\neq A\ominus 1B$. In fact, $(0,0)\in A\ominus 1B$ and $(0,0)\not\in\big{(}A\ominus\sqrt{2}B\big{)}\oplus(\sqrt{2}-1)B$. Therefore
$\Psi_{\sqrt{2}-1}(A\ominus 1B)=\big{(}A\ominus\sqrt{2}B\big{)}\oplus(\sqrt{2}-1)B\neq A\ominus 1B.$
(a) $A$
(b) $A\ominus 1B$
Figure 10: The sets $A$ and $A\ominus 1B$ from Remark A.2.
The next lemma ensures that $\Psi_{s}\big{(}[A\oplus rB]\setminus A\big{)}=[A\oplus rB]\setminus A$ when $s$ is small enough.
Lemma A.2.
Let $M$ be a connected and complete manifold, $A\subset M$ a closed set and $r,s>0$ two positive constants such that $s\leq r/3$. If $\Psi_{s}(A^{c})=A^{c}$ then
$\Psi_{s}\big{(}[A\oplus rB]\setminus A\big{)}=[A\oplus rB]\setminus A.$
Proof.
If $[A\oplus rB]\setminus A=\emptyset$, the result is obvious. Contrarily, assume that $[A\oplus rB]\setminus A\neq\emptyset$. Let $x\in[A\oplus rB]\setminus A=[A\oplus rB]\cap A^{c}$. Since $A$ is a closed set, then $0<\mathcal{d}(x,A)\leq r$. We distinguish three cases:
  1. Case 1:
    $30<\mathcal{d}(x,A)\leq r/$ 3. Since $\Psi_{s}(A^{c})=A^{c}$ there exists a point $y\in A^{c}$ verifying $x\in B_{s}[y]\subset A^{c}$. Furthermore, every $z\in B_{s}[y]$ satisfies
    $\mathcal{d}(z,A)\leq\mathcal{d}(z,y)+\mathcal{d}(y,x)+\mathcal{d}(x,A)\leq r;$
    so $B_{s}[y]\subset[A\oplus rB]\setminus A$. Consequently, $x\in\Psi_{s}\big{(}[A\oplus rB]\setminus A\big{)}$.
  2. Case 2:
    $r/3<\mathcal{d}(x,A)\leq 2r/3$. Then, $B_{s}[x]\subset[A\oplus rB]\setminus A$ and it follows that $x\in\Psi_{s}\big{(}[A\oplus rB]\setminus A\big{)}$.
  3. Case 3:
    $2r/3<\mathcal{d}(x,A)\leq r$. Therefore, Propositions A.1d and A.2d guarantee that $A\oplus rB=\Psi_{s}(A\oplus rB)$. So, there exists a point $y\in A\oplus rB$ such that $x\in B_{s}[y]\subset A\oplus rB$. In addition, every $z\in B_{s}[y]$ verifies
    $\mathcal{d}(x,A)\leq\mathcal{d}(x,y)+\mathcal{d}(y,z)+\mathcal{d}(z,A)\leq 2s+\mathcal{d}(z,A)\Rightarrow\\ \Rightarrow\mathcal{d}(z,A)\geq\mathcal{d}(x,A)-2s>\frac{2}{3}r-2s>0.$
    It follows that $B_{s}[y]\subset[A\oplus rB]\setminus A$, and so $x\in\Psi_{s}\big{(}[A\oplus rB]\setminus A\big{)}$. ∎
Finally, the last lemma states the relation between Minkowski operations and set difference.
Lemma A.3.
Let $M$ be a connected and complete Riemannian manifold. Let $A\subset M$ be a closed set and $r,s>0$. Then
$\big{[}(A\oplus rB)\setminus A\big{]}\oplus sB=\big{[}A\oplus(r+s)B\big{]}\setminus[A\ominus sB].$
Proof.
The first inclusion can be derived directly:
$\big{[}(A\oplus rB)\setminus A\big{]}\oplus sB=$ $\big{[}(A\oplus rB)\cap A^{c}\big{]}\oplus sB$
$\subset$ $\big{[}A\oplus(r+s)B\big{]}\cap\big{[}A^{c}\oplus sB\big{]}$
$=$ $\big{[}A\oplus(r+s)B\big{]}\cap\big{[}A\ominus sB\big{]}^{c}$
$=$ $\big{[}A\oplus(r+s)B\big{]}\setminus\big{[}A\ominus sB\big{]},$
where we have applied Proposition A.3a. Let $x\in\big{[}A\oplus(r+s)B\big{]}\cap\big{[}A^{c}\oplus sB\big{]}$. We distinguish three cases:
  1. Case 1:
    $x\in A$. Since $x\in A^{c}\oplus sB$, there exists a point $y\in A^{c}$ such that $x\in B_{s}[y]$. If $y\in A\oplus rB$ the inclusion holds. Assume that $y\notin A\oplus rB$. Let $\alpha$ be a geodesic segment joining $x$ with $y$, $\alpha:[0,\mathcal{d}(x,y)]\rightarrow M$. Define
    $t_{1}=\max\Big{\{}t\in\big{[}0,\mathcal{d}(x,y)\big{]}\colon\alpha(t)\in A\Big{\}}.$
    By definition $\alpha(t_{1})\in A$. Since $y\notin A\oplus rB$, Proposition A.4a ensures that
    $r<\mathcal{d}\big{(}\alpha(t_{1}),y\big{)}=\mathcal{d}(x,y)-t_{1},$
    or equivalently $r+t_{1}<\mathcal{d}(x,y)$. Then, the point $\alpha(t_{1}+r)$ verifies $\alpha(t_{1}+r)\in(A\oplus rB)\setminus A$ and $\mathcal{d}\big{(}x,\alpha(t_{1}+r)\big{)}<\mathcal{d}(x,y)\leq s$, and so $x\in\big{[}(A\oplus rB)\setminus A\big{]}\oplus sB$.
  2. Case 2:
    $x\notin A$ and $x\in A\oplus rB$. It is straightforward that $x\in\big{[}(A\oplus rB)\setminus A\big{]}\oplus sB$.
  3. Case 3:
    $x\notin A\oplus rB$. Since $x\in(A\oplus rB)\oplus sB$ by Proposition A.3a, there exists a point $y\in A\oplus rB$ such that $x\in B_{s}[y]$. If $y\in A^{c}$ the result holds. Suppose that $y\in A$ and let $\alpha$ be a geodesic segment joining $x$ with $y$, $\alpha:\big{[}0,\mathcal{d}(x,y)\big{]}\rightarrow M$. Define
    $t_{2}=\min\Big{\{}t\in\big{[}0,\mathcal{d}(x,y)\big{]}\colon\alpha(t)\in A\Big{\}}.$
    By definition $\alpha(t_{2})\in A$. Since $x\notin A\oplus rB$, Proposition A.4a shows that
    $r<\mathcal{d}\big{(}x,\alpha(t_{2})\big{)}=t_{2}.$
    Then, $\alpha(t_{2}-r)\in(A\oplus rB)\setminus A$ and
    $\mathcal{d}\big{(}x,\alpha(t_{2}-r)\big{)}<\mathcal{d}(x,y)\leq s.$
    So, $x\in\big{[}(A\oplus rB)\setminus A\big{]}\oplus sB$. ∎
Appendix B Consistency of $L_{n}(\lambda)$ and $\lambda_{\gamma,n}$
This appendix contains all the proofs related with the consistency of the estimators $L_{n}(\lambda)$ and $\hat{\lambda}_{\gamma,n}$ and its convergence rates. This appendix is organized in two sections. Section B.1 is fully devoted to the proof of Theorem 4.2. Finally, Section B.2 contains the proofs of Theorems 5.1 and 5.2.
B.1 Proof of Theorem 4.2
This section is fully devoted to the proof of Theorem 4.2. The proof of this results relies on several auxiliary results that we must introduce. The basis of the HDR estimator $L_{n}(\lambda)$ defined in (3.6) consists in using $\mathcal{X}^{+}_{n}(\lambda)\cap\big{(}\mathcal{X}^{-}_{n}(\lambda)\oplus r_{n}B\big{)}^{c}$ as a discretization of the set $L(\lambda)\ominus r_{n}B$, and then recovering $L(\lambda)$ by inflating this discretization. Lemma B.1 below studies the consistency of the last step of this procedure. If $\mathcal{X}_{n}$ is a i.i.d. sample from a given density, will $(\mathcal{X}_{n}\cap A)\oplus rB$ eventually recover the set $A\oplus rB$? The answer is affirmative, assuming some general conditions on the set $A$.
Lemma B.1.
Let $M$ be a Riemannian manifold under Assumptions M and $f:M\rightarrow\mathbb{R}$ a density function. Let $\mathcal{X}_{n}=\{X_{1},\ldots,X_{n}\}$ be a i.i.d. sample from the density function $f$.
Let $A\subset M$ a bounded set such that $f(x)\geq b>0$ for all $x\in A$. Let $r,\delta,\varepsilon>0$ verifying $\varepsilon<\delta\leq r$.
  1. (a)
    If $\Psi_{\delta}(A)=A$, then
    $\mathbb{P}\big{[}A\oplus(r-\varepsilon)B\not\subset(\mathcal{X}_{n}\cap A)\oplus rB\big{]}\leq D\left(\frac{\varepsilon}{2},A\oplus rB\right)\times\\ \times\exp\left[-nab\min{\left\{r-\frac{\varepsilon}{2},\delta,\rho\right\}}^{(d-1)/2}\min\left\{\frac{\varepsilon}{2},\rho\right\}^{(d+1)/2}\right];$
    where $\rho$ and $a$ are the constants in Assumption (M2) and $D(\cdot,\cdot)$ is the functional defined in Assumption (M3).
  2. (b)
    The following inequality holds:
    $\mathbb{P}\Big{[}S\oplus(r-\varepsilon)B\not\subset(\mathcal{X}_{n}\cap S)\oplus rB\text{ for some }S\subset A\text{ such that }\Psi_{\delta}(S)=S\Big{]}\\ \leq D\left(\frac{\varepsilon}{4},A\oplus rB\right)^{2}\exp\left[-nab\min{\left\{\frac{\delta}{2},\rho\right\}}^{(d-1)/2}\min\left\{\frac{\varepsilon}{4},\rho\right\}^{(d+1)/2}\right];$
    where $\rho$ and $a$ are the constants in Assumption (M2) and $D(\cdot,\cdot)$ is the functional defined in Assumption (M3).
Proof.
Let $T\subset A\oplus(r-\varepsilon)B$ be a subset such that
$\mathcal{d}(x,y)>\frac{\varepsilon}{2},\quad\forall x,y\in T,x\neq y$
with maximal cardinality. This implies that $A\oplus(r-\varepsilon)B\subset T\oplus\varepsilon/2B$. Thus, given any subset $S\subset M$, it follows that
$T\subset S\Rightarrow A\oplus(r-\varepsilon)B\subset S\oplus\varepsilon/2B.$
Then, keeping in mind that $\big{[}(\mathcal{X}_{n}\cap A)\oplus(r-\varepsilon/2)B\big{]}\oplus\varepsilon/2B=(\mathcal{X}_{n}\cap A)\oplus rB$, one deduces
$\mathbb{P}\big{[}A\oplus(r-\varepsilon)B\not\subset$ $(\mathcal{X}_{n}\cap A)\oplus rB\big{]}$
$\leq$ $\mathbb{P}\big{[}T\not\subset(\mathcal{X}_{n}\cap A)\oplus(r-\varepsilon/2)B\big{]}$
$\leq$ $\sum_{t\in T}\mathbb{P}\big{[}t\notin(\mathcal{X}_{n}\cap A)\oplus(r-\varepsilon/2)B\big{]}$
$\leq$ $D\big{(}\varepsilon/2,A\oplus rB\big{)}\sup_{t\in A\oplus(r-\varepsilon)B}\mathbb{P}\big{(}B_{r-\varepsilon/2}[t]\cap\mathcal{X}_{n}\cap A=\emptyset\big{)};$
where the last inequality holds because $T\subset A\oplus(r-\varepsilon)B$. Then, showing that
$\sup_{t\in A\oplus(r-\varepsilon)B}\mathbb{P}\big{(}B_{r-\varepsilon/2}[t]\cap\mathcal{X}_{n}\cap A=\emptyset\big{)}\\ \leq\exp\left[-nab\min{\left\{r-\frac{\varepsilon}{2},\delta,\rho\right\}}^{(d-1)/2}\min\left\{\frac{\varepsilon}{2},\rho\right\}^{(d+1)/2}\right]$ (B.1)
will complete the proof of part (a).
Fix a point $t\in A\oplus(r-\varepsilon)B$. It satisfies that
$\mathbb{P}\big{(}B_{r-\varepsilon/2}[t]\cap\mathcal{X}_{n}\cap A=\emptyset\big{)}=\Big{[}1-\mathbb{P}\big{(}X_{1}\in A\cap B_{r-\varepsilon/2}[t]\big{)}\Big{]}^{n}.$
Since $t\in A\oplus(r-\varepsilon)B$, there exists a point $x\in A$ such that $\mathcal{d}(x,t)\leq r-\varepsilon$. Furthermore, $\Psi_{\delta}(A)=A$, so there exists another point $y\in A$ verifying $x\in B_{\delta}[y]\subset A$. Therefore, $\mathcal{d}(t,y)\leq\delta+r-\varepsilon$ and
$\mathbb{P}\big{(}B_{r-\varepsilon/2}[t]\cap\mathcal{X}_{n}\cap A=\emptyset\big{)}\leq$ $\Big{[}1-\mathbb{P}\big{(}X_{1}\in B_{\delta}[y]\cap B_{r-\varepsilon/2}[t]\big{)}\Big{]}^{n}$
$\leq$ $\exp\Big{[}-n\mathbb{P}\big{(}X_{1}\in B_{\delta}[y]\cap B_{r-\varepsilon/2}[t]\big{)}\Big{]}$
$\leq$ $\exp\Big{[}-nb\mathrm{Vol}_{M}\big{(}B_{\delta}[y]\cap B_{r-\varepsilon/2}[t]\big{)}\Big{]}.$
Following Assumption (M2), one deduces
$\mathbb{P}\big{(}B_{r-\varepsilon/2}[t]\cap\mathcal{X}_{n}\cap A=\emptyset\big{)}\\ \leq\exp\left[-nab\min{\left\{r-\frac{\varepsilon}{2},\delta,\rho\right\}}^{(d-1)/2}\min\left\{\frac{\varepsilon}{2},\rho\right\}^{(d+1)/2}\right].$
This inequality holds for all $t\in A\oplus(r-\varepsilon)B$, so (B.1) holds too, concluding the proof of part (a).
Denote $\mathcal{S}_{\delta}(A)=\{S\subset A\colon\Psi_{\delta}(S)=S\}$. To prove part (b), the aim is to show:
$\mathbb{P}\left[\bigcup_{S\in\mathcal{S}_{\delta}(A)}\big{\{}S\oplus(r-\varepsilon)B\not\subset(\mathcal{X}_{n}\cap S)\oplus rB\big{\}}\right]\leq D\left(\frac{\varepsilon}{4},A\oplus rB\right)^{2}\times\\ \times\exp\left[-nab\min{\left\{\frac{\delta}{2},\rho\right\}}^{(d-1)/2}\min\left\{\frac{\varepsilon}{4},\rho\right\}^{(d+1)/2}\right].$
The key fact is that one does not have to check the condition
$S\oplus(r-\varepsilon)B\not\subset(\mathcal{X}_{n}\cap S)\oplus rB$
for all $S\in\mathcal{S}_{\delta}(A)$. It is enough to check it for a finite number of sets in $\mathcal{S}_{\delta/2}(A)$.
Let $T\subset A\ominus\delta B$ such that
$\mathcal{d}(x,y)>\frac{\varepsilon}{4},\quad\forall x,y\in T,x\neq y$
with maximal cardinality. Consider the following two statements:
  1. (S)
    It exists a $S\in\mathcal{S}_{\delta}(A)$ such that $S\oplus(r-\varepsilon)B\not\subset(\mathcal{X}_{n}\cap S)\oplus rB$.
  2. (T)
    It exists a $t\in T$ such that $B_{\delta-\varepsilon/4}[t]\oplus(r-\varepsilon/2)B\not\subset\big{(}\mathcal{X}_{n}\cap B_{\delta-\varepsilon/4}[t]\big{)}\oplus rB$.
We are going to show that (S) implies (T).
Suppose that there exists a set $S\in\mathcal{S}_{\delta}(A)$ such that
$S\oplus(r-\varepsilon)B\not\subset(\mathcal{X}_{n}\cap S)\oplus rB.$
Hence, there exists a point $x\in S\oplus(r-\varepsilon)B$ verifying $x\not\in(\mathcal{X}_{n}\cap S)\oplus rB$. From $S\in\mathcal{S}_{\delta}(A)$ and Proposition A.3a, it follows that
$x\in S\oplus(r-\varepsilon)B=\big{[}(S\ominus\delta B)\oplus\delta B\big{]}\oplus(r-\varepsilon)B=(S\ominus\delta B)\oplus(r+\delta-\varepsilon)B.$
So, there exists a point $s\in(S\ominus\delta B)$ such that $x\in B_{r+\delta-\varepsilon}[s]$. The definition of $T$ yields that $T\oplus\varepsilon/4B\supset A\ominus\delta B\supset S\ominus\delta B$. Therefore, there exists $t\in T$ satisfying $\mathcal{d}(s,t)\leq\varepsilon/4$. We will show that this point $t$ verifies the required condition in (T).
On the one hand:
$\mathcal{d}(x,t)\leq\mathcal{d}(x,s)+\mathcal{d}(s,t)\leq r+\delta-\frac{3}{4}\varepsilon.$
Consequently,
$x\in B_{r+\delta-3\varepsilon/4}[t]=B_{\delta-\varepsilon/4}[t]\oplus(r-\varepsilon/2)B.$ (B.2)
On the other hand, $\mathcal{d}(s,t)\leq\varepsilon/4$ and the triangular inequality guarantee that $B_{\delta-\varepsilon/4}[t]\subset B_{\delta}[s]\subset S$ where the last content follows from $s\in S\ominus\delta B$. Hence,
$\big{(}\mathcal{X}_{n}\cap B_{\delta-\varepsilon/4}[t]\big{)}\oplus rB\subset(\mathcal{X}_{n}\cap S)\oplus rB.$
and so
$x\not\in\big{(}\mathcal{X}_{n}\cap B_{\delta-\varepsilon/4}[t]\big{)}\oplus rB.$ (B.3)
Finally, from (B.2) and (B.3) one deduces
$B_{\delta-\varepsilon/4}[t]\oplus(r-\varepsilon/2)B\not\subset\big{(}\mathcal{X}_{n}\cap B_{\delta-\varepsilon/4}[t]\big{)}\oplus rB$
and therefore (S) implies (T).
Taking this into account, it follows that
$\mathbb{P}\Bigg{[}\bigcup_{S\in\mathcal{S}_{\delta}(A)}\big{\{}S\oplus(r-\varepsilon)B\not\subset(\mathcal{X}_{n}\cap S)\oplus rB\big{\}}\Bigg{]}\\ \leq\mathbb{P}\left[\bigcup_{t\in T}\big{\{}B_{\delta-\varepsilon/4}[t]\oplus(r-\varepsilon/2)B\not\subset\big{(}\mathcal{X}_{n}\cap B_{\delta-\varepsilon/4}[t]\big{)}\oplus rB\big{\}}\right]\\ \leq\sum_{t\in T}\mathbb{P}\Big{[}B_{\delta-\varepsilon/4}[t]\oplus(r-\varepsilon/2)B\not\subset\big{(}\mathcal{X}_{n}\cap B_{\delta-\varepsilon/4}[t]\big{)}\oplus rB\Big{]}\\ \leq D\left(\frac{\varepsilon}{4},A\oplus rB\right)\sup_{t\in A\ominus\delta B}\mathbb{P}\Big{[}B_{\delta-\varepsilon/4}[t]\oplus(r-\varepsilon/2)B\not\subset\big{(}\mathcal{X}_{n}\cap B_{\delta-\varepsilon/4}[t]\big{)}\oplus rB\Big{]},$
where the last inequality follows from $A\ominus\delta B\subset A\oplus rB$ and $T\subset A\ominus\delta B$. So, it only remains to show that
$\sup_{t\in A\ominus\delta B}\mathbb{P}\Big{[}B_{\delta-\varepsilon/4}[t]\oplus(r-\varepsilon/2)B\not\subset\big{(}\mathcal{X}_{n}\cap B_{\delta-\varepsilon/4}[t]\big{)}\oplus rB\Big{]}\\ \leq D\left(\frac{\varepsilon}{4},A\oplus rB\right)\exp\left[-nab\min{\left\{\frac{\delta}{2},\rho\right\}}^{(d-1)/2}\min\left\{\frac{\varepsilon}{4},\rho\right\}^{(d+1)/2}\right]$ (B.4)
to conclude the proof of part (b).
Let $t$ be an arbitrary point in $A\ominus\delta B$. One gets that:
  • $B_{\delta-\varepsilon/4}[t]\subset B_{\delta}[t]\subset A$, then $f(x)\geq b>0$ for all $x\in B_{\delta-\varepsilon/4}[t]$.
  • Since $\varepsilon<\delta$, then $\delta/2<\delta-\varepsilon/4$ and Proposition A.2a ensures that $\Psi_{\delta/2}\big{(}B_{\delta-\varepsilon/4}[t]\big{)}=B_{\delta-\varepsilon/4}[t]$.
Hence, part (a) guarantees that:
$\mathbb{P}\Big{[}B_{\delta-\varepsilon/4}[t]\oplus(r-\varepsilon/2)B\not\subset\big{(}\mathcal{X}_{n}\cap B_{\delta-\varepsilon/4}[t]\big{)}\oplus rB\Big{]}\\ \leq D\left(\frac{\varepsilon}{4},B_{\delta-\varepsilon/4}[t]\oplus rB\right)\times\\ \times\exp\left[-nab\min{\left\{r-\frac{\varepsilon}{4},\frac{\delta}{2},\rho\right\}}^{(d-1)/2}\min\left\{\frac{\varepsilon}{4},\rho\right\}^{(d+1)/2}\right]\\ \leq D\left(\frac{\varepsilon}{4},A\oplus rB\right)\exp\left[-nab\min{\left\{\frac{\delta}{2},\rho\right\}}^{(d-1)/2}\min\left\{\frac{\varepsilon}{4},\rho\right\}^{(d+1)/2}\right].$
Since this bound holds for every $t\in A\ominus\delta B$, (B.4) also holds, finishing the proof. ∎
Since $(\mathcal{X}_{n}\cap A)\oplus rB\subset A\oplus rB$, one deduces that
$\mathcal{d}_{H}\big{(}(\mathcal{X}_{n}\cap A)\oplus rB,A\oplus rB\big{)}>\varepsilon$ $\Rightarrow A\oplus rB\not\subset(\mathcal{X}_{n}\cap A)\oplus(r+\varepsilon)B$
$\Rightarrow A\oplus(r-\varepsilon)B\not\subset(\mathcal{X}_{n}\cap A)\oplus rB.$
Therefore,
$\mathbb{P}\Big{[}\mathcal{d}_{H}\big{(}(\mathcal{X}_{n}\cap A)\oplus rB,A\oplus rB\big{)}>\varepsilon\Big{]}\leq\mathbb{P}\Big{[}A\oplus(r-\varepsilon)B\not\subset(\mathcal{X}_{n}\cap A)\oplus rB\Big{]}.$
So, Lemma B.1a states that the probability of $\mathcal{d}_{H}\big{(}(\mathcal{X}_{n}\cap A)\oplus rB,A\oplus rB\big{)}>\varepsilon$ goes to zero exponentially with $n$ for every $\varepsilon>0$. This fact and Borel-Cantelli Lemma guarantee that $\mathcal{d}_{H}\big{(}(\mathcal{X}_{n}\cap A)\oplus rB,A\oplus rB\big{)}$ converges to zero almost surely.
Three conditions are required on the set $A$ for Lemma B.1a to hold. The first one is a technical detail: $A$ is needed to be bounded. This ensures that $D\left(\varepsilon/2,A\oplus rB\right)<+\infty$ and therefore the result is nontrivial.
The second condition is that there exists some positive constant $b$ such that $f\geq b$ in $A$. It could be interesting to find an analogous result without this restriction, allowing $f$ to vanish in a small subset of $A$ (in terms of measure). But we focus just on HDR estimation, so this supposition is not limiting and no further exploration will be done in that sense.
The last condition is the cornerstone of the result. Assuming that $\Psi_{\delta}(A)=A$ for some $\delta>0$ implies that $A$ is a set with a non-zero measure and so $\mathbb{P}(A)\neq 0$. Moreover, this condition forces the set to be smooth, without any sharp edge or too narrow part. This fact combined with Assumption (M2) guarantees that the discretization $\mathcal{X}_{n}\cap A$ will eventually be fine enough to recover the set $A\oplus rB$ with the required convergence rate.
As noticed before, the proposed estimator recovers $L(\lambda)$ by inflating a discretization of the set $L(\lambda)\ominus r_{n}B$. But there is a problem: since $r_{n}$ is a random variable, $L(\lambda)\ominus r_{n}B$ is a random set. Lemma B.1a assumes that $A$ is a fixed set, and therefore, one cannot apply this result directly to show the consistency of $L_{n}(\lambda)$. Lemma B.1b offers a solution for this issue, transforming the result of Lemma B.1a into a uniform one.
The next lemma guarantees that $\mathcal{X}_{n}^{+}(\lambda)$ and $\mathcal{X}_{n}^{-}(\lambda)$ will eventually be equal to $\mathcal{X}_{n}\cap L(\lambda)$ and $\mathcal{X}_{n}\cap L(\lambda)^{c}$, respectively.
Lemma B.2.
Under the hypotheses of Theorem 4.2, if $D_{n}<\delta$ and $kD_{n}<\delta$, the following four inclusions hold for all $\lambda\in[l,u]$:
$\mathcal{X}_{n}\cap\big{(}L(\lambda)\ominus kD_{n}B\big{)}\subset\mathcal{X}_{n}^{+}(\lambda);$ (B.5)
$\mathcal{X}_{n}\cap\big{(}L(\lambda)^{c}\ominus kD_{n}B\big{)}\subset\mathcal{X}_{n}^{-}(\lambda);$ (B.6)
$\mathcal{X}_{n}^{+}(\lambda)\subset L(\lambda)\oplus kD_{n}B;$ (B.7)
$\mathcal{X}_{n}^{-}(\lambda)\subset L(\lambda)^{c}\oplus kD_{n}B.$ (B.8)
Proof.
The key is dealing with the error made when $f$ is estimated by $f_{n}$. Given any sample point, $X_{i}\in\mathcal{X}_{n}$, one derives that
$X_{i}\in L(\lambda+D_{n})$ $\Rightarrow f(X_{i})\geq\lambda+D_{n}\Rightarrow\lambda\leq f(X_{i})-D_{n}\leq f_{n}(X_{i})\Rightarrow X_{i}\in\mathcal{X}_{n}^{+}(\lambda),$
$X_{i}\notin L(\lambda-D_{n})$ $\Rightarrow f(X_{i})<\lambda-D_{n}\Rightarrow\lambda>f(X_{i})+D_{n}\geq f_{n}(X_{i})\Rightarrow X_{i}\in\mathcal{X}_{n}^{-}(\lambda).$
Consequently $\mathcal{X}_{n}\cap L(\lambda+D_{n})\subset\mathcal{X}_{n}^{+}$ and $\mathcal{X}_{n}\cap L(\lambda-D_{n})^{c}\subset\mathcal{X}_{n}^{-}$.
Since $D_{n}<\delta$ and $kD_{n}<\delta$, one deduces from Assumption (L4) that
$\mathcal{d}_{H}\big{(}L(\lambda),L(\lambda-D_{n})\big{)}<kD_{n}$ $\Rightarrow L(\lambda)\oplus kD_{n}B\supset L(\lambda-D_{n})\Rightarrow$
$\Rightarrow L(\lambda)^{c}\ominus kD_{n}B\subset L(\lambda-D_{n})^{c}$
$\mathcal{d}_{H}\big{(}L(\lambda),L(\lambda+D_{n})\big{)}<kD_{n}$ $\Rightarrow L(\lambda)\subset L(\lambda+D_{n})\oplus kD_{n}B\Rightarrow$
$\Rightarrow L(\lambda)\ominus kD_{n}B$
$\subset\big{[}L(\lambda+D_{n})\oplus kD_{n}B\big{]}\ominus kD_{n}B$
$\qquad=\Big{[}\Psi_{kD_{n}}\big{(}L(\lambda+D_{n})^{c}\big{)}\Big{]}^{c}=L(\lambda+D_{n});$
when the last equality follows from Assumption (L2). All the above implies that
$X_{i}\in L(\lambda)\ominus kD_{n}B\Rightarrow X_{i}\in\mathcal{X}_{n}^{+}(\lambda),$
$X_{i}\in L(\lambda)^{c}\ominus kD_{n}B\Rightarrow X_{i}\in\mathcal{X}_{n}^{-}(\lambda);$
for all $X_{i}\in\mathcal{X}_{n}$. The first two inclusions follow directly from these ones:
$\mathcal{X}_{n}\cap\big{(}L(\lambda)\ominus kD_{n}B\big{)}\subset\mathcal{X}_{n}^{+}(\lambda);$
$\mathcal{X}_{n}\cap\big{(}L(\lambda)^{c}\ominus kD_{n}B\big{)}\subset\mathcal{X}_{n}^{-}(\lambda).$
The other two inclusions are derived from these two taking complements:
$\mathcal{X}_{n}^{+}(\lambda)=\mathcal{X}_{n}\setminus\mathcal{X}_{n}^{-}(\lambda)\subset\mathcal{X}_{n}\setminus\big{(}L(\lambda)^{c}\ominus kD_{n}B\big{)}\subset L(\lambda)\oplus kD_{n}B;$
$\mathcal{X}_{n}^{-}(\lambda)=\mathcal{X}_{n}\setminus\mathcal{X}_{n}^{+}(\lambda)\subset\mathcal{X}_{n}\setminus\big{(}L(\lambda)\ominus kD_{n}B\big{)}\subset L(\lambda)^{c}\oplus kD_{n}B.\qed$
Lastly, the proof of Theorem 4.2 ultimately relies on the Lemmas B.3 and B.4 below.
Lemma B.3.
Under the hypotheses of Theorem 4.2, consider the random set:
$S_{L}(\lambda)=L(\lambda)\ominus(r_{n}+kD_{n})B.$
Then, there exists a natural (and non-random) number $n_{L}$ such that, if $n\geq n_{L}$, then
$S_{L}(\lambda)\oplus(r_{n}-\varepsilon_{n})B\subset\big{(}\mathcal{X}_{n}\cap S_{L}(\lambda)\big{)}\oplus r_{n}B$ (B.9)
implies
$L(\lambda)\ominus(\varepsilon_{n}+kD_{n})B\subset L_{n}(\lambda)$ (L)
for all $\lambda\in[l,u]$, $D_{n}<\min\{\delta,\delta/k\}$, and $r_{n}\in(\delta,r-\delta]$.
Proof.
Assume that $D_{n}<\min\{\delta,\delta/k\}$ and $r_{n}\in(\delta,r-\delta]$. Let $n_{L}\in\mathbb{N}$ such that $\varepsilon_{n}<\delta$ and for all $n\geq n_{L}$. Then, Assumptions (L1) and (L2), Proposition A.3b, and Lemma A.1 ensure that
$S_{L}(\lambda)\oplus(r_{n}-\varepsilon_{n})B$ $=\big{(}L(\lambda)\ominus(r_{n}+kD_{n})B\big{)}\oplus(r_{n}-\varepsilon_{n})B$
$=\Psi_{r_{n}-\varepsilon_{n}}\big{(}L(\lambda)\ominus(\varepsilon_{n}+kD_{n})B\big{)}$
$=L(\lambda)\ominus(\varepsilon_{n}+kD_{n})B.$
From (B.5) it follows
$\mathcal{X}_{n}\cap S_{L}(\lambda)=\mathcal{X}_{n}\cap\big{[}L(\lambda)\ominus(r_{n}+kD_{n})B\big{]}\subset\mathcal{X}_{n}\cap\big{[}L(\lambda)\ominus kD_{n}B\big{]}\subset\mathcal{X}^{+}_{n}(\lambda).$
And from (B.8) one deduces
$\mathcal{X}_{n}^{-}(\lambda)\oplus r_{n}B\subset L(\lambda)^{c}\oplus(r_{n}+kD_{n})B\Rightarrow\\ \Rightarrow S_{L}(\lambda)=L(\lambda)\ominus(r_{n}+kD_{n})B\subset\big{(}\mathcal{X}_{n}^{-}(\lambda)\oplus r_{n}B\big{)}^{c}.$
Consequently:
$\mathcal{X}_{n}\cap S_{L}(\lambda)=\mathcal{X}_{n}\cap S_{L}(\lambda)\cap S_{L}(\lambda)\subset\mathcal{X}^{+}_{n}(\lambda)\cap\big{(}\mathcal{X}_{n}^{-}(\lambda)\oplus r_{n}B\big{)}^{c}.$
[…]
$L(\lambda)\ominus(\varepsilon_{n}+kD_{n})B$ $=S_{L}(\lambda)\oplus(r_{n}-\varepsilon_{n})B$
$\subset\big{(}\mathcal{X}_{n}\cap S_{L}(\lambda)\big{)}\oplus r_{n}B$
$\subset\Big{(}\mathcal{X}^{+}_{n}(\lambda)\cap\big{(}\mathcal{X}_{n}^{-}(\lambda)\oplus r_{n}B\big{)}^{c}\Big{)}\oplus r_{n}B$
$=L_{n}(\lambda).$
And therefore (B.9) implies (L) for all $n\geq n_{L}$. ∎
Lemma B.4.
Under the hypotheses of Theorem 4.2, let
$\xi=\min_{x\in L(l)}\mathcal{d}\big{(}x,L(l/2)^{c}\big{)}=d\big{(}L(l),L(l/2)^{c}\big{)};$
which exists since $L(l)$ is compact. Consider the random set:
$S_{U}(\lambda)=\big{(}L(\lambda)\oplus\xi B\big{)}\setminus\big{(}L(\lambda)\oplus kD_{n}B\big{)}.$
Then, $\xi>0$ and there exists a natural (and non-random) number $n_{U}\in\mathbb{N}$ such that, if $n\geq n_{U}$, then
$S_{U}(\lambda)\oplus(r_{n}-\varepsilon_{n})B\subset\big{(}\mathcal{X}_{n}\cap S_{U}(\lambda)\big{)}\oplus r_{n}B$ (B.10)
implies
$L_{n}(\lambda)\subset L(\lambda)\oplus(\varepsilon_{n}+kD_{n})B.$ (U)
for all $\lambda\in[l,u]$, $D_{n}<\min\{\xi,\delta,\xi/2k,\delta/2k\}$, and $r_{n}\in(\delta,r-\delta]$.
Proof.
First, let us show that $\xi>0$. By compactness, there exists $x_{\mathrm{min}}\in L(l)$ such that
$\mathcal{d}\big{(}x_{\mathrm{min}},L(l/2)^{c}\big{)}=\min_{x\in L(l)}\mathcal{d}\big{(}x,L(l/2)^{c}\big{)}=\xi.$
Since $f$ is continuous at $x_{\mathrm{min}}\in L(l)$, there exists $\epsilon>0$ such that every $y\in B_{\epsilon}(x_{\mathrm{min}})$ verifies:
$\frac{l}{2}-f(x_{\mathrm{min}})<f(y)-f(x_{\mathrm{min}})\Rightarrow\frac{l}{2}<f(y).$
Therefore $B_{\epsilon}(x_{\mathrm{min}})\subset L(l/2)$ and so $\xi=\mathcal{d}\big{(}x_{\mathrm{min}},L(l/2)^{c}\big{)}\geq\epsilon>0$. Note that from the definition of $\xi$ it follows that
$L(\lambda)\oplus\xi B\subset L(l)\oplus\xi B\subset L\left(\frac{l}{2}\right)$
holds for all $\lambda\in[l,u]$.
Now, assume that $D_{n}<\min\{\xi,\delta,\xi/2k,\delta/2k\}$ and $r_{n}\in(\delta,r-\delta]$. Let $n_{U}\in\mathbb{N}$ such that $\varepsilon_{n}<\delta/2$ for all $n\geq n_{U}$. Lemma A.3 guarantees that
$S_{U}(\lambda)\oplus(r_{n}-\varepsilon_{n})B=\Big{[}L(\lambda)\oplus(r_{n}-\varepsilon_{n}+\xi)B\Big{]}\setminus\Big{[}\big{(}L(\lambda)\oplus kD_{n}B\big{)}\ominus(r_{n}-\varepsilon_{n})B\Big{]}.$
Moreover, Proposition A.3b yields
$\big{(}L(\lambda)\oplus kD_{n}B\big{)}\ominus(r_{n}-\varepsilon_{n})B=$ $\Big{[}\Psi_{kD_{n}}\big{(}L(\lambda)^{c}\big{)}\Big{]}^{c}\ominus(r_{n}-\varepsilon_{n}-kD_{n})B$
$=$ $L(\lambda)\ominus(r_{n}-\varepsilon_{n}-kD_{n})B;$
where the last equality is guaranteed by Assumption (L2). Therefore,
$S_{U}(\lambda)\oplus(r_{n}-\varepsilon_{n})B=\big{[}L(\lambda)\oplus(r_{n}-\varepsilon_{n}+\xi)B\big{]}\setminus\big{[}L(\lambda)\ominus(r_{n}-\varepsilon_{n}-kD_{n})B\big{]},$
and
$\big{[}S_{U}(\lambda)\oplus(r_{n}-\varepsilon_{n})B\big{]}^{c}=\big{[}L(\lambda)^{c}\ominus(r_{n}-\varepsilon_{n}+\xi)B\big{]}\cup\big{[}L(\lambda)\ominus(r_{n}-\varepsilon_{n}-kD_{n})B\big{]}.$ (B.11)
Besides, it follows from (B.6) that
$\mathcal{X}_{n}\cap S_{U}(\lambda)=\mathcal{X}_{n}\cap\Big{[}\big{(}L(\lambda)\oplus\xi B\big{)}\setminus\big{(}L(\lambda)\oplus kD_{n}B\big{)}\Big{]}\\ \subset\mathcal{X}_{n}\cap\big{(}L(\lambda)^{c}\ominus kD_{n}B\big{)}\subset\mathcal{X}_{n}^{-}(\lambda).$
Taking all the above into consideration, (B.10) implies
$S_{U}(\lambda)\oplus(r_{n}-\varepsilon_{n})B\subset\mathcal{X}_{n}^{-}(\lambda)\oplus r_{n}B\Rightarrow(\mathcal{X}_{n}^{-}(\lambda)\oplus r_{n}B)^{c}\subset\big{[}S_{U}(\lambda)\oplus(r_{n}-\varepsilon_{n})B\big{]}^{c}.$
And from (B.7) and (B.11) one deduces that
$\mathcal{X}_{n}^{+}$ $(\lambda)\cap\big{(}\mathcal{X}_{n}^{-}(\lambda)\oplus r_{n}B\big{)}^{c}$
$\subset$ $\Big{(}L(\lambda)\oplus kD_{n}B\Big{)}\cap\Big{(}\big{[}L(\lambda)^{c}\ominus(r_{n}-\varepsilon_{n}+\xi)B\big{]}\cup\big{[}L(\lambda)\ominus(r_{n}-\varepsilon_{n}-kD_{n})B\big{]}\Big{)}$
$=$ $L(\lambda)\ominus(r_{n}-\varepsilon_{n}-kD_{n})B;$
where the last inequality holds since $r_{n}-\varepsilon_{n}+\xi>\xi>kD_{n}$ for all $n\geq n_{U}$. This in turn implies
$L_{n}(\lambda)\subset$ $\big{[}L(\lambda)\ominus(r_{n}-\varepsilon_{n}-kD_{n})B\big{]}\oplus r_{n}B$
$=$ $\Psi_{r_{n}-\varepsilon_{n}-kD_{n}}\big{[}L(\lambda)\big{]}\oplus(\varepsilon_{n}+kD_{n})B$
$=$ $L(\lambda)\oplus(\varepsilon_{n}+kD_{n})B;$
where we have applied Proposition A.3a and Assumption (L1). Consequently, (B.10) implies (U) for all $n\geq n_{U}$. ∎
We can finally proof Theorem 4.2.
Proof of Theorem 4.2.
Let
$\varepsilon_{n}=\left(\frac{c\log n}{n}\right)^{2/(d+1)}$ (B.12)
with $c>0$ a constant that will be specified later on. The basis of the proof is to bound
$\mathbb{P}\Big{[}\sup_{\lambda\in[l,u]}\mathcal{d}_{H}\big{(}L(\lambda),L_{n}(\lambda)\big{)}>\varepsilon_{n}+kD_{n},r_{n}\in(\delta,r-\delta]\text{ and }D_{n}<s\Big{]}$ (B.13)
where $k>0$ is the constant present in Assumption (L4) and $s>0$ is a fixed value satisfying some conditions (that will be detailed below). Once expression (B.13) is bounded, one applies Borel-Cantelli Lemma to ensure that, with probability one, there exists a $n_{0}\in\mathbb{N}$ such that, if $n\geq n_{0}$, then
$\sup_{\lambda\in[l,u]}\mathcal{d}_{H}\big{(}L(\lambda),L_{n}(\lambda)\big{)}\leq\varepsilon_{n}+kD_{n}\hskip 10.00002pt\text{or}\hskip 10.00002ptr_{n}\notin(\delta,r-\delta]\hskip 10.00002pt\text{or}\hskip 10.00002ptD_{n}\geq s.$
Hence, since $r_{n}\notin(\delta,r-\delta]$ and $D_{n}\geq s$ will eventually not occur by the assumptions considered, one concludes that Theorem 4.2 holds.
Let us then proceed to bound (B.13). Given any $\lambda\in[l,u]$, it is easy to check that
$L(\lambda)\ominus(\varepsilon_{n}+kD_{n})B\subset L_{n}(\lambda),$ (L)
$L_{n}(\lambda)\subset L(\lambda)\oplus(\varepsilon_{n}+kD_{n})B.$ (U)
imply $\mathcal{d}_{H}\big{(}L(\lambda),L_{n}(\lambda)\big{)}\leq\varepsilon_{n}+kD_{n}$ for large $n$. Suppose that $D_{n}<\delta/k$, and let $n_{1}\in\mathbb{N}$ be such that [^1] Notice that $n_{1}$ is not random since it only depends on the real sequence $\varepsilon_{n}$. every $n\geq n_{1}$ verifies $\varepsilon_{n}<r-\delta$. Then, if (L) is true, Assumption (L1) guarantees that
$L(\lambda)$ $=\big{[}L(\lambda)\ominus(\varepsilon_{n}+kD_{n})B\big{]}\oplus(\varepsilon_{n}+kD_{n})B$
$\subset L_{n}(\lambda)\oplus(\varepsilon_{n}+kD_{n})B.$
Consequently, if $D_{n}<\delta/k$, (L), and (U) hold, then $\mathcal{d}_{H}\big{(}L(\lambda),L_{n}(\lambda)\big{)}\leq\varepsilon_{n}+kD_{n}$ also holds for all $n\geq n_{1}$. Hence:
$\mathbb{P}\Big{[}\sup_{\lambda\in[l,u]}\mathcal{d}_{H}\big{(}L(\lambda),L_{n}(\lambda)\big{)}>\varepsilon_{n}+kD_{n},r_{n}\in(\delta,r-\delta]\text{ and }D_{n}<s\Big{]}\\ \leq\mathbb{P}\Big{[}\bigcup_{\lambda\in[l,u]}\big{\{}L_{n}(\lambda)\not\subset L(\lambda)\oplus(\varepsilon_{n}+kD_{n})B\big{\}},r_{n}\in(\delta,r-\delta]\text{ and }D_{n}<s\Big{]}\\ +\mathbb{P}\Big{[}\bigcup_{\lambda\in[l,u]}\big{\{}L(\lambda)\ominus(\varepsilon_{n}+kD_{n})B\not\subset L_{n}(\lambda)\big{\}},r_{n}\in(\delta,r-\delta]\text{ and }D_{n}<s\Big{]}$ (B.14)
for all $n\geq n_{1}$, where $s$ is any constant smaller than $\delta/k$.
The goal now is to translate the events (L) and (U) into events similar to those in Lemma B.1. We could then apply this result to bound these two probabilities.
For (L), define the (random) set
$S_{L}(\lambda)=L(\lambda)\ominus(r_{n}+kD_{n})B.$
Lemma B.3 guarantees that there exists a natural value $n_{L}\in\mathbb{N}$ such that, if $n\geq n_{L}$, then
$S_{L}(\lambda)\oplus(r_{n}-\varepsilon_{n})B\subset\big{(}\mathcal{X}_{n}\cap S_{L}(\lambda)\big{)}\oplus r_{n}B$
implies (L) for all $D_{n}<\min\{\delta,\delta/k\}$ and $r_{n}\in(\delta,r-\delta]$. Therefore:
$\mathbb{P}\Big{[}\bigcup_{\lambda\in[l,u]}\left\{L_{n}(\lambda)\not\subset L(\lambda)\oplus(\varepsilon_{n}+kD_{n})B\right\},r_{n}\in(\delta,r-\delta]\text{ and }D_{n}<s\Big{]}\\ \leq\mathbb{P}\Big{[}\bigcup_{\lambda\in[l,u]}\left\{S_{L}(\lambda)\oplus(r_{n}-\varepsilon_{n})B\not\subset\big{(}\mathcal{X}_{n}\cap S_{L}(\lambda)\big{)}\oplus r_{n}B\right\},\\ r_{n}\in(\delta,r-\delta]\text{ and }D_{n}<s\Big{]}$ (B.15)
for all $n\geq n_{L}$ and any constant $s\leq\min\{\delta,\delta/k\}$.
For (U), the idea is to show that the subsample $\mathcal{X}_{n}^{-}(\lambda)$ will eventually surround the set $L(\lambda)$ and, consequently, $L_{n}(\lambda)$ cannot be much larger than $L(\lambda)$. But, in order to prove this result using Lemma B.1, one needs to find a ring-shaped set surrounding $L(\lambda)$ where the density function $f$ does not vanish. For this purpose, take
$\xi=\min_{x\in L(l)}\mathcal{d}\big{(}x,L(l/2)^{c}\big{)}=\mathcal{d}\big{(}L(l),L(l/2)^{c}\big{)},$
which exists since $L(l)$ is compact. The required ring-shaped set is given by
$S_{U}(\lambda)=\big{(}L(\lambda)\oplus\xi B\big{)}\setminus\big{(}L(\lambda)\oplus kD_{n}B\big{)}.$
Lemma B.4 ensures that $\xi>0$ and there exists a $n_{U}\in\mathbb{N}$ such that, if $n\geq n_{U}$, then
$S_{U}(\lambda)\oplus(r_{n}-\varepsilon_{n})B\subset(\mathcal{X}_{n}\cap S_{U}(\lambda))\oplus r_{n}B$
implies (U) for $D_{n}<\min\{\xi,\delta,\xi/2k,\delta/2k\}$ and $r_{n}\in(\delta,r-\delta]$. Therefore:
$\mathbb{P}\Big{[}\bigcup_{\lambda\in[l,u]}\big{\{}L(\lambda)\ominus(\varepsilon_{n}+kD_{n})B\not\subset L_{n}(\lambda)\big{\}},r_{n}\in(\delta,r-\delta]\text{ and }D_{n}<s\Big{]}\\ \leq\mathbb{P}\Big{[}\bigcup_{\lambda\in[l,u]}\big{\{}S_{U}(\lambda)\oplus(r_{n}-\varepsilon_{n})B\not\subset\big{(}\mathcal{X}_{n}\cap S_{U}(\lambda)\big{)}\oplus r_{n}B\big{\}},\\ r_{n}\in(\delta,r-\delta]\text{ and }D_{n}<s\Big{]}$ (B.16)
for all $n\geq n_{U}$, where $s$ is smaller than $\min\{\xi,\delta,\xi/2k,\delta/2k\}$.
Consequently, taking [^2] Note that $n_{c}$ is not random since $n_{1}$, $n_{L}$, and $n_{U}$ only depend on the real sequence $\varepsilon_{n}$. $n_{c}=\max\{n_{1},n_{L},n_{U}\}$ and $s=\min\{\xi,\delta,\xi/2k,\delta/2k\}$, from (B.14), (B.15), and (B.16) one derives:
$\mathbb{P}\Big{[}\sup_{\lambda\in[l,u]}\mathcal{d}_{H}\big{(}L(\lambda),L_{n}(\lambda)\big{)}>\varepsilon_{n}+kD_{n},r_{n}\in(\delta,r-\delta]\text{ and }D_{n}<s\Big{]}\\ \leq\mathbb{P}\Big{[}\bigcup_{\lambda\in[l,u]}\big{\{}S_{L}(\lambda)\oplus(r_{n}-\varepsilon_{n})B\not\subset\big{(}\mathcal{X}_{n}\cap S_{L}(\lambda)\big{)}\oplus r_{n}B\big{\}},r_{n}\in(\delta,r-\delta]\text{ and }D_{n}<s\Big{]}\\ +\mathbb{P}\Big{[}\bigcup_{\lambda\in[l,u]}\big{\{}S_{U}(\lambda)\oplus(r_{n}-\varepsilon_{n})B\not\subset\big{(}\mathcal{X}_{n}\cap S_{U}(\lambda)\big{)}\oplus r_{n}B\big{\}},r_{n}\in(\delta,r-\delta]\text{ and }D_{n}<s\Big{]}$
for all $n\geq n_{c}$. The next step is to apply a monotonicity argument to get rid of the random sequence $r_{n}$ in the Minkowski sums. If $\delta<r_{n}$, then Proposition A.3a implies that, for any $S\subset M$,
$S\oplus(\delta-\varepsilon_{n})B\subset\big{(}\mathcal{X}_{n}\cap S\big{)}\oplus\delta B\Rightarrow S\oplus(r_{n}-\varepsilon_{n})B\subset\big{(}\mathcal{X}_{n}\cap S\big{)}\oplus r_{n}B.$
Hence
$\mathbb{P}\Big{[}\sup_{\lambda\in[l,u]}\mathcal{d}_{H}\big{(}L(\lambda),L_{n}(\lambda)\big{)}>\varepsilon_{n}+kD_{n},r_{n}\in(\delta,r-\delta]\text{ and }D_{n}<s\Big{]}\\ \leq\mathbb{P}\Big{[}\bigcup_{\lambda\in[l,u]}\big{\{}S_{L}(\lambda)\oplus(\delta-\varepsilon_{n})B\not\subset\big{(}\mathcal{X}_{n}\cap S_{L}(\lambda)\big{)}\oplus\delta B\big{\}},r_{n}\in(\delta,r-\delta]\text{ and }D_{n}<s\Big{]}\\ +\mathbb{P}\Big{[}\bigcup_{\lambda\in[l,u]}\big{\{}S_{U}(\lambda)\oplus(\delta-\varepsilon_{n})B\not\subset\big{(}\mathcal{X}_{n}\cap S_{U}(\lambda)\big{)}\oplus\delta B\big{\}},r_{n}\in(\delta,r-\delta]\text{ and }D_{n}<s\Big{]}$
for all $n\geq n_{c}$. Now, these two probabilities can be bounded by applying Lemma B.1.
Let $N_{L}\in\mathbb{N}$ such that if $n\geq N_{L}$ then $\varepsilon_{n}\leq\min\{\delta/2,\rho\}$, where $\rho$ is the constant from Assumption (M2). One checks that
  • $S_{L}(\lambda)=L(\lambda)\ominus(kD_{n}+r_{n})B\subset L(\lambda)\subset L(l/2).$
  • $\Psi_{\delta/2}\big{(}S_{L}(\lambda)\big{)}=\Psi_{\delta/2}\big{(}L(\lambda)\ominus(kD_{n}+r_{n})B\big{)}=L(\lambda)\ominus(kD_{n}+r_{n})B$ for all $n\geq N_{L}$ and $D_{n}<s$ by Lemma A.1.
So, Lemma B.1b guarantees that
$\mathbb{P}\Big{[}\bigcup_{\lambda\in[l,u]}$$\big{\{}S_{L}(\lambda)\oplus(\delta-\varepsilon_{n})B\not\subset\big{(}\mathcal{X}_{n}\cap S_{L}(\lambda)\big{)}\oplus\delta B\big{\}},r_{n}\in(\delta,r-\delta]\text{ and }D_{n}<s\Big{]}$
$\leq D\left(\frac{\varepsilon_{n}}{4},L(l/2)\oplus\delta B\right)^{2}\exp\left[-na\frac{l}{2}\min\left\{\frac{\delta}{4},\rho\right\}^{(d-1)/2}\left(\frac{\varepsilon_{n}}{4}\right)^{(d+1)/2}\right]$
for all $n\geq N_{L}$.
Similar reasoning applies to $S_{U}(\lambda)$. Let $N_{U}\in\mathbb{N},N_{U}\geq N_{L}$ such that $\varepsilon_{n}<\xi/6$ for all $n\geq N_{U}$. Then:
  • $S_{U}(\lambda)=\big{(}L(\lambda)\oplus\xi B\big{)}\setminus\big{(}L(\lambda)\oplus kD_{n}B\big{)}\subset L(\lambda)\oplus\xi B\subset L(l/2)$ by the definition of $\xi$.
  • $\Psi_{\xi/6}(S_{U}(\lambda))=\Psi_{\xi/6}\Big{[}\big{(}L(\lambda)\oplus\xi B\big{)}\setminus\big{(}L(\lambda)\oplus kD_{n}B\big{)}\Big{]}=\big{(}L(\lambda)\oplus\xi B\big{)}\setminus\big{(}L(\lambda)\oplus kD_{n}B\big{)}$ for all $n\geq N_{U}$ and $D_{n}<s$ by Lemma A.2.
Therefore, Lemma B.1b again ensures
$\mathbb{P}\Big{[}\bigcup_{\lambda\in[l,u]}\big{\{}S_{U}(\lambda)\oplus(\delta-\varepsilon_{n})B\not\subset\big{(}\mathcal{X}_{n}\cap S_{U}(\lambda)\big{)}\oplus\delta B\big{\}},r_{n}\in(\delta,r-\delta]\text{ and }D_{n}<s\Big{]}\\ \leq D\left(\frac{\varepsilon_{n}}{4},L(l/2)\oplus\delta B\right)^{2}\exp\left[-na\frac{l}{2}\min\left\{\frac{\xi}{12},\rho\right\}^{(d-1)/2}\left(\frac{\varepsilon_{n}}{4}\right)^{(d+1)/2}\right]$
for all $n\geq N_{U}$.
Let [^3] Again, note that $n_{p}$ is not random since both $n_{c}$ and $N_{U}$ only depend on the real sequence $\varepsilon_{n}$. $n_{p}=\max\{n_{c},N_{U}\}$. Finally, for all $n\geq n_{p}$ one deduces that
$\mathbb{P}\Big{[}\sup_{\lambda\in[l,u]}\mathcal{d}_{H}\big{(}$ $L(\lambda),L_{n}(\lambda)\big{)}>\varepsilon_{n}+kD_{n},r_{n}\in(\delta,r-\delta]\text{ and }D_{n}<s\Big{]}$
$\leq$ $D\left(\frac{\varepsilon_{n}}{4},L(l/2)\oplus\delta B\right)^{2}\exp\left[-na\frac{l}{2}\min\left\{\frac{\delta}{4},\rho\right\}^{(d-1)/2}\left(\frac{\varepsilon_{n}}{4}\right)^{(d+1)/2}\right]$
$+D\left(\frac{\varepsilon_{n}}{4},L(l/2)\oplus\delta B\right)^{2}\exp\left[-na\frac{l}{2}\min\left\{\frac{\xi}{12},\rho\right\}^{(d-1)/2}\left(\frac{\varepsilon_{n}}{4}\right)^{(d+1)/2}\right]$
$\leq$ $2D\left(\frac{\varepsilon_{n}}{4},L(l/2)\oplus\delta B\right)^{2}\exp\left[-na\frac{l}{2}\min\left\{\frac{\delta}{4},\frac{\xi}{12},\rho\right\}^{(d-1)/2}\left(\frac{\varepsilon_{n}}{4}\right)^{(d+1)/2}\right].$
Denote $G=a\frac{l}{2}\min\left\{\frac{\delta}{4},\frac{\xi}{12},\rho\right\}^{(d-1)/2}4^{-(d+1)/2}$. Then, the above rewrites as
$\mathbb{P}\Big{[}\sup_{\lambda\in[l,u]}\mathcal{d}_{H}\big{(}L(\lambda),L_{n}(\lambda)\big{)}>\varepsilon_{n}+kD_{n},r_{n}\in(\delta,r-\delta]\text{ and }D_{n}<s\Big{]}\\ \leq 2D\left(\frac{\varepsilon_{n}}{4},L(l/2)\oplus\delta B\right)^{2}\exp\left[-Gn\varepsilon_{n}^{(d+1)/2}\right]$
for all $n\geq n_{p}$. Since $\varepsilon_{n}<\rho$ for all $n\geq n_{p}$, Assumption (M3) ensures that
$\mathbb{P}\Big{[}\sup_{\lambda\in[l,u]}\mathcal{d}_{H}\big{(}L(\lambda),L_{n}(\lambda)\big{)}>\varepsilon_{n}+kD_{n},r_{n}\in(\delta,r-\delta]\text{ and }D_{n}<s\Big{]}\\ \leq 2K\varepsilon_{n}^{-2d}\exp\left[-Gn\varepsilon_{n}^{(d+1)/2}\right].$
Then, the definition of $\varepsilon_{n}$ in (B.12) yields
$\mathbb{P}\Big{[}\sup_{\lambda\in[l,u]}\mathcal{d}_{H}\big{(}L(\lambda),L_{n}(\lambda)\big{)}>\varepsilon_{n}+kD_{n},r_{n}\in(\delta,r-\delta]\text{ and }D_{n}<s\Big{]}\\ \leq\tilde{C}n^{4d/(d+1)-Gc}(\log n)^{-4d/(d+1)}$
where $\tilde{C}=2Cc^{-4d/(d+1)}$.
Choosing $c$ large enough guarantees that
$\sum_{n=n_{p}}^{+\infty}\mathbb{P}\Big{[}\sup_{\lambda\in[l,u]}\mathcal{d}_{H}\big{(}L(\lambda),L_{n}(\lambda)\big{)}>\varepsilon_{n}+kD_{n},r_{n}\in(\delta,r-\delta]\text{ and }D_{n}<s\Big{]}<+\infty.$
Therefore, the Borel-Cantelli Lemma ensures
$\mathbb{P}\bigg{(}\limsup_{n\in\mathbb{N}}\Big{[}\sup_{\lambda\in[l,u]}\mathcal{d}_{H}\big{(}L(\lambda),L_{n}(\lambda)\big{)}>\varepsilon_{n}+kD_{n},r_{n}\in(\delta,r-\delta]\text{ and }D_{n}<s\Big{]}\bigg{)}=0\Leftrightarrow\\ \mathbb{P}\bigg{(}\limsup_{n\in\mathbb{N}}\Big{[}\sup_{\lambda\in[l,u]}\mathcal{d}_{H}\big{(}L(\lambda),L_{n}(\lambda)\big{)}>\varepsilon_{n}+kD_{n}\Big{]}\cap\\ \cap\limsup_{n\in\mathbb{N}}\Big{[}r_{n}\in(\delta,r-\delta]\Big{]}\cap\limsup_{n\in\mathbb{N}}\Big{[}D_{n}<s\Big{]}\bigg{)}=0.$
Since it is assumed that
$\mathbb{P}\Big{[}\limsup_{n\in\mathbb{N}}\big{(}D_{n}<s\big{)}\Big{]}=\mathbb{P}\Big{[}\limsup_{n\in\mathbb{N}}\big{(}r_{n}\in(\delta,r-\delta]\big{)}\Big{]}=1,$
one concludes that
$\mathbb{P}\bigg{(}\limsup_{n\in\mathbb{N}}\Big{[}\sup_{\lambda\in[l,u]}\mathcal{d}_{H}\big{(}L(\lambda),L_{n}(\lambda)\big{)}>\varepsilon_{n}+kD_{n}\Big{]}\bigg{)}=0\Leftrightarrow\\ \mathbb{P}\bigg{(}\liminf_{n\in\mathbb{N}}\Big{[}\sup_{\lambda\in[l,u]}\mathcal{d}_{H}\big{(}L(\lambda),L_{n}(\lambda)\big{)}\leq\varepsilon_{n}+kD_{n}\Big{]}\bigg{)}=1.$
That it is to say, with probability $1$, there exists a $n_{0}\in\mathbb{N}$ such that
$\sup_{\lambda\in[l,u]}\mathcal{d}_{H}\big{(}L(\lambda),L_{n}(\lambda)\big{)}\leq\varepsilon_{n}+kD_{n}\leq C\max\left\{D_{n},\left(\frac{\log n}{n}\right)^{2/(d+1)}\right\}$
for all $n\geq n_{0}$, where $C=k+c^{2/(d+1)}$. ∎
B.2 Proofs of Theorems 5.1 and 5.2
Theorem 5.1 is a direct consequence of Theorem 4.2.
Proof of Theorem 5.1.
From the triangular inequality one deduces that
$\sup_{\gamma\in[\underline{\gamma},\overline{\gamma}]}\mathcal{d}_{H}\big{(}L(\lambda_{\gamma}),L_{n}(\tilde{\lambda}_{\gamma,n})\big{)}\leq\sup_{\gamma\in[\underline{\gamma},\overline{\gamma}]}\mathcal{d}_{H}\big{(}L(\lambda_{\gamma}),L(\tilde{\lambda}_{\gamma,n})\big{)}+\sup_{\gamma\in[\underline{\gamma},\overline{\gamma}]}\mathcal{d}_{H}\big{(}L(\tilde{\lambda}_{\gamma,n}),L_{n}(\tilde{\lambda}_{\gamma,n})\big{)}.$
If $T_{n}<\delta$, then Assumption (L4) implies
$\sup_{\gamma\in[\underline{\gamma},\overline{\gamma}]}\mathcal{d}_{H}\big{(}L(\lambda_{\gamma}),L(\tilde{\lambda}_{\gamma,n})\big{)}\leq kT_{n}.$
On the other hand, if $T_{n}<\min\{u-\lambda_{\overline{\gamma}},\lambda_{\underline{\gamma}}-l\}$, then $\tilde{\lambda}_{\gamma,n}\in[l,u]$ for all $\gamma\in[\underline{\gamma},\overline{\gamma}]$ and therefore
$\sup_{\gamma\in[\underline{\gamma},\overline{\gamma}]}\mathcal{d}_{H}\big{(}L(\tilde{\lambda}_{\gamma,n}),L_{n}(\tilde{\lambda}_{\gamma,n})\big{)}\leq\sup_{\lambda\in[l,u]}\mathcal{d}_{H}\big{(}L(\lambda),L_{n}(\lambda)\big{)}.$
Since $T_{n}\rightarrow 0$ almost surely, $T_{n}<\min\{u-\lambda_{\overline{\gamma}},\lambda_{\underline{\gamma}}-l,\delta\}$ happens almost surely. Consequently, all the above jointly with Theorem 4.2 guarantee that there exists a constant $C>0$ such that, almost surely, it holds that
$\sup_{\gamma\in[\underline{\gamma},\overline{\gamma}]}\mathcal{d}_{H}\big{(}L(\lambda_{\gamma}),L_{n}(\tilde{\lambda}_{\gamma,n})\big{)}\leq$ $kT_{n}+C\max\left\{D_{n},\left(\frac{\log n}{n}\right)^{\frac{2}{d+1}}\right\}$
$\leq$ $(k+C)\max\left\{D_{n},T_{n},\left(\frac{\log n}{n}\right)^{\frac{2}{d+1}}\right\}.\qed$
The proof of Theorem 5.2 requires Propositions B.1 and B.2 below. The first of them allows to bound the distance between $\lambda_{\gamma}$ and $\hat{\lambda}_{\gamma,n}$ in a special type of metric called Lévy metric (check Cadre et al.,, 2013 for more details about this metric).
Proposition B.1.
Let $M$ be a Riemannian manifold and $f$ a density function on $M$. Let $X_{1},\ldots,X_{n}$ be a random sample from $f$. Let $f_{n}$ be an estimator of $f$.
Consider the random variables:
$D_{n}=\sup_{x\in M}\left|f(x)-f_{n}(x)\right|,\qquad S_{n}=\sup_{\lambda\in\mathbb{R}}\left|\mathbb{P}\big{[}L(\lambda)\big{]}-\mathbb{P}_{n}\big{[}L(\lambda)\big{]}\right|.$
Hence:
$\lambda_{\gamma-M_{n}}-M_{n}\leq\hat{\lambda}_{\gamma,n}\leq\lambda_{\gamma+M_{n}}+M_{n},$
for all $\gamma\in(0,1)$, where $M_{n}=\max\{D_{n},S_{n}\}$.
Proof.
The definition of $D_{n}$ directly yields
$L(\lambda+D_{n})\subset f_{n}^{-1}\big{(}[\lambda,+\infty)\big{)}\subset L(\lambda-D_{n}),$
for every $\lambda\in\mathbb{R}$. So,
$\mathbb{P}_{n}\big{[}L(\lambda+D_{n})\big{]}\leq\mathbb{P}_{n}\big{[}f_{n}^{-1}\big{(}[\lambda,+\infty)\big{)}\big{]}\leq\mathbb{P}_{n}\big{[}L(\lambda-D_{n})\big{]},\quad\forall\lambda\in\mathbb{R}.$ (B.17)
On the other hand, from the definition of $S_{n}$, it follows that
$\mathbb{P}\big{[}L(\lambda)\big{]}-S_{n}\leq\mathbb{P}_{n}\big{[}L(\lambda)\big{]}\leq\mathbb{P}\big{[}L(\lambda)\big{]}+S_{n},\quad\forall\lambda\in\mathbb{R}.$ (B.18)
Consequently, (B.17) and (B.18) guarantee that
$\mathbb{P}\big{[}L(\lambda+D_{n})\big{]}-S_{n}\leq\mathbb{P}_{n}\big{[}f_{n}^{-1}\big{(}[\lambda,+\infty)\big{)}\big{]}\leq\mathbb{P}\big{[}L(\lambda-D_{n})\big{]}+S_{n},\quad\forall\lambda\in\mathbb{R}.$
Taking into account that $\mathbb{P}\big{[}L(\lambda)\big{]}$ is non-increasing in $\lambda$, this yields
$\mathbb{P}\big{[}L(\lambda+M_{n})\big{]}-M_{n}\leq\mathbb{P}_{n}\big{[}f_{n}^{-1}\big{(}[\lambda,+\infty)\big{)}\big{]}\leq\mathbb{P}\big{[}L(\lambda-M_{n})\big{]}+M_{n},\quad\forall\lambda\in\mathbb{R}.$
From the previous inequality one deduces that the following two contents
$\left\{\lambda\in\mathbb{R}\colon\mathbb{P}_{n}\big{[}f_{n}^{-1}\big{(}[\lambda,+\infty)\big{)}\big{]}\geq 1-\gamma\right\}\subset$$\left\{\lambda\in\mathbb{R}\colon\mathbb{P}\big{[}L(\lambda-M_{n})\big{]}\geq 1-\gamma-M_{n}\right\},$
$\left\{\lambda\in\mathbb{R}\colon\mathbb{P}_{n}\big{[}f_{n}^{-1}\big{(}[\lambda,+\infty)\big{)}\big{]}\geq 1-\gamma\right\}\supset$$\left\{\lambda\in\mathbb{R}\colon\mathbb{P}\big{[}L(\lambda+M_{n})\big{]}\geq 1-\gamma+M_{n}\right\};$
hold for all $\gamma\in(0,1)$. Hence,
$\hat{\lambda}_{\gamma,n}=$ $\sup\left\{\lambda\in\mathbb{R}\colon\mathbb{P}_{n}\big{[}f_{n}^{-1}\big{(}[\lambda,+\infty)\big{)}\big{]}\geq 1-\gamma\right\}$
$\leq$ $\sup\left\{\lambda\in\mathbb{R}\colon\mathbb{P}\big{[}L(\lambda-M_{n})\big{]}\geq 1-\gamma-M_{n}\right\}$
$=$ $\sup\left\{\lambda\in\mathbb{R}\colon\mathbb{P}\big{[}L(\lambda)\big{]}\geq 1-\gamma-M_{n}\right\}+M_{n}=\lambda_{\gamma+M_{n}}+M_{n},$
$\hat{\lambda}_{\gamma,n}=$ $\sup\left\{\lambda\in\mathbb{R}\colon\mathbb{P}_{n}\big{[}f_{n}^{-1}\big{(}[\lambda,+\infty)\big{)}\big{]}\geq 1-\gamma\right\}$
$\geq$ $\sup\left\{\lambda\in\mathbb{R}\colon\mathbb{P}\big{[}L(\lambda+M_{n})\big{]}\geq 1-\gamma+M_{n}\right\}$
$=$ $\sup\left\{\lambda\in\mathbb{R}\colon\mathbb{P}\big{[}L(\lambda)\big{]}\geq 1-\gamma+M_{n}\right\}-M_{n}=\lambda_{\gamma-M_{n}}-M_{n},$
also hold for all $\gamma\in(0,1)$, completing the proof ∎
Once proven that the distance between $\lambda_{\gamma}$ and $\hat{\lambda}_{\gamma,n}$ is bounded by the maximum of $D_{n}$ and $S_{n}$, one needs to show how fast these quantities converge to zero. Proposition B.2 ensures that $S_{n}$ is smaller than $(\log n/n)^{1/2}$ almost surely.
Proposition B.2.
Let $M$ be a Riemannian manifold and $f$ a density function on $M$. Let $X_{1},\ldots,X_{n}$ be a random sample from $f$. Then
$\sup_{\lambda\in\mathbb{R}}\left|\mathbb{P}\big{[}L(\lambda)\big{]}-\mathbb{P}_{n}\big{[}L(\lambda)\big{]}\right|\leq\sqrt{\frac{\log n}{n}}$
almost surely, where
$\mathbb{P}_{n}\big{[}L(\lambda)\big{]}=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\big{[}f(X_{i})\geq\lambda\big{]}.$
Proof.
Consider the random variable $Y=-f(X)$ and its cumulative distribution function defined by
$F_{Y}(t)=\mathbb{P}[Y\leq t]=\mathbb{P}[-f(X)\leq t]=\mathbb{P}[f(X)\geq-t]=\mathbb{P}[L(-t)].$
One can estimate $F_{Y}$ from the random sample $-f(X_{1}),\ldots,-f(X_{n})$ with the empirical distribution function:
$F_{Y,n}(t)=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\big{[}-f(X_{i})\leq t\big{]}=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\big{[}f(X_{i})\geq-t\big{]}=\mathbb{P}_{n}\big{[}L(-t)\big{]}.$
Therefore
$\sup_{\lambda\in\mathbb{R}}\left|\mathbb{P}\big{[}L(\lambda)\big{]}-\mathbb{P}_{n}\big{[}L(\lambda)\big{]}\right|=\sup_{t\in\mathbb{R}}\left|F_{Y}(t)-F_{Y,n}(t)\right|.$
The Dvoretzky–Kiefer–Wolfowitz inequality (Massart,, 1990) yields
$\mathbb{P}\left(\sup_{t\in\mathbb{R}}\left|F_{Y}(t)-F_{Y,n}(t)\right|>\sqrt{\frac{\log n}{n}}\right)\leq 2\exp(-2\log n)=2n^{-2}.$
Consequently
$\sum_{n=1}^{+\infty}\mathbb{P}\left(\sup_{\lambda\in\mathbb{R}}\left|\mathbb{P}\big{[}L(\lambda)\big{]}-\mathbb{P}_{n}\big{[}L(\lambda)\big{]}\right|>\sqrt{\frac{\log n}{n}}\right)\leq 2\sum_{n=1}^{+\infty}n^{-2}<+\infty.$
And the Borel-Cantelli Lemma ensures that
$\mathbb{P}\left[\limsup_{n\in\mathbb{N}}\left(\sup_{\lambda\in\mathbb{R}}\left|\mathbb{P}\big{[}L(\lambda)\big{]}-\mathbb{P}_{n}\big{[}L(\lambda)\big{]}\right|>\sqrt{\frac{\log n}{n}}\right)\right]=0\Rightarrow\\ \Rightarrow\mathbb{P}\left[\liminf_{n\in\mathbb{N}}\left(\sup_{\lambda\in\mathbb{R}}\left|\mathbb{P}\big{[}L(\lambda)\big{]}-\mathbb{P}_{n}\big{[}L(\lambda)\big{]}\right|\leq\sqrt{\frac{\log n}{n}}\right)\right]=1;$
showing that $\sup_{\lambda\in\mathbb{R}}\left|\mathbb{P}\big{[}L(\lambda)\big{]}-\mathbb{P}_{n}\big{[}L(\lambda)\big{]}\right|\leq\sqrt{\frac{\log n}{n}}$ holds almost surely. ∎
Theorem 5.2 then follows from Propositions B.1 and B.2.
Proof of Theorem 5.2.
Let
$S_{n}=\sup_{\lambda\geq 0}\left|\mathbb{P}\big{[}L(\lambda)\big{]}-\mathbb{P}_{n}\big{[}L(\lambda)\big{]}\right|$
and $M_{n}=\max\{D_{n},S_{n}\}$. Proposition B.1 guarantees that
$\lambda_{\gamma-M_{n}}-M_{n}\leq\hat{\lambda}_{\gamma,n}\leq\lambda_{\gamma+M_{n}}+M_{n},\quad\forall\gamma\in[\underline{\gamma},\overline{\gamma}].$
If $M_{n}<\delta$, then Assumption P yields
$\sup_{\gamma\in[\underline{\gamma},\overline{\gamma}]}\left|\lambda_{\gamma}-\lambda_{\gamma\pm M_{n}}\right|\leq kM_{n}.$
Therefore, if $M_{n}<\delta$, it follows that
$\lambda_{\gamma}-(1+k)M_{n}\leq\hat{\lambda}_{\gamma,n}\leq\lambda_{\gamma}+(1+k)M_{n},\quad\forall\gamma\in[\underline{\gamma},\overline{\gamma}];$
and, consequently,
$\sup_{\gamma\in[\underline{\gamma},\overline{\gamma}]}\left|\lambda_{\gamma}-\hat{\lambda}_{\gamma,n}\right|\leq(1+k)M_{n}.$
Proposition B.2 ensures that $S_{n}$ converges to zero almost surely. Since $D_{n}$ also converges to zero almost surely by assumption, $M_{n}\rightarrow 0$ almost surely. Hence $M_{n}<\delta$ occurs almost surely, and then
$\sup_{\gamma\in[\underline{\gamma},\overline{\gamma}]}\left|\lambda_{\gamma}-\hat{\lambda}_{\gamma,n}\right|\leq(1+k)M_{n},$
holds almost surely too. In addition, Proposition B.2 also shows that $S_{n}\leq\sqrt{\frac{\log n}{n}}$ almost surely. Therefore,
$\sup_{\gamma\in[\underline{\gamma},\overline{\gamma}]}\left|\lambda_{\gamma}-\hat{\lambda}_{\gamma,n}\right|\leq(1+k)\max\left\{D_{n},\sqrt{\frac{\log n}{n}}\right\}$
almost surely, concluding the proof. ∎
Appendix C Relation between Assumptions M, L and P and the smoothness of $f$ and $M$
As explained in Sections 4.1 and 4.2, some assumptions must be made on the manifold $M$ that supports the data and the HDRs of the population for the consistency of the estimator $L_{n}(\lambda)$. These assumptions, that we refer as Assumptions M and Assumptions L, are the following:
Assumptions M.
$M$ is a $d$-dimensional Riemannian manifold verifying:
  1. (M1)
    $M$ is complete and connected.
  2. (M2)
    Let $r_{1},r_{2}>0$. Let $x_{1},x_{2}\in M$ such that $\mathcal{d}(x_{1},x_{2})<r_{1}+r_{2}$. There exist two constants, $a,\rho>0$, (both not depending on $x_{1}$ and $x_{2}$) such that
    $\mathrm{Vol}_{M}\big{(}B_{r_{1}}[x_{1}]\cap B_{r_{2}}[x_{2}]\big{)}\geq a\min\{r_{1},r_{2},\rho\}^{(d-1)/2}\min\{\varepsilon,\rho\}^{(d+1)/2}.$
    for all $\varepsilon\leq r_{1}+r_{2}-\mathcal{d}(x_{1},x_{2})$.
  3. (M3)
    For every bounded subset, $A\subset M$, define:
    $D(\varepsilon,A)=\max\left\{\mathrm{card}(T)\colon T\subset A,\mathcal{d}(x,y)>\varepsilon\textit{ for all }x,y\in T,x\neq y\right\}.$
    There exist two constants, $\rho>0$ (which does not depend on $A$) and $K>0$, such that $D(\varepsilon,A)\leq K\varepsilon^{-d}$ for all $\varepsilon<\rho$.
Assumptions L.
Let $M$ be a connected Riemannian manifold, $f:M\rightarrow\mathbb{R}$ a continuous density function, and $l$ and $u$ two constants such that $0<l\leq u<\sup f$. There exist a radius, $r>0$, and two positive constants $k$ and $\delta$, with $k>0$ and $20<\delta<r/$ 2, verifying
  1. (L1)
    $L(\lambda)=\Psi_{r}[L(\lambda)]$ for all $\lambda\in[l-\delta,u+\delta]$.
  2. (L2)
    $L(\lambda)^{c}=\Psi_{\delta}\big{[}L(\lambda)^{c}\big{]}$ for all $\lambda\in[l-\delta,u+\delta]$.
  3. (L3)
    $L(\lambda)$ is a compact set for all $\lambda\in[l-\delta,u+\delta]$.
  4. (L4)
    $\mathcal{d}_{H}\big{(}L(\lambda),L(\lambda\pm\eta)\big{)}\leq k\eta$ for all $\lambda\in[l,u]$ when $\eta\in[0,\delta]$.
In addition, proving the uniform consistency of the level estimator $\lambda_{\gamma,n}$ (see equation (5.5)) requires a further assumption.
Assumption P.
Define $\lambda_{\gamma}=\sup\Big{\{}\lambda\in\mathbb{R}\colon\mathbb{P}\big{[}L(\lambda)\big{]}\geq 1-\gamma\Big{\}}.$ There exist $0<\underline{\gamma}\leq\overline{\gamma}$ such that
$\sup_{\gamma\in[\underline{\gamma},\overline{\gamma}]}\left|\lambda_{\gamma}-\lambda_{\gamma\pm\eta}\right|\leq k\eta$
for all $\eta\in[0,\delta]$, where $\delta$ and $k$ are the same constants as Assumptions L.
In this appendix, these Assumptions M, L and P are carefully investigated, and it is shown that they are related to the differentiability of the density function $f$ and the curvature of $M$. The contents of this appendix are divided in three different sections, each one devoted to a group of assumptions. Specifically, Section C.1 contains the proofs of Proposition 4.1 and Theorem 4.1, Section C.2 includes the proofs of Propositions 4.2 and 4.3, and Section C.3 is entirely dedicated to the proof of Proposition 5.1.
Since this supplement involves more technical results about Riemannian geometry than the main paper, some additional notation must be introduced before proceeding. Let $M$ be a connected and complete $d$-dimensional Riemannian manifold, and $\langle\cdot,\cdot\rangle_{M}$ its Riemannian metric. Given a point $x\in M$ and a tangent vector $v\in T_{x}M$, the norm of $v$ is denoted as $\|v\|_{M}=\sqrt{\langle v,v\rangle_{M}}$ (the notation $\left\|v\right\|$ will be also used when there is no possibility of confusion). Denote by $\exp_{x}:T_{x}M\rightarrow M$ the exponential map restricted at $x$ (see Lee,, 2018, Ch. 5, pp. 126–131). Denote by $i_{M}(x)$ the injectivity radius of $M$ at $x$, that is, the supremum of all $\delta>0$ such that $\exp_{x}$ is a diffeomorphism between $B_{\delta}(0)\subset T_{x}M$ and its image. Then, denote the injectivity radius of $M$ by $i_{M}=\sup_{x\in M}i_{M}(x)$. Given a two Riemannian manifolds, $M$ and $N$, and a smooth map $F:M\rightarrow N$, denote by $dF_{x}:T_{x}M\rightarrow T_{f(x)}N$ the differential of $F$ at $x$. Given a smooth function $f:M\rightarrow\mathbb{R}$, the gradient of $f$ is denoted by $\nabla f$.
C.1 About Assumptions M
This section is dedicated to the proofs of Proposition 4.1 and Theorem 4.1 in Section 4.1. For Proposition 4.1, an auxiliary lemma is required first.
Lemma C.1.
Let $M$ be a Riemannian manifold verifying Assumption (M2). Then,
$\mathrm{Vol}_{M}\big{(}B_{r}[x]\big{)}\geq a\min\{r,\rho\}^{d};$
for any $x\in M$, where $a$ and $\rho$ are the constants in Assumption (M2).
Proof.
This result is a direct consequence of Assumption (M2), taking $x_{1}=x_{2}=x$ and $r_{1}=r_{2}=\varepsilon=r$. ∎
Proof of Proposition 4.1.
Let $A$ be a bounded subset of $M$. Hence, there exist a point $z\in M$ and a radius $r>0$ such that $A\subset B_{r}(z)$. Let $\varepsilon\leq\rho$, and $T$ a subset of $A$ verifying:
$\mathcal{d}(x,y)>\varepsilon\text{ for all }x,y\in T,x\neq y.$ (C.1)
Then:
$\bigcup_{x\in T}B_{\varepsilon/2}(x)\subset B_{r+\rho}(z).$
Since $B_{\varepsilon/2}(x)\cap B_{\varepsilon/2}(y)=\emptyset$ for all $x,y\in T,x\neq y$, one deduces that
$\mathrm{Vol}_{M}\big{(}B_{r+\rho}(z)\big{)}\geq\mathrm{Vol}_{M}\left(\bigcup_{x\in T}B_{\varepsilon/2}(x)\right)=\sum_{x\in T}\mathrm{Vol}_{M}\big{(}B_{\varepsilon/2}(x)\big{)}.$
Then, Lemma C.1 yields
$\mathrm{Vol}_{M}\big{(}B_{r+\rho}(z)\big{)}\geq\sum_{x\in T}\mathrm{Vol}_{M}\big{(}B_{\varepsilon/2}(x)\big{)}\geq\mathrm{card}(T)\frac{a\varepsilon^{d}}{2^{d}}.$
Consequently,
$\mathrm{card}(T)\leq K\varepsilon^{-d},\qquad\text{with }K=\frac{2^{d}\mathrm{Vol}_{M}\big{(}B_{r+\rho}(z)\big{)}}{a}.$
The previous inequality holds for every $T\subset A$ verifying (C.1), so:
$D(\varepsilon,A)\leq K\varepsilon^{-d}.\qed$
The proof of Theorem 4.1 relies on the Rauch Theorem, which states that bounding the sectional curvatures of a manifold implies in turn bounding the norm of the differential of the exponential map. The version of the Rauch Theorem included here is taken form Dyer et al., (2015) (Lemma 8, p. 105). In the statement of the theorem the identification between the tangent spaces of a tangent space and the tangent space itself is implicitly used.
Theorem C.1 (Rauch Theorem, Dyer et al.,, 2015).
Assume that all sectional curvatures of $M$ are bounded by $\kappa^{-}\leq\kappa_{M}\leq\kappa^{+}$. Taking $v\in T_{x}M$ with $\left\|v\right\|=1$, one has for any $w\in T_{x}M,w\perp v$ that
$\frac{S_{\kappa^{+}}(r)}{r}\left\|w\right\|\leq\left\|(d\exp_{x})_{rv}w\right\|\leq\frac{S_{\kappa^{-}}(r)}{r}\left\|w\right\|$
for all $x\in M$ and all $r<\min\left\{i_{M},\dfrac{\pi}{2\sqrt{\max\{\left|\kappa^{-}\right|,\left|\kappa^{+}\right|\}}}\right\}$; where
$S_{\kappa}(r)=\left\{\begin{array}[]{ll}\frac{1}{\sqrt{\kappa}}\sin(\sqrt{\kappa}r),&\kappa>0\\ r,&\kappa=0\\ \frac{1}{\sqrt{\kappa}}\sinh(\sqrt{-\kappa}r),&\kappa<0.\end{array}\right.$
Since $d\exp_{x}$ preserves the length of any radial vector $v\in T_{x}M$, a bound on the distortion of the length of a vector orthogonal to $v$ means a bound for all tangent vectors. Furthermore, given any $\kappa>0$, one can show using Taylor’s theorem that:
$S_{\kappa}(r)\geq r-\frac{\kappa r^{3}}{6},$
for all $r\geq 0$. Therefore, the next result holds:
Corollary C.1.
Suppose that all sectional curvatures of $M$ are bounded by $0\leq\kappa_{M}<\kappa$. If $v\in T_{x}M$ is such that $\left\|v\right\|=r<\min\left\{i_{M},\frac{\pi}{2\sqrt{\kappa}}\right\}$, then for all $w\in T_{v}(T_{x}M)\simeq T_{x}M$ one has
$\left(1-\frac{\kappa r^{2}}{6}\right)\left\|w\right\|\leq\left\|(d\exp_{x})_{v}w\right\|\leq\left\|w\right\|$
for any $x\in M$.
Corollary C.1 clarifies how the exponential map distorts distance and volume. The following two lemmas explain how this distortion actually works.
Lemma C.2.
Suppose that $M$ is under the conditions of Theorem 4.1. Then, given any $x\in M$, the exponential mapping:
$\exp_{x}:B_{\rho}(0)\subset T_{x}M\longrightarrow M$
is a contraction. That is:
$\mathcal{d}\big{(}\exp_{x}(v_{1}),\exp_{x}(v_{2})\big{)}\leq\left\|v_{1}-v_{2}\right\|,\quad\forall v_{1},v_{2}\in B_{\rho}(0);$
where $\rho=\min\left\{i_{M},\frac{\pi}{2\sqrt{\kappa}}\right\}$.
Proof.
Given any $v_{1},v_{2}\in B_{\rho}(0)\subset T_{x}M$, consider the curve:
$\begin{array}[]{rcl}\alpha:[0,1]&\longrightarrow&B_{\rho}(0);\\ t&\longmapsto&(1-t)v_{1}+tv_{2}.\end{array}$
Then $\exp_{x}\circ\alpha$ is a curve from $\exp_{x}(v_{1})$ to $\exp_{x}(v_{2})$. Since the distance between two points on $M$ is smaller that the length of any curve joining them, one deduces that
[…]
Corollary C.1 guarantees that $\left\|(d\exp_{x})_{\alpha(t)}(\alpha^{\prime}(t))\right\|\leq\left\|\alpha^{\prime}(t)\right\|$ for all $t\in[0,1]$. So
[…]
Lemma C.3.
Suppose that $M$ is a $d$-dimensional manifold under the conditions of Theorem 4.1. Let $x$ be an arbitrary point on $M$ and $\exp_{x}$ the exponential mapping.
Any measurable set $A\subset B_{\rho}(0)\subset T_{x}M\simeq\mathbb{R}^{d}$ satisfies that
$\mathrm{Vol}_{M}\big{(}\exp_{x}(A)\big{)}\geq\left(1-\frac{\kappa\rho^{2}}{6}\right)^{d}\mathrm{Vol}_{\mathbb{R}^{d}}(A);$
where $\rho=\min\left\{i_{M},\frac{\pi}{2\sqrt{\kappa}}\right\}$.
Proof.
The mapping
$\begin{array}[]{rcl}\exp_{x}:A\subset\mathbb{R}^{d}&\longrightarrow&M\\ y&\longmapsto&\exp_{x}(y)\end{array}$
is a parametrization of the set $\exp_{x}(A)\subset M$. So:
[…]
where $G(y)$ is a $(d\times d)$ matrix with coefficients
$G_{ij}(y)=\big{\langle}(d\exp_{x})_{y}e_{i},(d\exp_{x})_{y}e_{j}\big{\rangle}_{M},$
$\langle\cdot,\cdot\rangle_{M}$ is the Riemannian metric of $M$, and $e_{i}\in\mathbb{R}^{d}$ denotes the $i$-th vector of the standard basis, i.e., $e_{i}$ is the vector with a $1$ in the $i$-th coordinate and 00’s elsewhere.
Corollary C.1 ensures that
$\left(1-\frac{\kappa\left\|y\right\|^{2}}{6}\right)\left\|w\right\|\leq\left\|(d\exp_{x})_{y}w\right\|\Rightarrow\left(1-\frac{\kappa\left\|y\right\|^{2}}{6}\right)^{2}\left\|w\right\|^{2}\leq\left\|(d\exp_{x})_{y}w\right\|^{2}\Rightarrow\\ \Rightarrow\left(1-\frac{\kappa\rho^{2}}{6}\right)^{2}\leq\left(1-\frac{\kappa\left\|y\right\|^{2}}{6}\right)^{2}\leq\frac{\left\|(d\exp_{x})_{y}w\right\|^{2}}{\left\|w\right\|^{2}}=\frac{w^{\prime}G(y)w}{w^{\prime}w}$
holds for all $y\in B_{\rho}(0)$ and all $w\in\mathbb{R}^{d}\simeq T_{x}M$. Hence, all the eigenvalues of $G(y)$ are larger than $\left(1-\kappa\rho^{2}/6\right)^{2}$, and, therefore,
$\det(G)\geq\left(1-\frac{\kappa\rho^{2}}{6}\right)^{2d}.$
Finally, one deduces that
[…]
Lemmas C.2 and C.3 imply Theorem 4.1. First, one proves that Assumption (M2) holds when $r_{1}$ and $r_{2}$ are small enough, and one then slowly extends the result for any radii. These results are collected in Lemmas C.4 and C.5 below.
Lemma C.4.
Suppose that $M$ is a $d$-dimensional manifold under the conditions of Theorem 4.1.
Let $20<r_{1},r_{2}\leq\rho/$ 2 with $\rho=\min\left\{i_{M},\frac{\pi}{2\sqrt{\kappa}}\right\}$. Let $x_{1},x_{2}$ be two points on $M$ such that $\mathfrak{e}=r_{1}+r_{2}-\mathcal{d}(x_{1},x_{2})>0$.
Then:
$\mathrm{Vol}_{M}\big{(}B_{r_{1}}[x_{1}]\cap B_{r_{2}}[x_{2}]\big{)}\geq ar^{(d-1)/2}\mathfrak{e}^{(d+1)/2}$
where $r=\min\{r_{1},r_{2}\}$ and $a>0$ is a constant.
Proof.
Consider the exponential map centered in $x_{1}$:
$\begin{array}[]{rcl}\exp_{x_{1}}:B_{\rho}(0)\subset\mathbb{R}^{d}&\longrightarrow&M\\ v&\longmapsto&\exp_{x_{1}}(v).\end{array}$
Let $v_{2}\in B_{\rho}(0)$ a preimage of $x_{2}$. Lemma C.2 guarantees that $\exp_{x_{1}}$ is a contraction in $B_{\rho}(0)$. Consequently,
$\exp_{x_{1}}\big{(}B_{r_{1}}[0]\cap B_{r_{2}}[v_{2}]\big{)}\subset B_{r_{1}}[x_{1}]\cap B_{r_{2}}[x_{2}].$
Hence, Lemma C.3 guarantees that
$\mathrm{Vol}_{M}\big{(}B_{r_{1}}[x_{1}]\cap B_{r_{2}}[x_{2}]\big{)}$ $\geq\mathrm{Vol}_{M}\Big{(}\exp_{x_{1}}\big{(}B_{r_{1}}[0]\cap B_{r_{2}}[v_{2}]\big{)}\Big{)}$
$\geq\left(1-\frac{\kappa\rho^{2}}{6}\right)^{d}\mathrm{Vol}_{\mathbb{R}^{d}}\big{(}B_{r_{1}}[0]\cap B_{r_{2}}[v_{2}]\big{)}.$
And Lemma 1 from Walther, (1997) yields
$\mathrm{Vol}_{M}\big{(}B_{r_{1}}[x_{1}]\cap B_{r_{2}}[x_{2}]\big{)}\geq\left(1-\frac{\kappa\rho^{2}}{6}\right)^{d}c_{d}r^{(d-1)/2}\mathfrak{e}^{(d+1)/2};$
where
$c_{d}=\frac{\pi^{(d-1)/2}}{2^{(d-3)/2}(d+1)\Gamma\big{(}1+(d-1)/2\big{)}}.$
Then, taking $a=\left(1-\frac{\kappa\rho^{2}}{6}\right)^{d}c_{d}$ ends the proof. ∎
Lemma C.5.
Suppose that $M$ is a $d$-dimensional manifold under the conditions of Theorem 4.1.
Let $r_{1},r_{2}>0$, and let $x_{1},x_{2}\in M$ such that $\mathfrak{e}=r_{1}+r_{2}-\mathcal{d}(x_{1},x_{2})>0$. Assume that $\mathfrak{e}\leq\rho/2$ with $\rho=\min\left\{i_{M},\frac{\pi}{2\sqrt{\kappa}}\right\}$.
Then:
$\mathrm{Vol}_{M}\big{(}B_{r_{1}}[x_{1}]\cap B_{r_{2}}[x_{2}]\big{)}\geq ar^{(d-1)/2}\mathfrak{e}^{(d+1)/2}$
where $r=\min\{r_{1},r_{2},\rho/2\}$ and $a>0$ is a constant.
Proof.
Since $\mathfrak{e}\leq r_{1}+r_{2}$ and $\mathfrak{e}\leq\rho/2$, it follows that
$\mathfrak{e}\leq\min\{r_{1}+r_{2},\rho/2\}\leq\tilde{r_{1}}+\tilde{r_{2}};$ (C.2)
where $\tilde{r_{1}}=\min\{r_{1},\rho/2\}$ and $\tilde{r_{2}}=\min\{r_{2},\rho/2\}$.
Let $\alpha$ be a geodesic segment from $x_{1}$ to $x_{2}$, $\alpha:\big{[}0,\mathcal{d}(x_{1},x_{2})\big{]}\rightarrow M$. From (C.2) one easily deduces that
$r_{1}-\tilde{r}_{1}\leq\mathcal{d}(x_{1},x_{2})\quad\text{and}\quad\mathcal{d}(x_{1},x_{2})-r_{2}+\tilde{r}_{2}\geq 0.$
$r_{1}-\tilde{r}_{1}$
$r_{2}-\tilde{r}_{2}$
$x_{1}$
$x_{2}$
$\tilde{x}_{1}$
$\tilde{x}_{2}$
Figure 11: The points $x_{1},\tilde{x}_{1},x_{2}$ and $\tilde{x}_{2}$ and the distance between them.
Therefore, one can define $\tilde{x}_{1}=\alpha\big{(}r_{1}-\tilde{r}_{1}\big{)}$ and $\tilde{x}_{2}=\alpha\big{(}\mathcal{d}(x_{1},x_{2})-(r_{2}-\tilde{r}_{2})\big{)}$ (check Figure 11). Then:
$\mathcal{d}(\tilde{x}_{1},\tilde{x}_{2})=\mathcal{d}(x_{1},x_{2})-(r_{2}-\tilde{r}_{2})-(r_{1}-\tilde{r}_{1})=\tilde{r}_{1}+\tilde{r}_{2}-\mathfrak{e}.$
Moreover, given any $x\in B_{\tilde{r}_{1}}(\tilde{x}_{1})\cap B_{\tilde{r}_{2}}(\tilde{x}_{2})$ one deduces
  • $\mathcal{d}(x,x_{1})\leq\mathcal{d}(x,\tilde{x}_{1})+\mathcal{d}(\tilde{x}_{1},x_{1})\leq\tilde{r}_{1}+r_{1}-\tilde{r}_{1}=r_{1}\Rightarrow x\in B_{r_{1}}(x_{1})$.
  • $\mathcal{d}(x,x_{2})\leq\mathcal{d}(x,\tilde{x}_{2})+\mathcal{d}(\tilde{x}_{2},x_{2})\leq\tilde{r}_{2}+r_{2}-\tilde{r}_{2}=r_{2}\Rightarrow x\in B_{r_{2}}(x_{2})$.
So, $B_{\tilde{r}_{1}}(\tilde{x}_{1})\cap B_{\tilde{r}_{2}}(\tilde{x}_{2})\subset B_{{r}_{1}}({x}_{1})\cap B_{{r}_{2}}({x}_{2})$ and
$\mathrm{Vol}_{M}\big{(}B_{{r}_{1}}({x}_{1})\cap B_{{r}_{2}}({x}_{2})\big{)}\geq\mathrm{Vol}_{M}\big{(}B_{\tilde{r}_{1}}(\tilde{x}_{1})\cap B_{\tilde{r}_{2}}(\tilde{x}_{2})\big{)}.$
Finally, Lemma C.4 guarantees that there exists a constant $a>0$ such that
$\mathrm{Vol}_{M}\big{(}B_{{r}_{1}}({x}_{1})\cap B_{{r}_{2}}({x}_{2})\big{)}$ $\geq\mathrm{Vol}_{M}\big{(}B_{\tilde{r}_{1}}(\tilde{x}_{1})\cap B_{\tilde{r}_{2}}(\tilde{x}_{2})\big{)}$
$\geq a\min\{r_{1},r_{2},\rho/2\}^{(d-1)/2}\mathfrak{e}^{(d+1)/2},$
concluding the proof. ∎
Proof of Theorem 4.1.
Take $\rho=\min\left\{i_{M},\frac{\pi}{2\sqrt{\kappa}}\right\}$. If $r_{1}+r_{2}-\mathcal{d}(x_{1},x_{2})\leq\rho/2$, the result is a direct consequence of Lemma C.5. Suppose then that $r_{1}+r_{2}-\mathcal{d}(x_{1},x_{2})>\rho/2$. We distinguish two cases:
  1. Case 1:
    If $r_{1}-\rho/4\geq\mathcal{d}(x_{1},x_{2})$, take
    $\tilde{r}_{1}=\mathcal{d}(x_{1},x_{2})+\rho/4,\quad\tilde{r}_{2}=\min\{r_{2},\rho/4\},\quad\mathfrak{e}=\tilde{r}_{1}+\tilde{r}_{2}-\mathcal{d}(x_{1},x_{2}).$
    Hence:
    $\mathfrak{e}=\mathcal{d}(x_{1},x_{2})+\rho/4+\min\{r_{2},\rho/4\}-\mathcal{d}(x_{1},x_{2})=\rho/4+\min\{r_{2},\rho/4\}\leq\rho/2.$
    So, Lemma C.5 guarantees that there exists a constant $a>0$ such that
    $\mathrm{Vol}_{M}\big{(}B_{\tilde{r}_{1}}[x_{1}]\cap B_{\tilde{r}_{2}}[x_{2}]\big{)}\geq a\min\{\tilde{r}_{1},\tilde{r}_{2},\rho/2\}^{(d-1)/2}\mathfrak{e}^{(d+1)/2}.$
    Furthermore, $\rho/4\leq\tilde{r}_{1}\leq r_{1}$ and $\rho/4\leq\tilde{r}_{2}\leq r_{2}$. Therefore:
    $\mathrm{Vol}_{M}\big{(}B_{r_{1}}[x_{1}]\cap B_{r_{2}}[x_{2}]\big{)}\geq$ $a\min\{\tilde{r}_{1},\tilde{r}_{2},\rho/2\}^{(d-1)/2}\mathfrak{e}^{(d+1)/2}$
    $\geq$ $a\min\{r_{1},r_{2},\rho/4\}^{(d-1)/2}\mathfrak{e}^{(d+1)/2}$
    $\geq$ $a\min\{r_{1},r_{2},\rho/4\}^{(d-1)/2}\min\{\varepsilon,\rho/4\}^{(d+1)/2};$
    where the last inequality follows from $\mathfrak{e}>\rho/4$.
  2. Case 2:
    If $r_{1}-\rho/4<\mathcal{d}(x_{1},x_{2})$, take
    $\tilde{r}_{2}=\rho/2+\mathcal{d}(p_{1},p_{2})-r_{1},\quad\mathfrak{e}=r_{1}+\tilde{r}_{2}-d(x_{1},x_{2}).$
    Hence:
    $\mathfrak{e}=r_{1}+\rho/2+\mathcal{d}(x_{1},x_{2})-r_{1}-\mathcal{d}(x_{1},x_{2})=\rho/2.$
    Then, Lemma C.5 ensures that there exists a constant $a>0$ such that
    $\mathrm{Vol}_{M}\big{(}B_{r_{1}}[x_{1}]\cap B_{\tilde{r}_{2}}[x_{2}]\big{)}\geq a\min\{r_{1},\tilde{r}_{2},\rho/2\}^{(d-1)/2}(\rho/2)^{(d+1)/2}.$
    From $r_{1}+r_{2}-\mathcal{d}(x_{1},x_{2})>\rho/2$ one deduces
    $\tilde{r}_{2}=\rho/2+\mathcal{d}(x_{1},x_{2})-r_{1}<r_{2}.$
    So,
    $\mathrm{Vol}_{M}\big{(}B_{r_{1}}[x_{1}]\cap B_{r_{2}}[x_{2}]\big{)}\geq a\min\{r_{1},\tilde{r}_{2},\rho/2\}^{(d-1)/2}(\rho/2)^{(d+1)/2}.$
    In addition, it follows from $r_{1}-\rho/4<\mathcal{d}(x_{1},x_{2})$ that:
    $\tilde{r}_{2}=\rho/2+\mathcal{d}(x_{1},x_{2})-r_{1}=\mathcal{d}(x_{1},x_{2})-(r_{1}-\rho/4)+\rho/4>\rho/4.$
    Consequently,
    $\mathrm{Vol}_{M}\big{(}B_{r_{1}}[x_{1}]\cap B_{r_{2}}[x_{2}]\big{)}\geq$ $a\min\{r_{1},\tilde{r}_{2},\rho/2\}^{(d-1)/2}(\rho/2)^{(d+1)/2}$
    $\geq$ $a\min\{r_{1},r_{2},\rho/4\}^{(d-1)/2}(\rho/2)^{(d+1)/2}$
    $\geq$ $a\min\{r_{1},r_{2},\rho/4\}^{(d-1)/2}\min\{\varepsilon,\rho/4\}^{(d+1)/2},$
    which ends the proof. ∎
C.2 About Assumptions L
This section contains the proofs of Propositions 4.2 and 4.3 in Section 4.2. We start with Proposition 4.2.
Proof of Proposition 4.2.
$M$ is a complete Riemannian manifold by assumption. So, the Hopf-Rinow Theorem (see Gallot et al.,, 2004, Cor. 2.105) ensures that every bounded and closed subset of $M$ is a compact set. $L(\lambda)$ is a closed set since $f$ is continuous. We will prove that is bounded by contradiction.
If $L(\lambda)=\emptyset$, $L(\lambda)$ is compact. Suppose that $L(\lambda)\neq\emptyset$ is not bounded. Let $x_{1}\in L(\lambda)$. Since $L(\lambda)$ is not bounded, there exists a point $x_{2}\in L(\lambda)$ such that $\mathcal{d}(x_{1},x_{2})>4\delta$. Iterating this argument, one gets a sequence $\{x_{n}\}_{n\in\mathbb{N}}\subset L(\lambda)$ verifying that
$\min_{1\leq i<n}\mathcal{d}(x_{n},x_{i})>4\delta,\quad\forall n\geq 2.$
By hypothesis, $L(\lambda)=\Psi_{\delta}\big{[}L(\lambda)\big{]}$. Then, for each $x_{n}\in L(\lambda)$ there exists a point $y_{n}\in L(\lambda)$ such that $x_{n}\in B_{\delta}[y_{n}]\subset L(\lambda)$. Thus, given $i,j\in\mathbb{N},i\neq j$:
$4\delta<\mathcal{d}(x_{i},x_{j})\leq\mathcal{d}(x_{i},y_{i})+\mathcal{d}(y_{i},y_{j})+\mathcal{d}(y_{j},x_{j})\leq\mathcal{d}(y_{i},y_{j})+2\delta.$
Consequently, $\mathcal{d}(y_{i},y_{j})>2\delta$ and $B_{\delta}[y_{i}]\cap B_{\delta}[y_{j}]=\emptyset$. Therefore, Lemma C.1 yields
$\mathrm{Vol}_{M}\big{[}L(\lambda)\big{]}\geq\mathrm{Vol}_{M}\bigg{(}\bigcup_{n\in N}B_{\delta}[y_{i}]\bigg{)}=\sum_{n\in\mathbb{N}}\mathrm{Vol}_{M}\big{(}B_{\delta}[y_{i}]\big{)}\geq\sum_{n\in\mathbb{N}}a\min\{r,\rho\}^{d}=+\infty;$
where $a,\rho>0$ are the constants present in Assumption (M2). But then
$\mathbb{P}\big{[}L(\lambda)\big{]}\geq\lambda\mathrm{Vol}_{M}\big{[}L(\lambda)\big{]}=+\infty,$
and that is a contradiction. ∎
The proof of Proposition 4.3 makes use of the integral flow generated by $\nabla f$ (see Lee,, 2012, Ch. 9, for an introduction to flows) to establish an explicit relation between the level sets of $f$ and the geodesic distance of $M$.
Proof of Proposition 4.3.
Let $\eta\leq\delta$ and $\lambda\in[l,u]$. By definition $L(\lambda+\eta)\subset L(\lambda)$. So, in order to prove $\mathcal{d}_{H}\big{(}L(\lambda),L(\lambda+\eta)\big{)}\leq\eta/c$, we only need to show that
$L(\lambda)\subset L(\lambda+\eta)\oplus(\eta/c)B.$
Let $p\in L(\lambda)$. Consider the following vector field in $U$:
$X=\frac{\nabla f}{\left\|\nabla f\right\|}.$
Let $\alpha$ be the maximal integral curve of $X$ such that $\alpha(0)=p$:
$\alpha:I\rightarrow U;\qquad\alpha^{\prime}(t)=\frac{\nabla f\big{(}\alpha(t)\big{)}}{\left\|\nabla f\big{(}\alpha(t)\big{)}\right\|},\quad\forall t\in I.$
Notice that $\left\|\alpha^{\prime}(t)\right\|=1$ for all $t\in I$. Then,
$\mathcal{d}\big{(}p,\alpha(t)\big{)}=\mathcal{d}\big{(}\alpha(0),\alpha(t)\big{)}\leq\left|t\right|,\quad\forall t\in I.$ (C.3)
Furthermore, given $t\in I$, we deduce:
$f\big{(}\alpha(t)\big{)}-\lambda$ $=f\big{(}\alpha(t)\big{)}-f\big{(}\alpha(0)\big{)}$
[…]
Hence,
$f\big{(}\alpha(t)\big{)}\geq\lambda+ct,\quad\forall t\in I.$ (C.4)
We distinguish two cases.
  • If $\eta/c\in I$, then (C.4) ensures that $\alpha(\eta/c)\in L(\lambda+\eta)$. Moreover, (C.3) guarantees that $\mathcal{d}\big{(}p,\alpha(\eta/c)\big{)}\leq\eta/c$. Hence, $p\in L(\lambda+\eta)\oplus(\eta/c)B$.
  • If $\eta/c\notin I$, then $\eta/c$ is an upper bound of $I$. The Escape Lemma (Theorem 9.20, Lee,, 2012) guarantees that $\alpha\big{(}I\cap[0,\eta/c]\big{)}\not\subset f^{-1}\big{(}[\lambda,\lambda+\eta]\big{)}$. I.e., there exists a $\tilde{t}\in I\cap[0,\eta/c]$ such that $\alpha(\tilde{t})\notin f^{-1}\big{(}[\lambda,\lambda+\eta]\big{)}$. Since $\tilde{t}\geq 0$, (C.4) ensures that $\alpha(\tilde{t})\in L(\lambda+\eta)$. In addition, $\mathcal{d}\big{(}p,\alpha(\tilde{t})\big{)}\leq\eta/c$ by (C.3). Consequently, $p\in L(\lambda+\eta)\oplus(\eta/c)B$.
The proof of $\mathcal{d}_{H}\big{(}L(\lambda),L(\lambda-\eta)\big{)}\leq\eta/c$ is totally analogous. ∎
C.3 About Assumption P
This section is fully devoted to the proof of Proposition 5.1 in Section 5.2. The basic idea is to make use of the flow generated by $\nabla f$ again, this time to connect the level sets of $f$ and the volume measure of $M$. This allows to prove that the function $\lambda\mapsto\mathbb{P}\big{[}L(\lambda)\big{]}$ is derivable, and its derivative is continuous and bounded away from zero. Then, one can translate that bound to the derivative of the function $\gamma\mapsto\lambda_{\gamma}$ through the Inverse Function Theorem.
First, Proposition C.1 below ensures the derivability of $\mathbb{P}\big{[}L(\lambda)\big{]}$.
Proposition C.1.
Let $M$ be a Riemannian manifold under Assumptions M, and let $f:M\rightarrow\mathbb{R}$ be a density function. Let $0<l\leq u<\sup f$, and let $U$ be an open set such that $f^{-1}\big{(}[l,u]\big{)}\subset U$. Suppose that $f$ is a $\mathcal{C}^{1}$ function in $U$ and there exists a positive constant $c>0$ satisfying such that $\left\|\nabla f\right\|\geq c$ in $U$.
Therefore, the function $H(\lambda)=P\big{[}L(\lambda)\big{]}$ is derivable in $(l,u)$, its derivative $H^{\prime}(\lambda)$ is continuous in $(l,u)$, and $H^{\prime}(\lambda)<0$ for all $\lambda\in(l,u)$.
Proof.
Let $\lambda\in(l,u)$. Then,
[…]
The basis of the proof is to rewrite the set $L(l)\setminus L(\lambda)$ in a more convenient way. Consider the following vector field in $U$
$X=\frac{\nabla f}{\left\|\nabla f\right\|^{2}}.$
Let $\Upsilon$ be the smooth maximal flow over $U$ with $X$ as infinitesimal generator:
$\Upsilon:\mathcal{D}\subset U\times\mathbb{R}\rightarrow U.$
Take $\mathcal{O}=\big{(}f^{-1}(l)\times\mathbb{R}\big{)}\cap\mathcal{D}$. The next lemma (that will be proven later on) guarantees that the restriction of the flow $\Upsilon$ to this set has nice properties.
Lemma C.6.
Consider the set $\mathcal{O}=\big{(}f^{-1}(l)\times\mathbb{R}\big{)}\cap\mathcal{D}$. Then:
  1. (a)
    The map $\Upsilon|_{\mathcal{O}}$ is a diffeomorphism between $\mathcal{O}$ and an open submanifold of $M$.
  2. (b)
    $f^{-1}(l)\times[0,u-l]\subset\mathcal{O}$.
  3. (c)
    The map $\Upsilon|_{\mathcal{O}}$ takes horizontal fibers to level sets of $f$. That is,
    $\Upsilon|_{\mathcal{O}}\big{(}f^{-1}(l)\times\{\eta\}\big{)}=f^{-1}(l+\eta),$
    for all $\eta\in[0,u-l]$.
In fact, $\Upsilon|_{\mathcal{O}}$ is the required reparametrization of $L(l)\setminus L(\lambda)$. For the sake of simplicity, denote $\Upsilon=\Upsilon|_{\mathcal{O}}$. Lemma C.6 guarantees that
[…]
where $G\equiv G(p,t)$ is the $m\times m$ matrix with coefficients
$G_{ij}:\mathcal{O}\longrightarrow M,\quad G_{ij}=\langle d\Upsilon E_{i},d\Upsilon E_{j}\rangle_{M}$
and $E_{1},\ldots,E_{n}$ is an orthonormal frame defined in a neighborhood of $(p,t)\in\mathcal{O}$ (the determinant of $G$ does not depend on the choice of this frame). Keep in mind that, since $d\Upsilon$ and $E_{i}$ are continuous, the functions $G_{ij}$ are also continuous in $\mathcal{O}$ for all $i,j\in\{1,\ldots,n\}$, and then $\sqrt{\left|\det G\right|}$ is also continuous in $\mathcal{O}$.
Lemma C.6c ensures that $f\big{(}\Upsilon(p,t)\big{)}=l+t$ for all $(p,t)\in f^{-1}(l)\times[0,\lambda-l)$. So,
[…]
Therefore,
[…]
for all $\lambda\in(l,u)$. So, $H$ is derivable at $\lambda$ and
[…]
Since $\sqrt{\left|\det G\right|}$ is continuous in $\mathcal{O}$ and $f^{-1}(l)\times\{\lambda-l\}\subset\mathcal{O}$ for all $\lambda\in(l,u)$ by Lemma C.6b, $H^{\prime}$ is continuous in $(l,u)$. Moreover, $\Upsilon$ is a diffeomorphism, so $\left|\det G\right|>0$ in $\mathcal{O}$. Hence, $H^{\prime}(\lambda)<0$ for all $\lambda\in(l,u)$.
The only thing left is to prove Lemma C.6.
Proof of Lemma C.6.
By assumption, $\nabla f\neq 0$ in $f^{-1}(l)$. So, $f^{-1}(l)$ is a regular submanifold of $M$ with codimension $1$. Furthermore, $X$ is not tangent to $f^{-1}(l)$ anywhere since the gradient is orthogonal to all the level sets. Hence, the Flowout Theorem (Lee,, 2012, Th. 9.20) guarantees that $\Upsilon|_{\mathcal{O}}$ is a diffeomorphism between $\mathcal{O}$ and an open submanifold of $M$.
The proof of (b) and (c) requires a deeper understanding of the relation between $f$ and the flow $\Upsilon$. Consider the set $\mathcal{D}^{(p)}=\left\{t\in\mathbb{R}\colon(p,t)\in\mathcal{D}\right\}$ and the map $\Upsilon^{(p)}:\mathcal{D}^{(p)}\rightarrow U$ defined by $\Upsilon^{(p)}(t)=\Upsilon(p,t)$. Given any $(p,t)\in\mathcal{D}$, we deduce
$f\big{(}\Upsilon(p,t)\big{)}-f(p)=$ $f\big{(}\Upsilon^{(p)}(t)\big{)}-f\big{(}\Upsilon^{(p)}(0)\big{)}$
[…]
So,
$f\big{(}\Upsilon(p,t)\big{)}=f(p)+t,\quad\forall(p,t)\in\mathcal{D}.$ (C.5)
Furthermore, it holds that:
$\big{[}l-f(p),u-f(p)\big{]}\subset\mathcal{D}^{(p)},\quad\forall p\in f^{-1}\big{(}[l,u]\big{)}.$ (C.6)
The latter result is proven by contradiction. Assume that $\big{[}l-f(p),u-f(p)\big{]}\not\subset\mathcal{D}^{(p)}$. Then, either $l-f(p)$ is a lower bound of $\mathcal{D}^{(p)}$, or $u-f(p)$ is an upper bound of $\mathcal{D}^{(p)}$. The reasoning is a little bit different (but analogous) in each case.
  • If $l-f(p)$ is a lower bound of $\mathcal{D}^{(p)}$, then the Escape Lemma (Lee,, 2012, Lemma 9.19) guarantees that $\Upsilon^{(p)}\big{(}\mathcal{D}^{(p)}\cap(l-f(p),0]\big{)}$ is not contained in any compact subset of $U$. But, at the same time, (C.5) ensures that
    $\Upsilon^{(p)}\big{(}\mathcal{D}^{(p)}\cap(l-f(p),0]\big{)}\subset f^{-1}\big{(}[l,u]\big{)},$
    which is clearly a contradiction.
  • If $u-f(p)$ is an upper bound of $\mathcal{D}^{(p)}$, then the Escape Lemma again ensures that $\Upsilon^{(p)}\big{(}\mathcal{D}^{(p)}\cap[0,u-f(p))\big{)}$ is not contained in any compact subset of $U$. However, it follows from (C.5) that
    $\Upsilon^{(p)}\big{(}\mathcal{D}^{(p)}\cap[0,u-f(p))\big{)}\subset f^{-1}\big{(}[l,u]\big{)},$
    and that is a contradiction.
Once (C.5) and (C.6) are known, the proof of (b) and (c) is straightforward. In fact, (b) is a direct consequence of (C.6). Moreover, (C.5) yields
$\Upsilon|_{\mathcal{O}}\big{(}f^{-1}(l)\times\{\eta\}\big{)}=\Upsilon\big{(}f^{-1}(l)\times\{\eta\}\big{)}\subset f^{-1}(l+\eta).$
Take $p\in f^{-1}(l+\eta)$. From (C.6), it follows that $-\eta\in\mathcal{D}^{(p)}$. Hence, $\Upsilon(p,-\eta)$ exists and (C.5) ensures that
$f\big{(}\Upsilon(p,-\eta)\big{)}=f(p)-\eta=l+\eta-\eta=l.$
Then, $\Upsilon(p,-\eta)\in f^{-1}(l)$. Furthermore,
$\Upsilon\big{(}\Upsilon(p,-\eta),\eta\big{)}=\Upsilon(p,0)=p,$
which shows that $p\in\Upsilon\big{(}f^{-1}(l)\times\{\eta\}\big{)}$, completing the proof. ∎
Finally, the proof of Proposition 5.1 uses the Inverse Function Theorem to translate this bound of the derivative of $\mathbb{P}\big{[}L(\lambda)\big{]}$ into a bound of the derivative of the function $\gamma\mapsto\lambda_{\gamma}$.
Proof of Proposition 5.1.
Proposition C.1 and Inverse Function Theorem guarantee that $H(\lambda)=\mathbb{P}\big{[}L(\lambda)\big{]}$ has a derivable inverse function, $H^{-1}:\big{(}H(u),H(l)\big{)}\rightarrow(l,u)$. We will show that this inverse function of $H$ is closely related to $\lambda_{\gamma}$.
Let $\gamma\in\big{[}\underline{\gamma}-\delta,\overline{\gamma}+\delta\big{]}$. From the definition of $\lambda_{\gamma}$ (see (5.5)) it follows that $H(\lambda_{\gamma})\geq 1-\gamma$. Assume by contradiction that $H(\lambda_{\gamma})>1-\gamma$. Since $\lambda_{\gamma}\in[\lambda_{\underline{\gamma}-\delta},\lambda_{\overline{\gamma}+\delta}]\subset(l,u)$, $H$ is continuous in $\lambda_{\gamma}$. Then, there exists a constant $\epsilon>0$ such that
$H(\lambda_{\gamma}+\eta)>1-\gamma,\quad\forall\eta\in[0,\epsilon]$
which contradicts the definition of $\lambda_{\gamma}$. Therefore, $H(\lambda_{\gamma})=1-\gamma$ and consequently $\lambda_{\gamma}=H^{-1}(1-\gamma)$ for all $\gamma\in\big{[}\underline{\gamma}-\delta,\overline{\gamma}+\delta\big{]}$. Hence, $\lambda_{\gamma}$ is derivable with respect to $\gamma$ and its derivative is
$\frac{d\lambda_{\gamma}}{d\gamma}=(H^{-1})^{\prime}(1-\gamma)=\frac{-1}{H^{\prime}(\lambda_{\gamma})},$
which is a continuous function in $\big{[}\underline{\gamma}-\delta,\overline{\gamma}+\delta\big{]}$.
Now consider the function:
$F(\gamma,\eta)=\left\{\begin{array}[]{ll}\dfrac{\lambda_{\gamma+\eta}-\lambda_{\gamma}}{\eta},&\text{if }\eta\in[-\delta,\delta]\setminus\{0\};\\[14.22636pt] \dfrac{-1}{H^{\prime}(\lambda_{\gamma})},&\text{if }\eta=0.\end{array}\right.$
From the continuity of $\lambda_{\gamma}$ and $H^{\prime}(\lambda_{\gamma})$, and the differentiability of $\lambda_{\gamma}$, it follows that $F$ is a continuous function in $[\underline{\gamma},\overline{\gamma}]\times[-\delta,\delta]$. Then, by compactness, there exists a constant such that $\sup_{(\gamma,\eta)\in[\underline{\gamma},\overline{\gamma}]\times[-\delta,\delta]}\left|F(\gamma,\eta)\right|\leq k$, which implies
$\sup_{\gamma\in[\underline{\gamma},\overline{\gamma}]}\left|\lambda_{\gamma+\eta}-\lambda_{\gamma}\right|\leq k\left|\eta\right|$
for all $\eta\in[-\delta,\delta]$. ∎